### v1

# Channel-Imposed Fusion: A Simple yet Effective Method for Medical Time Series Classification

Abstract
The automatic classification of medical time series signals, such as electroencephalogram (EEG) and electrocardiogram (ECG), plays a pivotal role in clinical decision support and early detection of diseases. Although Transformer based models have achieved notable performance by implicitly modeling temporal dependencies through self-attention mechanisms, their inherently complex architectures and opaque reasoning processes undermine their trustworthiness in high stakes clinical settings. In response to these limitations, this study shifts focus toward a modeling paradigm that emphasizes structural transparency, aligning more closely with the intrinsic characteristics of medical data. We propose a novel method, Channel Imposed Fusion (CIF), which enhances the signal-to-noise ratio through cross-channel information fusion, effectively reduces redundancy, and improves classification performance. Furthermore, we integrate CIF with the Temporal Convolutional Network (TCN), known for its structural simplicity and controllable receptive field, to construct an efficient and explicit classification framework. Experimental results on multiple publicly available EEG and ECG datasets demonstrate that the proposed method not only outperforms existing state-of-the-art (SOTA) approaches in terms of various classification metrics, but also significantly enhances the transparency of the classification process, offering a novel perspective for medical time series classification.
1 Introduction
Medical time series (MedTS) data, such as electroencephalogram (EEG) and electrocardiogram (ECG) signals, are widely used in clinical settings to monitor patient health and play a crucial role in diagnosing neurological and cardiovascular diseases Arif et al. (2024); Xiao et al. (2023); Zhu et al. (2025). Accurate classification of these signals enables early anomaly detection, personalized treatment, and optimized therapy planning, ultimately improving patient outcomes and healthcare efficiency Liu et al. (2024a); Tian et al. (2023). With advances in deep learning, CNN-based models like EEGNet Lawhern et al. (2018) can automatically extract informative features from raw signals, significantly improving classification performance.
In recent years, Transformer models Vaswani et al. (2017), powered by the self-attention mechanism Bahdanau et al. (2014), have achieved remarkable progress in time-series modeling, particularly in capturing long-range dependencies and global contextual information Liu et al. (2021); Zhou et al. (2021). By converting sequential data into high-dimensional token embeddings, Transformers are capable of modeling complex temporal dependencies implicitly. However, despite their strong performance across a range of time-series tasks, several limitations have emerged when applying Transformer-based architectures to medical time-series classification, particularly in the following aspects: (1) Conflict Between Implicit Temporal Modeling and Explicit Physiological Mechanisms. Conventional Transformer architectures employ self attention mechanisms to implicitly represent temporal dependencies . However, essential physiological dynamics in medical signals, including conduction delays between ECG leads and synchronization of EEG rhythms, require explicit structural representations. Although methods such as Medformer Wang et al. (2024) have adopted strategies involving multiple scales of embedding, their attention distributions still fail to directly correspond to the underlying biological processes. This leads to fundamental limitations in model interpretability. (2) Mismatch Between Medical Domain Knowledge and Architectures Designed for General Purposes. Mainstream time series modeling methods such as Autoformer Wu et al. (2021), Informer Zhou et al. (2021), and FEDformer Zhou et al. (2022) have achieved strong performance in general domains like weather forecasting and finance. However, their architectural designs fail to account for the unique structural characteristics of medical time series. Medical signals often exhibit high inter-channel redundancy, spatial dependencies among physiological leads, and distinct local morphological patterns. These important features are easily weakened or overlooked in the high-dimensional and implicit modeling processes of conventional Transformer architectures.
To address these limitations, we rethink the modeling paradigm from the perspective of data representation and break away from the conventional approach of implicitly discovering unknown temporal relationships. Unlike existing methods that rely on implicit learning, we advocate for explicitly modeling temporal relationships that have clear physical or physiological foundations. Based on this principle, we propose the Channel-Imposed Fusion (CIF) method, which encodes prior causal structures directly into the feature representation. Specifically, CIF constructs new features through a linear combination of the form , where and represent signals from two different channels, and and are predefined coefficients derived from domain knowledge. By embedding such structured prior knowledge into the model, CIF not only enhances the ability to capture critical pathological patterns but also significantly improves robustness and interpretability when dealing with noisy or irregular medical time-series data. This approach provides a more intuitive and effective framework for medical time-series analysis, contributing to the development of reliable clinical decision support systems.
Based on the design philosophy of the CIF method, we aim for the model to maintain high efficiency while ensuring good interpretability and simplifying the structure as much as possible. To this end, we propose a simple yet effective Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN) architecture. This architecture utilizes a hierarchical causal convolution module, which explicitly encodes the coupling relationships between channels and temporal dynamics, effectively avoiding the issues of poor interpretability and structural complexity caused by the implicit modeling of temporal dependencies in Transformer-based methods. Moreover, it is worth noting that the CIF method is not limited to the HM-BiTCN architecture; it has high portability and can be seamlessly integrated with existing Transformer architectures to enhance the model’s adaptability and interpretability in medical time-series data. Our contributions are as follows:
(1) Proposing the Channel-Imposed Fusion (CIF) method: We propose the CIF method to explicitly model the relationships between channel dimensions in medical time-series data, particularly for signals with clear physiological structures, such as EEG and ECG. (2) Designing the HM-BiTCN model based on CIF: We leverage the CIF method to construct the Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN), which explicitly models channel coupling and temporal dependencies. By adopting this TCN-based approach, HMBiTCN breaks the dominance of transformer models (SOTA) in MedTS classification and surpasses them across multiple datasets. (3) Providing methodological transferability: The CIF method is not only applicable to the HM-BiTCN model but can also be seamlessly integrated into existing Transformer architectures.
2 Related Work
Medical Time Series Classification. Medical time series analysis diverges fundamentally from general time series forecasting Wu et al. (2022); Lu et al. (2024) by prioritizing pathological signature decoding over temporal extrapolation, with modalities like EEG Tang et al. (2021); Yang et al. (2023); Qu et al. (2020), ECG Xiao et al. (2023); Wang et al. (2023); Kiyasseh et al. (2021), and EMG [Xiong et al. (2021); Dai et al. (2022)] encoding distinct clinical semantics. Early methods were dominated by compact CNNs such as EEGNet Lawhern et al. (2018), which employs depthwise separable convolutions to efficiently extract spatio–temporal features while providing preliminary interpretability via feature-map visualization. Subsequently, temporal convolutional networks (TCNs) Bai et al. (2018); Lin et al. (2019) leveraging dilated causal convolutions achieved parallelizable computation and extended receptive fields, surpassing LSTM-based approaches Zhou et al. (2016); Shen and Lee (2016); Hochreiter and Schmidhuber (1997) on multiple medical signal classification benchmarks. Hybrid architectures such as EEG-Conformer Song et al. (2022) combined convolutional front-ends with Transformer self-attention to capture both local and global dependencies and enabled attention-based interpretability. More recently, fine-grained Transformer models such as Medformer Wang et al. (2024) introduced cross-channel tokenization and dual-stage self-attention, setting new SOTA accuracy on several public datasets. The latest MedGNN Fan et al. (2025) further augments attention mechanisms with multi-resolution graph learning to jointly model spatial multi-scale channel dependencies and temporal dynamics. Although Transformer models have significantly improved classification accuracy, their complexity, especially due to deep architectures and multi-layer attention mechanisms, results in reduced interpretability. In clinical applications, where model transparency and reliability are essential, it is imperative to strike a balance between high accuracy and explainability.
Implicit modeling strategies in Transformer-based temporal methods. In time series analysis, Transformer-based models implicitly learn complex dependencies through diversified architectural designs: the vanilla Transformer Vaswani et al. (2017) first introduced multi-head self-attention and sinusoidal positional encoding to model temporal correlations globally; Informer Zhou et al. (2021) employs ProbSparse attention to select key time steps and compress sequence length, thereby reducing the computational cost of long-range dependencies; Reformer Kitaev et al. (2019) incorporates Locality-Sensitive Hashing (LSH) to reduce attention complexity to , making it suitable for ultra-long sequences; Autoformer Wu et al. (2021) proposes an Auto-Correlation mechanism that aggregates periodic subsequences to enhance the implicit capture of cyclic patterns; FEDformer Zhou et al. (2022) performs seasonal–trend decomposition in the frequency domain and uses compressed Fourier coefficients to enable cross-frequency attention interactions; Crossformer Zhang and Yan (2022) designs a two-stage attention mechanism across time and feature dimensions to implicitly fuse multivariate spatiotemporal couplings; iTransformer Liu et al. (2024b) innovatively treats time steps as channel dimensions and applies standard attention to implicitly learn nonlinear inter-variable relationships; PatchTST Nie et al. (2023) segments continuous time steps into patch-based tokens and uses a combination of local and global attention to capture multi-scale temporal patterns; Medformer Wang et al. (2024) introduces multi-granularity patch embeddings and cross-channel attention for medical signals, implicitly modeling the heterogeneous couplings of physiological metrics; and MedGNN Fan et al. (2025) combines graph attention with frequency-differential networks to incorporate medical topological priors into implicit spatiotemporal dependency learning. Although these variants enhance time series modeling via attention-based techniques, their reliance on high-dimensional weight matrices to implicitly encode dependencies between time steps and variables hinders interpretability. This limitation is particularly severe in medical applications, where it undermines the trustworthiness of high-risk tasks such as disease diagnosis and risk prediction.
3 Preliminaries and Problem Formulation
We define the MedTS classification task as a supervised learning problem over time series data with hierarchical subject-level structure. The dataset is represented as , where is the -th sample with time steps and channels, is a multi-label vector over categories, and is the subject ID. Samples from the same subject share the same label, i.e., , where represents the set of samples from subject . We consider two dataset partitioning strategies: (i) Subject-Dependent Split, where the dataset is split at the sample level such that samples from the same subject may appear in both training and test sets, which can cause information leakage and yield overly optimistic performance estimates; and (ii) Subject-Independent Split, where the dataset is split at the subject level, ensuring that each subject appears only in one of the train, validation, or test sets, simulating real-world diagnostic scenarios but introducing challenges due to inter-subject variability. The goal is to train an encoder to obtain a latent representation , which is then passed to a classifier to predict the label .
4 Method
4.1 Channel-Imposed Fusion
In this section, we analyze the data fusion process using Singular Value Decomposition (SVD) Golub and Reinsch (1971); Klema and Laub (1980). Suppose the original time-series data matrix has dimensions , where represents the temporal length and denotes the number of channels. We partition along the channel dimension into two submatrices: the first channels form , while the last channels form . The goal of fusion is to combine these two submatrices into a new representation:
where and are learnable parameters that control the contribution of the front and back segments in the fused matrix. To understand this fusion process, we first apply SVD to both submatrices:
where are the left singular vectors representing temporal patterns, are diagonal matrices containing the singular values that indicate the importance of each channel, and are the right singular vectors capturing channel relationships. The fused matrix is then constructed by linearly combining these two matrices. Depending on the degree of similarity between and , the fusion process either reduces redundancy or increases data diversity. When the correlation between and is high, their singular values and left singular vectors and are similar. This implies that both matrices capture similar temporal patterns, and the fused matrix primarily retains these common structures. Mathematically, if , the fused matrix can be approximated as:
In this case, is largely determined by the shared temporal patterns , while the weighted sum of and reflects the contribution of each channel. Since the temporal patterns are similar, the fusion process does not introduce significant new information, effectively reducing redundancy. On the other hand, when and have low correlation, their singular values and left singular vectors and differ significantly. This means that the two matrices capture distinct temporal patterns, and the fused matrix will combine these different components, enhancing data diversity. Mathematically, if and are dissimilar, the fused matrix can be expressed as:
In this scenario, the fused matrix incorporates both the temporal patterns from and , effectively combining the complementary information from both matrices. This leads to an increase in data diversity, as each matrix contributes distinct temporal patterns and channel relationships to the final fused representation. Thus, the fusion process allows for either redundancy reduction by preserving shared patterns or diversity enhancement by integrating complementary information, depending on the correlation between the front and back segments.
Optimizing SNR via CIF: Let two observed signals be and , where are zero-mean useful signals with variance and correlation coefficient , and are zero-mean noise components with variance and correlation coefficient . Assume that the signal and noise components are mutually uncorrelated. When applying the CIF operation via a linear combination , the output signal-to-noise ratio (SNR) becomes (details in Appendix A):
where . When the numerator exceeds the denominator (i.e., ), the SNR improves. This phenomenon occurs in two distinct modes: the Difference Mode ( and ) and the Cooperative Mode ( and ). In the Difference Mode, the correlation of noise sources (e.g., ocular artifacts) is higher than that of the signal, enhancing the SNR by suppressing correlated noise. In the Cooperative Mode, task-related brain regions exhibit synchronized activity while noise remains uncorrelated, leading to an increase in SNR by constructively accumulating and amplifying task-relevant signals. Both modes optimize the SNR by adjusting parameters and , showcasing CIF’s ability to suppress noise and enhance the signal under varying correlations.
Gradient Coupling Mechanism: The coefficients and are optimized through gradient descent, where their updates are coupled due to shared dependencies on and . The gradients as:
Specifically, the gradient computations of the loss function with respect to and include information from each other’s signals, requiring their updates to simultaneously consider the influence of both forward and backward channels. This coupling mechanism enables the model to learn cross-channel interaction patterns, enhancing feature representation and generalization capability.
In Appendix B, we visualize scenarios with highly and generally correlated channel information.
4.2 HM-BiTCN Structure Design and Theoretical Analysis
Modeling short-term and long-term dependencies in time series data is challenging. Traditional CNNs excel at capturing local features but have limited receptive fields, hindering long-range dependency learning. Transformer-based methods effectively model long-term dependencies, but their complex design lacks interpretability, which is a key issue in medical time-series classification.To address these limitations, TCNs use causal convolutions for explicit temporal modeling and dilated convolutions to expand the receptive field, overcoming the constraints of traditional CNNs. Building on the advantages of TCN, we propose the HM-BiTCN, which combines the benefits of dilated convolutions, bidirectional causal convolution, and residual connections. This approach allows for better capture of temporal dependencies while preserving causality. Our goal is to simplify the model structure through explicit design principles while improving its interpretability.
4.2.1 Dilated Convolution
Dilated convolution expands the receptive field without significantly increasing computational cost Yu and Koltun (2015). For a 1D input sequence , its output is defined as , where is the current time step, is the kernel size, is the dilation factor, and is the weight at the -th position in the kernel. Increasing effectively enlarges the receptive field, enabling the network to capture longer-term temporal dependencies. When stacking multiple dilated convolutional layers, the receptive field grows progressively. For the -th layer, the receptive field can be expressed as , where is the dilation factor of the -th layer. By gradually increasing , the network captures temporal dependencies across both global and local scales, offering an effective way to model long-term dependencies in time series.
4.2.2 Bidirectional Causal Convolution Structure
In addition to dilated convolutions, HM-BiTCN introduces a bidirectional causal convolution structure, inspired by prior bidirectional temporal modeling approaches Hanson et al. (2018); Hu et al. (2024); Yin et al. (2025). Unlike traditional TCNs that use only forward causal convolutions, our architecture applies causal convolutions in both forward and backward directions, enabling the model to capture dependencies from both past and future contexts while strictly preserving causality. The forward causal convolution processes the input sequence in chronological order, producing output , which depends only on current and past inputs. For the backward causal convolution, we first reverse the input sequence as , and then apply a causal convolution over this flipped sequence. This ensures that the model captures future-directed dependencies without introducing information leakage. The output is given by . These two operations are implemented using separate convolutional layers (convforward and convbackward), and their outputs are summed to form the final bidirectional result: . By integrating both directions under strict causality constraints, HM-BiTCN achieves superior temporal dependency modeling compared to unidirectional causal approaches.
4.2.3 Multi-Scale Feature Learning and Residual Connections
To further improve the model’s capacity to capture dependencies at different temporal scales, HM-BiTCN incorporates multi-scale feature learning and residual connections. Multi-scale Feature Learning: In HM-BiTCN, we employ a hierarchy of dilation factors that decrease layer by layer to capture temporal dependencies at multiple scales. Lower layers use larger dilation factors to expand the receptive field, aggregating long-range information and smoothing short-term noise in highly redundant medical time series; higher layers use smaller dilation factors to focus on local dependencies and capture fine-grained features. This coarse-to-fine, global-to-local design enables the network to extract broad patterns in its initial layers and refine precise details in its later layers, thereby enhancing adaptability across a wide range of time series tasks. Residual connections: Residual connections He et al. (2016) are introduced between the dilated convolutional layers to facilitate the efficient flow of information through the network. The residual connection is defined as , where is the convolutional output, and is the input. This design alleviates the vanishing gradient problem and improves the overall stability of the network during training.
5 Experiments
Datasets. (1) APAVA Escudero et al. (2006) is an EEG dataset where each sample is assigned a binary label indicating whether the subject has Alzheimer’s disease. (2) TDBrain van Dijk et al. (2022) is an EEG dataset with a binary label assigned to each sample, indicating whether the subject has Parkinson’s disease. (3) ADFTD Miltiadous et al. (2023b, a) is an EEG dataset with a three-class label for each sample, categorizing the subject as Healthy, having Frontotemporal Dementia, or Alzheimer’s disease. (4) PTB PhysioBank (2000) is an ECG dataset where each sample is labeled with a binary indicator of Myocardial Infarction. (5) PTB-XL Wagner et al. (2020) is an ECG dataset with a five-class label for each sample, representing various heart conditions. Table 4 provides information on the processed datasets.The data processing methodology is the same as that of Medformer Wang et al. (2024) and MedGNN Fan et al. (2025).
Baselines. We compare with 12 state-of-the-art time series transformer methods: Autoformer Wu et al. (2021), Crossformer Zhang and Yan (2022), FEDformer Zhou et al. (2022), Informer Zhou et al. (2021), iTransformer Liu et al. (2024b), MTST Zhang et al. (2024), Nonformer Liu et al. (2022), PatchTST Nie et al. (2023), Reformer Kitaev et al. (2019), vanilla Transformer Vaswani et al. (2017), Medformer Wang et al. (2024), MedGNN Fan et al. (2025).
Implementation. We employ six evaluation metrics: accuracy, precision (macro-averaged), recall (macro-averaged), F1 score (macro-averaged), AUROC (macro-averaged), and AUPRC (macro-averaged). The training process is conducted with five random seeds (41-45) on fixed training, validation, and test sets to compute the mean and standard deviation of the models. All experiments were conducted using an NVIDIA RTX 3090 GPU and implemented with PyTorch version 1.11.0 Paszke et al. (2017).
5.1 Results of Subject-Dependent
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| ADFTD (3-Classes) Reported | Autoformer Wu et al. (2021) | 87.831.62 | 87.631.66 | 87.221.97 | 87.381.79 | 96.590.88 | 93.821.64 |
| Crossformer Zhang and Yan (2022) | 89.351.32 | 89.001.44 | 88.791.37 | 88.881.40 | 97.520.58 | 95.451.03 | |
| FEDformer Zhou et al. (2022) | 77.632.37 | 76.762.17 | 76.682.48 | 76.602.46 | 91.671.34 | 84.942.11 | |
| Informer Zhou et al. (2021) | 90.930.90 | 90.740.71 | 90.501.14 | 90.600.94 | 98.190.27 | 96.510.49 | |
| iTransformer Liu et al. (2024b) | 64.900.25 | 62.530.27 | 62.210.26 | 62.250.33 | 81.520.29 | 68.870.49 | |
| MTST Zhang et al. (2024) | 65.080.69 | 63.850.80 | 62.710.64 | 63.030.58 | 81.360.56 | 69.340.89 | |
| Nonformer Liu et al. (2022) | 96.120.47 | 95.940.56 | 95.990.38 | 95.960.47 | 99.590.09 | 99.080.16 | |
| PatchTST Nie et al. (2023) | 66.260.40 | 65.080.41 | 64.970.51 | 64.950.42 | 83.070.45 | 71.700.61 | |
| Reformer Kitaev et al. (2019) | 91.511.75 | 91.151.79 | 91.651.56 | 91.141.83 | 98.850.35 | 97.880.60 | |
| Transformer Vaswani et al. (2017) | 97.000.43 | 96.870.53 | 96.860.36 | 96.860.44 | 99.750.04 | 99.420.07 | |
| Medformer Wang et al. (2024) | 97.620.34 | 97.530.33 | 97.480.40 | 97.500.36 | 99.830.05 | 99.620.12 | |
| MedGNN Fan et al. (2025) | 98.420.04 | 98.310.02 | 98.290.05 | 98.300.12 | 99.930.11 | - | |
| Medformer + CIF | 98.870.26 | 98.770.27 | 98.860.27 | 98.810.27 | 99.960.01 | 99.920.03 | |
| MedGNN + CIF | 99.600.09 | 99.600.11 | 99.580.09 | 99.590.10 | 99.990.01 | 99.970.01 |
To further validate our proposed method, we reproduced 12 baseline approaches. Table 1 presents a comparison of the reported experimental results for these methods, while Table 5 in Appendix E provides detailed results from our reproduction of the 12 methods. The experimental results demonstrate that by incorporating our proposed CIF method into the Medformer model, we not only achieve significantly better performance than the original methods, but also outperform the current state-of-the-art (SOTA) approach, MedGNN. These findings strongly confirm the effectiveness and superiority of our method.
5.2 Results of Subject-Independent
In the Subject-Independent setup, the training, validation, and test sets are split based on subjects. All subjects and their corresponding samples are distributed into these three sets according to a predetermined ratio or subject IDs. Samples from the same subject exclusively appear in one of the three sets. This setup simulates real-world MedTS-based disease diagnosis, aiming to train a model on subjects with known labels and then test it on unseen subjects to determine if they have a specific disease. All five datasets are evaluated under this setup.
Table 2 presents the results reported by various methods in the subject-independent setting, while Table 6 shows the results of our reproduction of these methods. Our method achieves the highest scores in six metrics across four out of the five datasets. For the PTB-XL dataset, compared to the results reported by other methods, our approach achieves the highest AUROC and AUPRC score and ranks second in Accuracy. When compared to the results of our reproduction, our method ranks first in Accuracy, Precision, AUROC, and AUPRC, and second in Recall. Additionally, it is worth noting that in the subject-independent setup, the F1 score of ADFTD is 55.42%, which is significantly lower than the 99.59% achieved in the subject-dependent setup. This comparison highlights the challenges of the subject-independent setup, which better simulates real-world scenarios.
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| APAVA (2-Classes) Reported | Autoformer Wu et al. (2021) | 68.641.82 | 68.482.10 | 68.772.27 | 68.061.94 | 75.943.61 | 74.384.05 |
| Crossformer Zhang and Yan (2022) | 73.771.95 | 79.294.36 | 68.861.70 | 68.931.85 | 72.393.33 | 72.053.65 | |
| FEDformer Zhou et al. (2022) | 74.942.15 | 74.591.50 | 73.563.55 | 73.513.39 | 83.721.97 | 82.942.37 | |
| Informer Zhou et al. (2021) | 73.114.40 | 75.176.06 | 69.174.56 | 69.475.06 | 70.464.91 | 70.755.27 | |
| iTransformer Liu et al. (2024b) | 74.551.66 | 74.772.10 | 71.761.72 | 72.301.79 | 85.591.55 | 84.391.57 | |
| MTST Zhang et al. (2024) | 71.141.59 | 79.300.97 | 65.272.28 | 64.013.16 | 68.872.34 | 71.061.60 | |
| Nonformer Liu et al. (2022) | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |
| PatchTST Nie et al. (2023) | 67.031.65 | 78.761.28 | 59.912.02 | 55.973.10 | 65.650.28 | 67.990.76 | |
| Reformer Kitaev et al. (2019) | 78.702.00 | 82.503.95 | 75.001.61 | 75.931.82 | 73.941.40 | 76.041.14 | |
| Transformer Vaswani et al. (2017) | 76.304.72 | 77.645.95 | 73.095.01 | 73.755.38 | 72.506.60 | 73.237.60 | |
| Medformer Wang et al. (2024) | 78.740.64 | 81.110.84 | 75.400.66 | 76.310.71 | 83.200.91 | 83.660.92 | |
| MedGNN Fan et al. (2025) | 82.600.35 | 87.700.22 | 78.930.09 | 80.250.16 | 85.930.26 | - | |
| HM-BiTCN + CIF | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | |
| TDBrain (2-Classes) Reported | Autoformer Wu et al. (2021) | 87.333.79 | 88.063.56 | 87.333.79 | 87.263.84 | 93.812.26 | 93.322.42 |
| Crossformer Zhang and Yan (2022) | 81.562.19 | 81.972.25 | 81.562.19 | 81.502.20 | 91.201.78 | 91.511.71 | |
| FEDformer Zhou et al. (2022) | 78.131.98 | 78.521.91 | 78.131.98 | 78.042.01 | 86.561.86 | 86.481.99 | |
| Informer Zhou et al. (2021) | 89.022.50 | 89.432.14 | 89.022.50 | 88.982.54 | 96.640.68 | 96.750.63 | |
| iTransformer Liu et al. (2024b) | 74.671.06 | 74.711.06 | 74.671.06 | 74.651.06 | 83.371.14 | 83.731.27 | |
| MTST Zhang et al. (2024) | 76.963.76 | 77.243.59 | 76.963.76 | 76.883.83 | 85.274.46 | 82.815.64 | |
| Nonformer Liu et al. (2022) | 87.882.48 | 88.861.84 | 87.882.48 | 87.782.56 | 97.050.68 | 96.990.68 | |
| PatchTST Nie et al. (2023) | 79.253.79 | 79.604.09 | 79.253.79 | 79.203.77 | 87.954.96 | 86.366.67 | |
| Reformer Kitaev et al. (2019) | 87.922.01 | 88.641.40 | 87.922.01 | 87.852.08 | 96.300.54 | 96.400.45 | |
| Transformer Vaswani et al. (2017) | 87.171.67 | 87.991.68 | 87.171.67 | 87.101.68 | 96.280.92 | 96.340.81 | |
| Medformer Wang et al. (2024) | 89.620.81 | 89.680.78 | 89.620.81 | 89.620.81 | 96.410.35 | 96.510.33 | |
| MedGNN Fan et al. (2025) | 91.040.09 | 91.150.12 | 91.040.20 | 91.040.08 | 96.740.04 | - | |
| HM-BiTCN + CIF | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 | |
| ADFTD (3-Classes) Reported | Autoformer Wu et al. (2021) | 45.251.48 | 43.671.94 | 42.962.03 | 42.591.85 | 61.021.82 | 43.102.30 |
| Crossformer Zhang and Yan (2022) | 50.452.31 | 45.571.63 | 45.881.82 | 45.501.70 | 66.452.03 | 48.332.05 | |
| FEDformer Zhou et al. (2022) | 46.300.59 | 46.050.76 | 44.221.38 | 43.911.37 | 62.621.75 | 46.111.44 | |
| Informer Zhou et al. (2021) | 48.451.96 | 46.541.68 | 46.061.84 | 45.741.38 | 65.871.27 | 47.601.30 | |
| iTransformer Liu et al. (2024b) | 52.601.59 | 46.791.27 | 47.281.29 | 46.791.13 | 67.261.16 | 49.531.21 | |
| MTST Zhang et al. (2024) | 45.602.03 | 44.701.33 | 45.051.30 | 44.311.74 | 62.500.81 | 45.160.85 | |
| Nonformer Liu et al. (2022) | 49.951.05 | 47.710.97 | 47.461.50 | 46.961.35 | 66.231.37 | 47.331.78 | |
| PatchTST Nie et al. (2023) | 44.370.95 | 42.401.13 | 42.061.48 | 41.971.37 | 60.081.50 | 42.491.79 | |
| Reformer Kitaev et al. (2019) | 50.781.17 | 49.641.49 | 49.891.67 | 47.940.69 | 69.171.58 | 51.731.94 | |
| Transformer Vaswani et al. (2017) | 50.472.14 | 49.131.83 | 48.011.53 | 48.091.59 | 67.931.59 | 48.932.02 | |
| Medformer Wang et al. (2024) | 53.271.54 | 51.021.57 | 50.711.55 | 50.651.51 | 70.931.19 | 51.211.32 | |
| MedGNN Fan et al. (2025) | 56.120.11 | 55.070.09 | 55.470.34 | 55.000.24 | 74.680.33 | - | |
| HM-BiTCN + CIF | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 | |
| PTB (2-Classes) Reported | Autoformer Wu et al. (2021) | 73.352.10 | 72.112.89 | 63.243.17 | 63.693.84 | 78.543.48 | 74.253.53 |
| Crossformer Zhang and Yan (2022) | 80.173.79 | 85.041.83 | 71.256.29 | 72.757.19 | 88.553.45 | 87.313.25 | |
| FEDformer Zhou et al. (2022) | 76.052.54 | 77.583.61 | 66.103.55 | 67.144.37 | 85.934.31 | 82.595.42 | |
| Informer Zhou et al. (2021) | 78.691.68 | 82.871.02 | 69.192.90 | 70.843.47 | 92.090.53 | 90.020.60 | |
| iTransformer Liu et al. (2024b) | 83.890.71 | 88.251.18 | 76.391.01 | 79.061.06 | 91.181.16 | 90.930.98 | |
| MTST Zhang et al. (2024) | 76.591.90 | 79.881.90 | 66.312.95 | 67.383.71 | 86.862.75 | 83.752.84 | |
| Nonformer Liu et al. (2022) | 78.660.49 | 82.770.86 | 69.120.87 | 70.901.00 | 89.372.51 | 86.672.38 | |
| PatchTST Nie et al. (2023) | 74.741.62 | 76.941.51 | 63.892.71 | 64.363.38 | 88.790.91 | 83.390.96 | |
| Reformer Kitaev et al. (2019) | 77.962.13 | 81.721.61 | 68.203.35 | 69.653.88 | 91.130.74 | 88.421.30 | |
| Transformer Vaswani et al. (2017) | 77.371.02 | 81.840.66 | 67.141.80 | 68.472.19 | 90.081.76 | 87.221.68 | |
| Medformer Wang et al. (2024) | 83.502.01 | 85.190.94 | 77.113.39 | 79.183.31 | 92.811.48 | 90.321.54 | |
| MedGNN Fan et al. (2025) | 84.530.28 | 87.350.45 | 77.900.66 | 80.400.62 | 93.310.46 | - | |
| HM-BiTCN + CIF | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 | |
| PTB-XL (5-Classes) Reported | Autoformer Wu et al. (2021) | 61.682.72 | 51.601.64 | 49.101.52 | 48.852.27 | 82.041.44 | 51.931.71 |
| Crossformer Zhang and Yan (2022) | 73.300.14 | 65.060.35 | 61.230.33 | 62.590.14 | 90.020.06 | 67.430.22 | |
| FEDformer Zhou et al. (2022) | 57.209.47 | 52.386.09 | 49.047.26 | 47.898.44 | 82.134.17 | 52.317.03 | |
| Informer Zhou et al. (2021) | 71.430.32 | 62.640.60 | 59.120.47 | 60.440.43 | 88.650.09 | 64.760.17 | |
| iTransformer Liu et al. (2024b) | 69.280.22 | 59.590.45 | 54.620.18 | 56.200.19 | 86.710.10 | 60.270.21 | |
| MTST Zhang et al. (2024) | 72.140.27 | 63.840.72 | 60.010.81 | 61.430.38 | 88.970.33 | 65.830.51 | |
| Nonformer Liu et al. (2022) | 70.560.55 | 61.570.66 | 57.750.72 | 59.100.66 | 88.320.36 | 63.400.79 | |
| PatchTST Nie et al. (2023) | 73.230.25 | 65.700.64 | 60.820.76 | 62.610.34 | 89.740.19 | 67.320.22 | |
| Reformer Kitaev et al. (2019) | 71.720.43 | 63.121.02 | 59.200.75 | 60.690.18 | 88.800.24 | 64.720.47 | |
| Transformer Vaswani et al. (2017) | 70.590.44 | 61.570.65 | 57.620.35 | 59.050.25 | 88.210.16 | 63.360.29 | |
| Medformer Wang et al. (2024) | 72.870.23 | 64.140.42 | 60.600.46 | 62.020.37 | 89.660.13 | 66.390.22 | |
| MedGNN Fan et al. (2025) | 73.870.18 | 66.260.29 | 61.130.29 | 62.540.20 | 90.210.15 | - | |
| HM-BiTCN + CIF | 73.730.30 | 65.410.67 | 60.701.08 | 61.890.91 | 90.530.22 | 67.750.75 |
5.2.1 Efficiency Analysis
We evaluate the model efficiency in terms of accuracy, training speed, and memory footprint using two datasets: APAVA and TDBRAIN. In Figure 3, a marker closer to the upper-left corner indicates higher accuracy and faster training speed, while a smaller marker area corresponds to lower memory usage. The results show that HM-BiTCN achieves the best overall performance among all baseline methods, demonstrating its high efficiency and reliability across different application scenarios.
5.3 Ablation Study
(1) Effectiveness of CIF: Table 3 demonstrates the excellent performance of combining HM-BiTCN with CIF, confirming the compatibility of the HM-BiTCN with CIF. See Appendix F.1 for details.
| Datasets | APAVA | ADFTD | PTB | TDBRAIN | ||||
| Metrics | Accuracy | F1 Score | Accuracy | F1 Score | Accuracy | F1 Score | Accuracy | F1 Score |
| w/ CIF | 85.161.55 | 84.821.49 | 58.560.93 | 55.420.82 | 88.291.45 | 85.591.96 | 93.131.41 | 93.121.42 |
| w/o CIF | 82.491.40 | 81.601.39 | 52.052.22 | 49.482.70 | 81.871.87 | 75.843.20 | 84.902.60 | 84.762.74 |
| Improvement | +2.67% | +3.22% | +6.51% | +5.94% | +6.42% | +9.75% | +8.23% | +8.36% |
(2) Hyperparameter Transfer and Adaptation: We evaluate the transferability of key hyperparameters (e.g., , , ) from HM-BiTCN to other models. If transferred settings underperform, we further fine-tune them for adaptation. See Appendices F.2 and F.3. Figure 4 illustrates the outstanding performance of CIF when combined with other models.
6 Conclusion
In this paper, we propose a simple yet effective temporal classification representation method called Channel-Imposed Fusion (CIF) and combine it with a straightforward TCN structure. This approach outperforms current SOTA methods on multiple medical time-series datasets. Additionally, the CIF method demonstrates strong adaptability, enabling integration with cutting-edge methods to further enhance their performance. Through the CIF method, we aim to show researchers in the field that, rather than solely modifying models to process high-level features, exploring the underlying core features of the data itself is also a research path worth delving into.
Appendix A Proof of SNR Optimization via CIF.
Consider two observed signals:
where:
are zero-mean useful signals with variance and correlation coefficient
are zero-mean noise components with variance and correlation coefficient
Signal and noise components are mutually uncorrelated
Derivation of Output SNR
For the linear combination :
1. Signal Power Calculation
2. Noise Power Calculation
3. SNR Ratio
The output SNR becomes:
where .
Optimization Conditions
SNR improvement occurs when:
Difference Mode ():
Cooperative Mode ():
Appendix B Visualization of the Channel-Imposed Fusion (CIF) Process.
In Figure 5 6, we visualize scenarios with highly correlated channel information and those with generally correlated channel information. These visualizations intuitively demonstrate the strong representation capability of our proposed method.
Appendix C Data Preprocessing
| Datasets | #-Subject | #-Sample | #-Class | #-Channel | #-Timestamps | Sampling Rate | Modality | File Size |
| APAVA | 23 | 5,967 | 2 | 16 | 256 | 256Hz | EEG | 186MB |
| ADFTD | 88 | 69,752 | 3 | 19 | 256 | 256Hz | EEG | 2.52GB |
| TDBrain | 72 | 6,240 | 2 | 33 | 256 | 256Hz | EEG | 571MB |
| PTB | 198 | 64,356 | 2 | 15 | 300 | 250Hz | ECG | 2.15GB |
| PTB-XL | 17,596 | 191,400 | 5 | 12 | 250 | 250Hz | ECG | 4.28GB |
Appendix D Implementation Details
We implement our method and all the baselines based on the Time-Series-Library project111https://github.com/thuml/Time-Series-Library from Tsinghua University Wu et al. [2023], which integrates all methods under the same framework and training techniques to ensure a relatively fair comparison. The 12 baseline time series transformer methods are Autoformer Wu et al. [2021], Crossformer Zhang and Yan [2022], FEDformer Zhou et al. [2022], Informer Zhou et al. [2021], iTransformer Liu et al. [2024b], MTST Zhang et al. [2024], Nonformer Liu et al. [2022], PatchTST Nie et al. [2023], Reformer Kitaev et al. [2019], vanilla Transformer Vaswani et al. [2017], Medformer Wang et al. [2024], and MedGNN Fan et al. [2025].
Autoformer Autoformer Wu et al. [2021] employs an auto-correlation mechanism to replace self-attention for time series forecasting. Additionally, they use a time series decomposition block to separate the time series into trend-cyclical and seasonal components for improved learning. The raw source code is available at https://github.com/thuml/Autoformer.
Crossformer Crossformer Zhang and Yan [2022] designs a single-channel patching approach for token embedding. They utilize two-stage self-attention to leverage both temporal features and channel correlations. A router mechanism is proposed to reduce time and space complexity during the cross-dimension stage. The raw code is available at https://github.com/Thinklab-SJTU/Crossformer.
FEDformer FEDformer Zhou et al. [2022] leverages frequency domain information using the Fourier transform. They introduce frequency-enhanced blocks and frequency-enhanced attention, which are computed in the frequency domain. A novel time series decomposition method replaces the layer norm module in the transformer architecture to improve learning. The raw code is available at https://github.com/MAZiqing/FEDformer.
Informer Informer Zhou et al. [2021] is the first paper to employ a one-forward procedure instead of an autoregressive method in time series forecasting tasks. They introduce ProbSparse self-attention to reduce complexity and memory usage. The raw code is available at https://github.com/zhouhaoyi/Informer2020.
iTransformer iTransformer Liu et al. [2024b] questions the conventional approach of embedding attention tokens in time series forecasting tasks and proposes an inverted approach by embedding the whole series of channels into a token. They also invert the dimension of other transformer modules, such as the layer norm and feed-forward networks. The raw code is available at https://github.com/thuml/iTransformer.
MTST MTST Zhang et al. [2024] uses the same token embedding method as Crossformer and PatchTST. It highlights the importance of different patching lengths in forecasting tasks and designs a method that can take different sizes of patch tokens as input simultaneously. The raw code is available at https://github.com/networkslab/MTST.
Nonformer Nonformer Liu et al. [2022] analyzes the impact of non-stationarity in time series forecasting tasks and its significant effect on results. They design a de-stationary attention module and incorporate normalization and denormalization steps before and after training to alleviate the over-stationarization problem. The raw code is available at https://github.com/thuml/Nonstationary_Transformers.
PatchTST PatchTST Nie et al. [2023] embeds a sequence of single-channel timestamps as a patch token to replace the attention token used in the vanilla transformer. This approach enlarges the receptive field and enhances forecasting ability. The raw code is available at https://github.com/yuqinie98/PatchTST.
Reformer Reformer Kitaev et al. [2019] replaces dot-product attention with locality-sensitive hashing. They also use a reversible residual layer instead of standard residuals. The raw code is available at https://github.com/lucidrains/reformer-pytorch.
Transformer Transformer Vaswani et al. [2017], commonly known as the vanilla transformer, is introduced in the well-known paper "Attention is All You Need." It can also be applied to time series by embedding each timestamp of all channels as an attention token. The PyTorch version of the code is available at https://github.com/jadore801120/attention-is-all-you-need-pytorch.
Medformer Medformer Wang et al. [2024] uses cross-channel patch embedding to model spatiotemporal dependencies.The raw code is available at https://github.com/DL4mHealth/Medformer
MedGNN MedGNN Fan et al. [2025] employs multi-resolution spatiotemporal graph learning to extract dynamic features across multiple time scales. The raw code is available at https://github.com/aikunyi/MedGNN.
D.1 Evaluation metrics
For all methods, the optimizer used is Adam, with a learning rate of 1e-4. The batch size is set to {32,32,128,128,128} for the datasets APAVA, TDBrain, ADFD, PTB, and PTB-XL, respectively. Training is conducted for 100 epochs, with early stopping triggered after 10 epochs without improvement in the F1 score on the validation set. We save the model with the best F1 score on the validation set and evaluate it on the test set. We employ six evaluation metrics: accuracy, precision (macro-averaged), recall (macro-averaged), F1 score (macro-averaged), AUROC (macro-averaged), and AUPRC (macro-averaged). Both subject-dependent and subject-independent setups are implemented for different datasets. Each experiment is run with 5 random seeds (41-45) and fixed training, validation, and test sets to compute the average results and standard deviations.
To comprehensively and fairly evaluate the performance of each model in the classification task, we select five evaluation metrics: Accuracy, Precision, Recall, F1 score, and AUROC. The definitions and specific calculation formulas for each metric are presented below:
Accuracy measures the proportion of correct predictions out of the total number of predictions. It’s calculated as:
This metric is useful when the classes are balanced but may be misleading in cases of class imbalance.
Precision focuses on the quality of positive predictions and measures the proportion of correctly predicted positive instances out of all instances predicted as positive. It’s especially useful when false positives need to be minimized. The formula is:
Recall measures the proportion of actual positive instances that were correctly identified. It’s important when false negatives are costly. The formula is:
It shows how well the model captures all relevant instances.
The F1 score is the harmonic mean of precision and recall, balancing the two when one is more important than the other. It’s particularly useful when dealing with imbalanced datasets, as it accounts for both false positives and false negatives. The formula is:
It gives a single metric that reflects both precision and recall performance.
Appendix E Reproduced Results of Existing Methods
We employ six evaluation metrics: accuracy, precision (macro-averaged), recall (macro-averaged), F1 score (macro-averaged), AUROC (macro-averaged), and AUPRC (macro-averaged). The training process is conducted with five random seeds (41-45) on fixed training, validation, and test sets to compute the mean and standard deviation of the models. All experiments are run on an NVIDIA RTX 3090 GPU.
E.1 Results of Subject-Dependent
In Table 5, we present the results of various models under the Subject-Dependent Setup on the ADFTD (3-Classes) dataset. The results indicate that our proposed CIF method significantly improves the performance of models, particularly in accuracy, F1 score, AUROC, and AUPRC metrics. Specifically, the combination of MedGNN + CIF outperforms MedGNN alone across all key performance metrics, and similarly, Medformer + CIF surpasses Medformer in every critical evaluation metric. These results convincingly demonstrate the effectiveness of the CIF method in enhancing model performance. By integrating the CIF method, not only are the models’ performances significantly improved across multiple evaluation metrics, but the broad applicability and substantial benefits of the method across different model architectures are also highlighted. Notably, MedGNN + CIF achieves a remarkable Accuracy of 99.60%, far surpassing other models, highlighting the advantage of our approach in handling imbalanced datasets. Furthermore, MedGNN + CIF shows significant improvements in F1 score and precision, indicating that the CIF method effectively enhances the model’s ability to recognize both positive and negative samples. Thus, our experimental results validate the effectiveness of the CIF method in boosting model generalization and precision.
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| ADFTD (3-Classes) Reproduced | Autoformer Wu et al. [2021] | 87.792.87 | 87.632.95 | 86.843.10 | 87.173.05 | 96.451.15 | 93.522.05 |
| Crossformer Zhang and Yan [2022] | 89.271.19 | 88.981.27 | 88.611.35 | 88.771.29 | 97.530.48 | 95.470.87 | |
| FEDformer Zhou et al. [2022] | 77.552.42 | 76.712.14 | 76.462.52 | 76.442.52 | 91.661.36 | 84.992.15 | |
| Informer Zhou et al. [2021] | 91.720.13 | 91.500.16 | 91.360.21 | 91.420.14 | 98.400.07 | 96.920.11 | |
| iTransformer Liu et al. [2024b] | 64.600.77 | 62.220.64 | 61.950.35 | 61.870.41 | 81.460.29 | 68.810.39 | |
| MTST Zhang et al. [2024] | 65.430.79 | 64.470.65 | 63.490.55 | 63.710.55 | 81.670.51 | 69.740.69 | |
| Nonformer Liu et al. [2022] | 96.000.72 | 95.720.80 | 95.900.69 | 95.810.75 | 99.560.13 | 99.020.23 | |
| PatchTST Nie et al. [2023] | 66.550.48 | 65.610.60 | 65.260.78 | 65.260.60 | 83.270.33 | 71.960.44 | |
| Reformer Kitaev et al. [2019] | 90.742.18 | 90.872.15 | 90.612.01 | 90.442.25 | 98.690.50 | 97.570.85 | |
| Transformer Vaswani et al. [2017] | 96.620.54 | 96.460.54 | 96.450.59 | 96.450.56 | 99.690.08 | 99.310.13 | |
| Medformer Wang et al. [2024] | 95.761.03 | 95.541.10 | 95.621.05 | 95.581.08 | 99.530.18 | 99.090.36 | |
| Medformer + CIF | 98.870.26 | 98.770.27 | 98.860.27 | 98.810.27 | 99.960.01 | 99.920.03 | |
| MedGNN Fan et al. [2025] | 99.440.21 | 99.430.21 | 99.410.22 | 99.420.21 | 99.980.01 | 99.970.01 | |
| MedGNN + CIF | 99.600.09 | 99.600.11 | 99.580.09 | 99.590.10 | 99.990.01 | 99.970.01 |
E.2 Results of Subject-Independent
Table 2 presents the results of various methods reported under the subject-independent setup, while Table 6 shows the results of our reproduction of these methods. Our approach achieves the highest scores on six metrics across four out of the five datasets. For the PTB-XL dataset, compared to the results reported by other methods, our method achieves the highest AUROC, with Accuracy ranking second. In comparison with our reproduced results, our method ranks first in Accuracy, AUROC, and AUPRC, and second in Precision and Recall. Additionally, it is noteworthy that under the subject-independent setup, the F1 score for ADFTD is 55.42%, significantly lower than the 99.59% F1 score achieved under the subject-dependent setup. This comparison highlights the challenge of the subject-independent setup, which better simulates real-world scenarios.
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| APAVA (2-Classes) Reproduced | Autoformer Wu et al. [2021] | 73.187.33 | 73.876.72 | 73.016.10 | 72.407.03 | 81.647.24 | 81.107.75 |
| Crossformer Zhang and Yan [2022] | 72.762.04 | 79.642.45 | 67.412.62 | 66.883.61 | 71.814.06 | 71.643.74 | |
| FEDformer Zhou et al. [2022] | 75.161.67 | 74.980.69 | 73.342.97 | 73.502.90 | 83.891.54 | 83.271.62 | |
| Informer Zhou et al. [2021] | 72.202.78 | 73.924.80 | 68.482.51 | 68.742.70 | 70.143.43 | 70.843.80 | |
| iTransformer Liu et al. [2024b] | 74.551.66 | 74.772.10 | 71.761.72 | 72.301.79 | 85.591.55 | 84.391.57 | |
| MTST Zhang et al. [2024] | 69.241.24 | 75.872.80 | 63.281.81 | 61.622.75 | 66.093.27 | 68.082.93 | |
| Nonformer Liu et al. [2022] | 71.814.20 | 71.314.40 | 70.153.38 | 70.383.74 | 71.542.73 | 72.792.50 | |
| PatchTST Nie et al. [2023] | 68.272.11 | 78.561.88 | 61.532.60 | 58.524.07 | 64.612.18 | 67.142.06 | |
| Reformer Kitaev et al. [2019] | 78.422.85 | 80.894.52 | 75.202.28 | 76.092.54 | 75.482.79 | 77.522.64 | |
| Transformer Vaswani et al. [2017] | 75.534.28 | 76.905.05 | 72.144.87 | 72.645.44 | 72.306.04 | 73.047.15 | |
| Medformer Wang et al. [2024] | 77.852.42 | 80.313.21 | 74.382.49 | 75.212.67 | 80.853.80 | 81.623.24 | |
| MedGNN Fan et al. [2025] | 77.405.77 | 82.774.46 | 73.247.06 | 73.299.01 | 81.312.94 | 82.802.91 | |
| HM-BiTCN + CIF | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | |
| TDBrain (2-Classes) Reproduced | Autoformer Wu et al. [2021] | 90.383.03 | 91.162.42 | 90.383.03 | 90.313.09 | 95.832.14 | 95.432.31 |
| Crossformer Zhang and Yan [2022] | 82.152.60 | 82.812.11 | 82.152.60 | 82.042.70 | 91.202.23 | 91.472.16 | |
| FEDformer Zhou et al. [2022] | 77.601.23 | 78.251.52 | 77.601.23 | 77.481.19 | 86.311.23 | 86.481.36 | |
| Informer Zhou et al. [2021] | 88.422.99 | 89.012.45 | 88.422.99 | 88.363.05 | 96.540.90 | 96.660.85 | |
| iTransformer Liu et al. [2024b] | 74.691.02 | 74.761.04 | 74.691.02 | 74.671.02 | 83.351.24 | 83.651.41 | |
| MTST Zhang et al. [2024] | 77.673.58 | 78.974.37 | 77.673.58 | 77.453.55 | 86.474.84 | 84.996.43 | |
| Nonformer Liu et al. [2022] | 88.102.39 | 88.761.74 | 88.102.39 | 88.042.47 | 96.560.91 | 96.361.21 | |
| PatchTST Nie et al. [2023] | 77.982.64 | 79.303.73 | 77.982.64 | 77.762.65 | 86.674.03 | 84.935.47 | |
| Reformer Kitaev et al. [2019] | 88.502.30 | 89.011.80 | 88.502.30 | 88.452.35 | 96.100.63 | 96.190.55 | |
| Transformer Vaswani et al. [2017] | 85.131.86 | 86.391.56 | 85.131.86 | 84.991.93 | 95.611.05 | 95.630.91 | |
| Medformer Wang et al. [2024] | 88.771.24 | 88.911.11 | 88.771.24 | 88.761.25 | 96.380.34 | 96.440.30 | |
| MedGNN Fan et al. [2025] | - | - | - | - | - | - | |
| HM-BiTCN + CIF | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 | |
| ADFTD (3-Classes) Reproduced | Autoformer Wu et al. [2021] | 46.902.89 | 45.592.37 | 44.912.23 | 44.342.52 | 63.492.44 | 45.632.29 |
| Crossformer Zhang and Yan [2022] | 50.181.97 | 45.971.84 | 46.301.73 | 45.901.84 | 66.681.67 | 48.651.89 | |
| FEDformer Zhou et al. [2022] | 45.750.78 | 45.711.29 | 44.271.28 | 43.511.00 | 62.641.64 | 45.881.35 | |
| Informer Zhou et al. [2021] | 48.421.99 | 46.941.60 | 46.410.99 | 45.760.43 | 65.991.14 | 47.491.07 | |
| iTransformer Liu et al. [2024b] | 52.851.36 | 46.971.05 | 47.311.03 | 46.840.78 | 67.460.96 | 49.900.89 | |
| MTST Zhang et al. [2024] | 45.771.70 | 44.391.73 | 43.701.82 | 43.361.98 | 61.381.57 | 44.011.60 | |
| Nonformer Liu et al. [2022] | 50.811.06 | 48.711.40 | 48.551.47 | 48.361.38 | 66.951.54 | 48.081.82 | |
| PatchTST Nie et al. [2023] | 43.320.53 | 41.950.38 | 41.451.26 | 40.751.62 | 60.210.29 | 42.490.57 | |
| Reformer Kitaev et al. [2019] | 51.282.60 | 49.682.75 | 49.642.02 | 48.452.06 | 69.202.53 | 51.743.24 | |
| Transformer Vaswani et al. [2017] | 50.530.94 | 49.310.87 | 48.571.23 | 48.421.28 | 67.980.90 | 49.071.35 | |
| Medformer Wang et al. [2024] | 53.701.18 | 51.511.32 | 50.491.48 | 50.351.53 | 70.481.17 | 50.911.13 | |
| MedGNN Fan et al. [2025] | 50.223.21 | 48.653.72 | 47.504.57 | 47.334.40 | 67.184.39 | 48.844.11 | |
| HM-BiTCN + CIF | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 | |
| PTB (2-Classes) Reproduced | Autoformer Wu et al. [2021] | 71.992.74 | 69.603.85 | 61.504.23 | 61.435.07 | 74.291.89 | 70.262.00 |
| Crossformer Zhang and Yan [2022] | 78.063.44 | 81.533.13 | 68.625.63 | 69.766.53 | 88.312.07 | 85.812.43 | |
| FEDformer Zhou et al. [2022] | 74.542.27 | 77.994.10 | 63.143.29 | 63.284.36 | 84.634.27 | 80.915.55 | |
| Informer Zhou et al. [2021] | 79.590.65 | 83.330.77 | 70.580.95 | 72.581.08 | 92.770.48 | 90.890.57 | |
| iTransformer Liu et al. [2024b] | 83.431.19 | 88.061.47 | 75.641.55 | 78.291.70 | 91.381.41 | 91.081.30 | |
| MTST Zhang et al. [2024] | 75.532.45 | 78.721.87 | 64.784.06 | 65.304.81 | 87.764.09 | 83.603.92 | |
| Nonformer Liu et al. [2022] | 78.931.46 | 82.481.53 | 69.682.15 | 71.502.56 | 90.540.59 | 87.781.46 | |
| PatchTST Nie et al. [2023] | 75.282.44 | 77.052.44 | 64.864.05 | 65.415.28 | 88.112.59 | 82.652.87 | |
| Reformer Kitaev et al. [2019] | 78.111.65 | 82.700.80 | 68.172.68 | 69.683.34 | 90.771.56 | 88.141.20 | |
| Transformer Vaswani et al. [2017] | 76.431.98 | 81.251.15 | 65.643.29 | 66.444.39 | 90.211.24 | 87.281.49 | |
| Medformer Wang et al. [2024] | 80.990.75 | 83.010.72 | 73.351.16 | 75.471.21 | 93.101.18 | 90.691.04 | |
| MedGNN Fan et al. [2025] | - | - | - | - | - | - | |
| HM-BiTCN + CIF | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 | |
| PTB-XL (5-Classes) Reproduced | Autoformer Wu et al. [2021] | 60.371.72 | 49.830.50 | 49.170.72 | 48.430.47 | 81.350.56 | 50.830.72 |
| Crossformer Zhang and Yan [2022] | 73.150.17 | 64.620.31 | 61.280.45 | 62.490.34 | 89.940.15 | 67.140.29 | |
| FEDformer Zhou et al. [2022] | 56.0110.27 | 51.615.58 | 48.777.64 | 47.238.76 | 82.014.20 | 52.237.14 | |
| Informer Zhou et al. [2021] | 71.280.27 | 62.280.53 | 59.180.54 | 60.380.38 | 88.570.09 | 64.570.18 | |
| iTransformer Liu et al. [2024b] | 69.180.33 | 59.440.48 | 54.840.25 | 56.440.23 | 86.650.17 | 60.270.44 | |
| MTST Zhang et al. [2024] | 72.140.29 | 64.010.61 | 59.310.27 | 61.070.32 | 88.820.19 | 65.520.52 | |
| Nonformer Liu et al. [2022] | 70.510.79 | 61.421.03 | 58.150.57 | 59.310.63 | 88.200.43 | 63.330.77 | |
| PatchTST Nie et al. [2023] | 73.130.20 | 65.380.54 | 60.610.56 | 62.390.27 | 89.660.16 | 66.970.24 | |
| Reformer Kitaev et al. [2019] | 71.240.46 | 61.780.79 | 59.670.99 | 60.510.65 | 88.670.21 | 64.290.29 | |
| Transformer Vaswani et al. [2017] | 70.460.42 | 61.560.58 | 57.860.65 | 59.160.46 | 88.180.14 | 63.250.27 | |
| Medformer Wang et al. [2024] | 72.940.15 | 64.390.32 | 59.980.49 | 61.610.33 | 89.670.10 | 66.250.25 | |
| MedGNN Fan et al. [2025] | - | - | - | - | - | - | |
| HM-BiTCN + CIF | 73.730.30 | 65.410.67 | 60.701.08 | 61.890.91 | 90.530.22 | 67.750.75 |
Appendix F Ablation Study
F.1 Module Study of CIF
We conduct experiments by integrating our Channel-Imposed Fusion principle with other methods.
In the CIF structure, there are four different hyperparameters. First, is the switch for forward and backward feature fusion. When , it replaces the forward features with the fused features; when , it replaces the backward features with the fused features. The second hyperparameter, , determines the number of selected channels. Additionally, there are two scaling factors associated with the channels: , the scaling factor for the first channels, and , the scaling factor for the remaining channels. Both and can either be learnable or fixed parameters.
We use
to represent learnable parameters, and
to represent non-learnable parameters.
The ablation study(Table 7 8 9) reveals that the performance of the CIF architecture is highly sensitive to its hyperparameters, with optimal configurations varying significantly across datasets. For the APAVA dataset (16 channels), the best results (85.16% accuracy, 94.06% AUROC) are achieved using backward fusion () on 9 selected channels (), where the first 7 channels apply fixed reinforcement weights (, non-learnable) and the rest use fixed suppression weights (, non-learnable). Conversely, the ADFTD dataset (19 channels) performs best (58.56% accuracy, 76.07% AUROC) with forward fusion () over 10 selected channels () using dynamically learned weights for negative suppression () and positive enhancement (). These differences, also observed in PTB (APAVA: 78–85% accuracy vs. ADFTD: 50–58%), emphasize the critical role of dataset-specific tuning. In particular, the number of selected channels () and fusion direction () are pivotal, while adaptive weighting () further improves robustness. Overall, the results highlight the necessity of careful hyperparameter optimization—especially in channel selection and fusion strategy—for maximizing CIF performance across diverse medical time series tasks.
| t | n | a | b | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| - | - | - | - | 52.052.22 | 50.453.00 | 50.402.55 | 49.482.70 | 69.432.84 | 50.993.15 |
| 1 | 10 | -0.20 |
-0.30 |
57.330.95 | 54.061.47 | 54.441.33 | 53.731.47 | 74.821.36 | 58.241.78 |
| 1 | 10 | -0.25 |
-0.40 |
57.081.46 | 53.310.98 | 53.241.14 | 52.870.98 | 73.771.28 | 56.861.47 |
| 1 | 11 | -0.20 |
-0.30 |
56.681.30 | 53.042.00 | 53.452.03 | 52.842.30 | 73.811.55 | 57.341.46 |
| 1 | 10 | -0.19 |
-0.27 |
58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 |
| 1 | 7 | 0.20 |
-0.60 |
56.360.58 | 53.911.10 | 54.341.03 | 53.780.91 | 74.190.69 | 56.621.28 |
| 1 | 9 | -0.20 |
-0.60 |
55.201.50 | 53.051.14 | 53.211.06 | 52.761.02 | 73.091.13 | 55.501.71 |
| 1 | 10 | -0.20 |
-0.20 |
57.110.92 | 55.041.43 | 55.531.57 | 54.581.29 | 75.131.33 | 58.471.43 |
| 1 | 11 | -0.20 |
-0.30 |
57.831.16 | 54.321.22 | 54.511.18 | 54.011.21 | 74.741.12 | 57.961.08 |
| 1 | 11 | -0.18 |
-0.25 |
57.001.32 | 53.002.28 | 53.171.94 | 52.322.19 | 73.971.89 | 57.001.93 |
| t | n | a | b | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| - | - | - | - | 81.871.87 | 86.501.24 | 73.492.90 | 75.843.20 | 94.200.29 | 93.040.45 |
| 1 | 8 | 0.21 |
-0.50 |
88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 |
| 1 | 8 | 0.20 |
-0.50 |
84.452.15 | 88.341.15 | 77.403.52 | 79.893.51 | 93.150.70 | 91.940.67 |
| 1 | 9 | 0.22 |
-0.50 |
85.140.22 | 87.421.08 | 79.050.52 | 81.420.35 | 93.821.19 | 92.731.27 |
| 1 | 10 | 0.22 |
-0.50 |
85.402.08 | 88.171.96 | 79.173.12 | 81.573.07 | 94.571.13 | 93.531.44 |
| 1 | 11 | 0.22 |
-0.50 |
84.470.84 | 88.020.59 | 77.501.26 | 80.101.28 | 92.770.89 | 92.020.78 |
| 1 | 8 | 0.23 |
-0.50 |
83.965.41 | 87.133.96 | 76.898.22 | 78.669.51 | 92.492.91 | 91.143.79 |
| 1 | 8 | 0.25 |
-0.50 |
85.562.09 | 88.961.77 | 79.033.03 | 81.613.14 | 93.611.11 | 93.071.03 |
| t | n | a | b | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| - | - | - | - | 82.491.40 | 82.381.79 | 81.201.32 | 81.601.39 | 91.101.63 | 91.301.71 |
| -1 | 5 | 1 |
1 |
81.481.96 | 81.992.38 | 79.522.00 | 80.212.07 | 90.931.74 | 91.171.77 |
| -1 | 6 | 1 |
1 |
84.002.37 | 84.332.61 | 82.372.44 | 83.022.50 | 92.672.57 | 92.872.55 |
| -1 | 7 | 1 |
1 |
81.772.31 | 82.673.24 | 79.782.04 | 80.492.26 | 89.922.60 | 90.232.53 |
| 1 | 9 | 1 |
1 |
85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 |
| -1 | 10 | 1 |
1 |
82.601.82 | 82.302.04 | 81.922.29 | 81.921.95 | 91.941.55 | 92.161.50 |
| -1 | 11 | 1 |
2 |
78.980.97 | 78.571.09 | 77.600.93 | 77.930.97 | 87.290.93 | 87.350.96 |
F.2 Generalizability of Hyperparameters
We transfer the key hyperparameters obtained from HM-BiTCN—such as the fusion coefficients (, ), and segment length ()—to other baseline models. This experiment evaluates whether these hyperparameters retain effectiveness when applied to different architectures, indicating their robustness and broader applicability.
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| APAVA (2-Classes) Reproduced | Autoformer | 73.187.33 | 73.876.72 | 73.016.10 | 72.407.03 | 81.647.24 | 81.107.75 |
| Autoformer + CIF | 75.962.68 | 76.333.47 | 75.111.10 | 74.971.90 | 83.421.99 | 83.152.63 | |
| Crossformer | 72.762.04 | 79.642.45 | 67.412.62 | 66.883.61 | 71.814.06 | 71.643.74 | |
| Crossformer + CIF | 82.322.60 | 85.351.83 | 79.213.17 | 80.293.30 | 90.391.58 | 90.021.47 | |
| FEDformer | 75.161.67 | 74.980.69 | 73.342.97 | 73.502.90 | 83.891.54 | 83.271.62 | |
| FEDformer + CIF | 77.202.17 | 76.972.07 | 77.022.60 | 76.552.40 | 86.701.73 | 86.531.76 | |
| Informer | 72.202.78 | 73.924.80 | 68.482.51 | 68.742.70 | 70.143.43 | 70.843.80 | |
| Informer + CIF | 79.782.07 | 82.293.03 | 76.552.05 | 77.532.23 | 78.561.33 | 78.581.14 | |
| iTransformer | 74.551.66 | 74.772.10 | 71.761.72 | 72.301.79 | 85.591.55 | 84.391.57 | |
| iTransformer + CIF | 74.950.87 | 74.400.75 | 73.791.78 | 73.811.43 | 84.300.88 | 82.490.96 | |
| MTST | 69.241.24 | 75.872.80 | 63.281.81 | 61.622.75 | 66.093.27 | 68.082.93 | |
| MTST + CIF | 76.202.39 | 81.461.11 | 71.673.06 | 72.063.73 | 77.653.37 | 77.982.73 | |
| Nonformer | 71.814.20 | 71.314.40 | 70.153.38 | 70.383.74 | 71.542.73 | 72.792.50 | |
| Nonformer + CIF | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |
| Reformer | 78.422.85 | 80.894.52 | 75.202.28 | 76.092.54 | 75.482.79 | 77.522.64 | |
| Reformer + CIF | 81.510.57 | 84.900.86 | 78.180.60 | 79.320.64 | 79.102.50 | 80.772.21 | |
| Transformer | 75.534.28 | 76.905.05 | 72.144.87 | 72.645.44 | 72.306.04 | 73.047.15 | |
| Transformer + CIF | 77.962.82 | 79.343.86 | 75.072.52 | 75.872.73 | 74.751.62 | 74.762.33 | |
| Medformer | 77.852.42 | 80.313.21 | 74.382.49 | 75.212.67 | 80.853.80 | 81.623.24 | |
| Medformer + CIF | 81.061.58 | 82.972.23 | 78.261.52 | 79.231.65 | 85.741.85 | 86.321.48 | |
| HM-BiTCN | 82.491.40 | 82.381.79 | 81.201.32 | 81.601.39 | 91.101.63 | 91.301.71 | |
| HM-BiTCN + CIF | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 |
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| TDBrain (2-Classes) Reproduced | Autoformer | 90.383.03 | 91.162.42 | 90.383.03 | 90.313.09 | 95.832.14 | 95.432.31 |
| Autoformer + CIF | 93.422.49 | 93.712.27 | 93.422.49 | 93.402.51 | 97.461.08 | 97.211.15 | |
| Crossformer | 82.152.60 | 82.812.11 | 82.152.60 | 82.042.70 | 91.202.23 | 91.472.16 | |
| Crossformer + CIF | 89.401.26 | 89.831.18 | 89.401.26 | 89.371.27 | 96.760.62 | 96.800.62 | |
| FEDformer | 77.601.23 | 78.251.52 | 77.601.23 | 77.481.19 | 86.311.23 | 86.481.36 | |
| FEDformer + CIF | 78.871.94 | 79.192.00 | 78.881.94 | 78.821.94 | 88.121.87 | 88.391.92 | |
| Informer | 88.422.99 | 89.012.45 | 88.422.99 | 88.363.05 | 96.540.90 | 96.660.85 | |
| Informer + CIF | 89.381.66 | 89.711.37 | 89.381.66 | 89.351.68 | 96.920.66 | 97.030.64 | |
| iTransformer | 74.691.02 | 74.761.04 | 74.691.02 | 74.671.02 | 83.351.24 | 83.651.41 | |
| iTransformer + CIF | 72.791.12 | 72.891.16 | 72.791.12 | 72.761.11 | 81.081.03 | 81.310.93 | |
| MTST | 77.673.58 | 78.974.37 | 77.673.58 | 77.453.55 | 86.474.84 | 84.996.43 | |
| MTST + CIF | 87.211.99 | 87.302.06 | 87.211.99 | 87.201.99 | 94.141.83 | 93.242.63 | |
| Nonformer | 88.102.39 | 88.761.74 | 88.102.39 | 88.042.47 | 96.560.91 | 96.361.21 | |
| Nonformer + CIF | 88.002.06 | 88.911.68 | 88.002.06 | 87.922.12 | 97.101.10 | 97.171.09 | |
| PatchTST | 77.982.64 | 79.303.73 | 77.982.64 | 77.762.65 | 86.674.03 | 84.935.47 | |
| PatchTST + CIF | 79.580.86 | 80.220.82 | 79.580.86 | 79.470.88 | 87.201.57 | 85.811.17 | |
| Reformer | 88.502.30 | 89.011.80 | 88.502.30 | 88.452.35 | 96.100.63 | 96.190.55 | |
| Reformer + CIF | 89.081.19 | 89.470.70 | 89.081.19 | 89.051.23 | 96.880.35 | 97.000.35 | |
| Transformer | 85.131.86 | 86.391.56 | 85.131.86 | 84.991.93 | 95.611.05 | 95.630.91 | |
| Transformer + CIF | 89.961.57 | 90.531.26 | 89.961.57 | 89.921.60 | 97.730.45 | 97.770.46 | |
| Medformer | 88.771.24 | 88.911.11 | 88.771.24 | 88.761.25 | 96.380.34 | 96.440.30 | |
| Medformer + CIF | 90.880.87 | 90.940.81 | 90.880.87 | 90.870.87 | 97.390.30 | 97.460.28 | |
| HM-BiTCN | 84.902.60 | 86.022.00 | 84.902.60 | 84.762.74 | 93.941.92 | 94.201.85 | |
| HM-BiTCN + CIF | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 |
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| ADFTD (3-Classes) Reproduced | Autoformer | 46.902.89 | 45.592.37 | 44.912.23 | 44.342.52 | 63.492.44 | 45.632.29 |
| Autoformer + CIF | 45.922.78 | 44.812.44 | 44.662.25 | 44.262.38 | 62.692.21 | 44.932.68 | |
| Crossformer | 50.181.97 | 45.971.84 | 46.301.73 | 45.901.84 | 66.681.67 | 48.651.89 | |
| Crossformer + CIF | 54.581.22 | 47.850.67 | 48.960.89 | 48.220.72 | 69.101.05 | 52.230.98 | |
| FEDformer | 45.750.78 | 45.711.29 | 44.271.28 | 43.511.00 | 62.641.64 | 45.881.35 | |
| FEDformer + CIF | 48.631.99 | 46.971.58 | 46.871.63 | 46.691.54 | 65.562.14 | 48.302.30 | |
| Informer | 48.421.99 | 46.941.60 | 46.410.99 | 45.760.43 | 65.991.14 | 47.491.07 | |
| Informer + CIF | 50.120.66 | 47.230.47 | 46.770.39 | 46.620.38 | 65.130.52 | 46.840.54 | |
| iTransformer | 52.851.36 | 46.971.05 | 47.311.03 | 46.840.78 | 67.460.96 | 49.900.89 | |
| iTransformer + CIF | 50.760.50 | 47.110.67 | 47.290.59 | 47.100.66 | 67.000.55 | 49.600.69 | |
| MTST | 45.771.70 | 44.391.73 | 43.701.82 | 43.361.98 | 61.381.57 | 44.011.60 | |
| MTST + CIF | 46.360.93 | 45.040.65 | 45.470.88 | 44.870.81 | 63.571.24 | 47.351.36 | |
| Nonformer | 50.811.06 | 48.711.40 | 48.551.47 | 48.361.38 | 66.951.54 | 48.081.82 | |
| Nonformer + CIF | 51.812.11 | 49.930.81 | 49.660.36 | 49.100.57 | 68.590.60 | 50.440.94 | |
| Reformer | 51.282.60 | 49.682.75 | 49.642.02 | 48.452.06 | 69.202.53 | 51.743.24 | |
| Reformer + CIF | 52.500.91 | 52.102.91 | 49.012.74 | 47.534.23 | 68.251.79 | 50.442.59 | |
| Transformer | 50.530.94 | 49.310.87 | 48.571.23 | 48.421.28 | 67.980.90 | 49.071.35 | |
| Transformer + CIF | 52.842.50 | 51.272.33 | 51.532.37 | 51.102.21 | 70.252.42 | 52.352.94 | |
| Medformer | 53.701.18 | 51.511.32 | 50.491.48 | 50.351.53 | 70.481.17 | 50.911.13 | |
| Medformer + CIF | 55.880.82 | 51.911.90 | 50.801.63 | 50.291.92 | 70.451.33 | 53.731.70 | |
| HM-BiTCN | 52.052.22 | 50.453.00 | 50.402.55 | 49.482.70 | 69.432.84 | 50.993.15 | |
| HM-BiTCN + CIF | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 |
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| PTB (2-Classes) | Autoformer | 71.992.74 | 69.603.85 | 61.504.23 | 61.435.07 | 74.291.89 | 70.262.00 |
| Autoformer + CIF | 77.710.63 | 77.151.29 | 70.321.62 | 71.771.60 | 81.204.15 | 78.134.80 | |
| Crossformer | 78.063.44 | 81.533.13 | 68.625.63 | 69.766.53 | 88.312.07 | 85.812.43 | |
| Crossformer + CIF | 84.554.80 | 85.963.63 | 78.647.49 | 80.257.89 | 92.172.99 | 91.163.70 | |
| FEDformer | 74.542.27 | 77.994.10 | 63.143.29 | 63.284.36 | 84.634.27 | 80.915.55 | |
| FEDformer + CIF | 77.931.86 | 80.653.50 | 68.592.20 | 70.232.63 | 86.003.51 | 83.424.76 | |
| Informer | 79.590.65 | 83.330.77 | 70.580.95 | 72.581.08 | 92.770.48 | 90.890.57 | |
| Informer + CIF | 83.632.32 | 85.311.80 | 77.283.67 | 79.373.45 | 93.870.48 | 92.290.63 | |
| iTransformer | 83.431.19 | 88.061.47 | 75.641.55 | 78.291.70 | 91.381.41 | 91.081.30 | |
| iTransformer + CIF | 81.622.96 | 87.482.03 | 72.744.41 | 74.994.91 | 90.253.29 | 89.803.28 | |
| MTST | 75.532.45 | 78.721.87 | 64.784.06 | 65.304.81 | 87.764.09 | 83.603.92 | |
| MTST + CIF | 81.801.60 | 84.192.05 | 74.282.05 | 76.522.25 | 91.302.48 | 88.413.30 | |
| Nonformer | 78.931.46 | 82.481.53 | 69.682.15 | 71.502.56 | 90.540.59 | 87.781.46 | |
| Nonformer + CIF | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |
| PatchTST | 75.282.44 | 77.052.44 | 64.864.05 | 65.415.28 | 88.112.59 | 82.652.87 | |
| PatchTST + CIF | 80.801.90 | 82.461.75 | 73.232.81 | 75.242.95 | 92.450.91 | 88.231.61 | |
| Reformer | 78.111.65 | 82.700.80 | 68.172.68 | 69.683.34 | 90.771.56 | 88.141.20 | |
| Reformer + CIF | 81.952.09 | 84.631.07 | 74.443.53 | 76.553.64 | 93.360.87 | 91.531.05 | |
| Transformer | 76.431.98 | 81.251.15 | 65.643.29 | 66.444.39 | 90.211.24 | 87.281.49 | |
| Transformer + CIF | 78.802.44 | 82.452.36 | 69.423.62 | 71.104.33 | 92.501.17 | 89.071.39 | |
| Medformer | 80.990.75 | 83.010.72 | 73.351.16 | 75.471.21 | 93.101.18 | 90.691.04 | |
| Medformer + CIF | 78.740.64 | 81.110.84 | 75.400.66 | 76.310.71 | 83.200.91 | 83.660.92 | |
| HM-BiTCN | 81.871.87 | 86.501.24 | 73.492.90 | 75.843.20 | 94.200.29 | 93.040.45 | |
| HM-BiTCN + CIF | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 |
During the process of transferring the optimized CIF hyperparameter configurations from the HM-BiTCN architecture to other model structures, we observed significant improvements in performance across most models and evaluation metrics. This result strongly demonstrates that the CIF module not only exhibits high generalizability but also possesses well-transferable hyperparameters that can be effectively adapted to various architectures. These hyperparameters enhance feature extraction and representation capabilities across different model designs. This finding further highlights the potential of the CIF architecture in diverse tasks and provides valuable insights for future model design and parameter sharing.
F.3 Fine-tuning Transferred Parameters
For models where the transferred hyperparameters from HM-BiTCN do not yield optimal results, we perform additional fine-tuning. This step investigates the adaptability of CIF-related parameters and explores how they can be optimized for other model structures.
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| TDBrain (2-Classes) Reproduced | iTransformer | 74.691.02 | 74.761.04 | 74.691.02 | 74.671.02 | 83.351.24 | 83.651.41 |
| iTransformer + CIF(HM-BiTCN) | 72.791.12 | 72.891.16 | 72.791.12 | 72.761.11 | 81.081.03 | 81.310.93 | |
| iTransformer + CIF(New) | 76.100.76 | 76.160.75 | 76.100.76 | 76.090.76 | 84.791.15 | 85.131.11 | |
| Nonformer | 88.102.39 | 88.761.74 | 88.102.39 | 88.042.47 | 96.560.91 | 96.361.21 | |
| Nonformer + CIF(HM-BiTCN) | 88.002.06 | 88.911.68 | 88.002.06 | 87.922.12 | 97.101.10 | 97.171.09 | |
| Nonformer + CIF (New) | 88.671.30 | 89.340.93 | 88.671.30 | 88.611.33 | 96.890.31 | 96.910.37 |
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| ADFTD (3-Classes) Reproduced | Autoformer | 46.902.89 | 45.592.37 | 44.912.23 | 44.342.52 | 63.492.44 | 45.632.29 |
| Autoformer + CIF((HM-BiTCN)) | 45.922.78 | 44.812.44 | 44.662.25 | 44.262.38 | 62.692.21 | 44.932.68 | |
| Autoformer + CIF(New) | 48.491.31 | 46.421.07 | 45.391.80 | 44.941.97 | 63.751.46 | 46.101.84 | |
| iTransformer | 52.851.36 | 46.971.05 | 47.311.03 | 46.840.78 | 67.460.96 | 49.900.89 | |
| iTransformer + CIF(HM-BiTCN) | 50.760.50 | 47.110.67 | 47.290.59 | 47.100.66 | 67.000.55 | 49.600.69 | |
| iTransformer + CIF(New) | 53.931.52 | 48.440.89 | 49.131.17 | 48.560.84 | 68.700.94 | 50.881.20 | |
| Reformer | 51.282.60 | 49.682.75 | 49.642.02 | 48.452.06 | 69.202.53 | 51.743.24 | |
| Reformer + CIF(HM-BiTCN) | 52.500.91 | 52.102.91 | 49.012.74 | 47.534.23 | 68.251.79 | 50.442.59 | |
| Reformer + CIF(New) | 53.111.71 | 52.141.42 | 52.421.90 | 51.732.01 | 70.471.64 | 51.881.92 | |
| Medformer | 53.701.18 | 51.511.32 | 50.491.48 | 50.351.53 | 70.481.17 | 50.911.13 | |
| Medformer + CIF(HM-BiTCN) | 55.880.82 | 51.911.90 | 50.801.63 | 50.291.92 | 70.451.33 | 53.731.70 | |
| Medformer + CIF(New) | 55.211.42 | 52.842.36 | 52.602.26 | 52.432.65 | 72.112.42 | 54.103.09 |
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| PTB (2-Classes) Reproduced | iTransformer | 83.431.19 | 88.061.47 | 75.641.55 | 78.291.70 | 91.381.41 | 91.081.30 |
| iTransformer + CIF(HM-BiTCN) | 81.622.96 | 87.482.03 | 72.744.41 | 74.994.91 | 90.253.29 | 89.803.28 | |
| iTransformer + CIF(New) | 83.561.15 | 86.001.03 | 76.781.94 | 79.101.87 | 90.400.75 | 89.800.93 | |
| Nonformer | 78.931.46 | 82.481.53 | 69.682.15 | 71.502.56 | 90.540.59 | 87.781.46 | |
| Nonformer + CIF(HM-BiTCN) | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |
| Nonformer + CIF(New) | 80.462.20 | 84.222.44 | 71.793.10 | 73.893.50 | 92.740.97 | 90.301.26 | |
| Medformer | 80.990.75 | 83.010.72 | 73.351.16 | 75.471.21 | 93.101.18 | 90.691.04 | |
| Medformer + CIF(HM-BiTCN) | 78.740.64 | 81.110.84 | 75.400.66 | 76.310.71 | 83.200.91 | 83.660.92 | |
| Medformer + CIF(New) | 83.980.81 | 85.210.62 | 78.001.30 | 80.111.22 | 93.320.97 | 90.971.69 |
For models that did not initially benefit from the direct transfer of CIF hyperparameters optimized for HM-BiTCN, we conducted further fine-tuning of the key CIF parameters. The results demonstrate that, after targeted adjustment, these models also achieved notable performance improvements. This process not only reinforces the adaptability of the CIF architecture across diverse models but also highlights the high tunability and flexibility of its hyperparameters. By appropriately modifying the number of selected channels, fusion direction, and scaling factors, CIF can be effectively tailored to different network structures, thereby maximizing its strengths in feature modeling and discriminative capability.
Appendix G Limitations and Future Work
Limitations:
Biomedical time series exhibit complex modal characteristics, which lead to significant efficiency bottlenecks when manually adjusting the prior parameters (t, n, a, b) with clear medical interpretations in the CIF model based on empirical experience. This limitation highlights the urgent need for developing novel automated hyperparameter optimization frameworks.
Future Work:
We plan to explore a more universal and generalizable time-series analysis approach, incorporating domain knowledge, structural modeling, and automated hyperparameter optimization. This integration should foster both deeper theoretical insights and stronger practical applicability, providing robust solutions for real-world medical problems.
Furthermore, integrating domain-specific prior knowledge into medical time-series analysis can reveal and model relationships between channels more precisely. By combining medical expertise, clinical experience, and existing pathological data, we can enhance model interpretability and predictive performance, supporting clinical decision-making and interventions. Future research could focus on incorporating such prior knowledge to advance intelligent medical analytics.
Finally, we must acknowledge that the development trends in the field of artificial intelligence highlight the importance of architectural innovation. Future research should focus on designing novel architectures that align more closely with the CIF method, combining the strengths of existing models. For example, the local feature extraction capabilities of CNNs LeCun et al. [1989], the temporal stability of TCNs Bai et al. [2018] for long sequences, the long-term dependency modeling of RNNs Rumelhart et al. [1986] and LSTMs Hochreiter and Schmidhuber [1997], the global modeling efficiency of Transformers Vaswani et al. [2017], the resource-efficient computation of Mamba Gu and Dao [2023], and the hybrid recurrence-attention structure of RWKV Peng et al. [2023]. By adapting and integrating these methods, we aim to build a powerful model that is not only deeply compatible with the CIF framework but also capable of efficiently handling complex medical time-series data.
Appendix H Broader Impacts
In this study, we take the first step toward a paradigm shift in medical time-series analysis, transitioning from a focus on model architecture improvements to emphasizing the intrinsic structure of the data. Our experimental results, which outperform several state-of-the-art methods across multiple datasets, validate the effectiveness of this approach. We emphasize that uncovering the essential characteristics in medical time-series data is a valuable research direction, particularly given the critical need for model interpretability in medical applications. However, practitioners should be cognizant of the method’s limitations.

### v2

@@
 Abstract
-The automatic classification of medical time series signals, such as electroencephalogram (EEG) and electrocardiogram (ECG), plays a pivotal role in clinical decision support and early detection of diseases.
-Although Transformer based models have achieved notable performance by implicitly modeling temporal dependencies through self-attention mechanisms, their inherently complex architectures and opaque reasoning processes undermine their trustworthiness in high stakes clinical settings.
-In response to these limitations, this study shifts focus toward a modeling paradigm that emphasizes structural transparency, aligning more closely with the intrinsic characteristics of medical data.
-We propose a novel method, Channel Imposed Fusion (CIF), which enhances the signal-to-noise ratio through cross-channel information fusion, effectively reduces redundancy, and improves classification performance.
-Furthermore, we integrate CIF with the Temporal Convolutional Network (TCN), known for its structural simplicity and controllable receptive field, to construct an efficient and explicit classification framework.
-Experimental results on multiple publicly available EEG and ECG datasets demonstrate that the proposed method not only outperforms existing state-of-the-art (SOTA) approaches in terms of various classification metrics, but also significantly enhances the transparency of the classification process, offering a novel perspective for medical time series classification.
+Medical time series (MedTS) such as EEG and ECG are critical for clinical diagnosis, yet existing deep learning approaches often struggle with two key challenges: the misalignment between domain-specific physiological knowledge and generic architectures, and the inherent low signal-to-noise ratio (SNR) of MedTS.
+To address these limitations, we shift from a conventional model-centric paradigm toward a data-centric perspective grounded in physiological principles.
+We propose Channel-Imposed Fusion (CIF), a method that explicitly encodes causal inter-channel relationships by linearly combining signals under domain-informed constraints, thereby enabling interpretable signal enhancement and noise suppression.
+To further demonstrate the effectiveness of data-centric design, we develop a simple yet powerful model, Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN), which, when combined with CIF, consistently outperforms Transformer-based approaches on multiple MedTS benchmarks and achieves new state-of-the-art performance on general time series classification datasets.
+Moreover, CIF is architecture-agnostic and can be seamlessly integrated into mainstream models such as Transformers, enhancing their adaptability to medical scenarios.
+Our work highlights the necessity of rethinking MedTS classification from a data-centric perspective and establishes a transferable framework for bridging physiological priors with modern deep learning architectures.
+The complete source code supporting this study is publicly available at the following Link: https://github.com/Xi-Mu-Yu/CIF.
 1 Introduction
@@
 (2023); Zhu et al.
-(2025).
+(2025); Wang et al.
+(2024b; 2025b).
 Accurate classification of these signals enables early anomaly detection, personalized treatment, and optimized therapy planning, ultimately improving patient outcomes and healthcare efficiency Liu et al.
@@
 In recent years, Transformer models Vaswani et al.
-(2017), powered by the self-attention mechanism Bahdanau et al.
-(2014), have achieved remarkable progress in time-series modeling, particularly in capturing long-range dependencies and global contextual information Liu et al.
+(2017), originally inspired by the self-attention mechanism Bahdanau et al.
+(2014), have achieved remarkable progress in time series modeling, particularly in capturing long-range dependencies and global contextual information Liu et al.
 (2021); Zhou et al.
 (2021).
-By converting sequential data into high-dimensional token embeddings, Transformers are capable of modeling complex temporal dependencies implicitly.
-However, despite their strong performance across a range of time-series tasks, several limitations have emerged when applying Transformer-based architectures to medical time-series classification, particularly in the following aspects: (1) Conflict Between Implicit Temporal Modeling and Explicit Physiological Mechanisms.
-Conventional Transformer architectures employ self attention mechanisms to implicitly represent temporal dependencies .
-However, essential physiological dynamics in medical signals, including conduction delays between ECG leads and synchronization of EEG rhythms, require explicit structural representations.
-Although methods such as Medformer Wang et al.
-(2024) have adopted strategies involving multiple scales of embedding, their attention distributions still fail to directly correspond to the underlying biological processes.
-This leads to fundamental limitations in model interpretability.
-(2) Mismatch Between Medical Domain Knowledge and Architectures Designed for General Purposes.
-Mainstream time series modeling methods such as Autoformer Wu et al.
-(2021), Informer Zhou et al.
-(2021), and FEDformer Zhou et al.
-(2022) have achieved strong performance in general domains like weather forecasting and finance.
-However, their architectural designs fail to account for the unique structural characteristics of medical time series.
-Medical signals often exhibit high inter-channel redundancy, spatial dependencies among physiological leads, and distinct local morphological patterns.
-These important features are easily weakened or overlooked in the high-dimensional and implicit modeling processes of conventional Transformer architectures.
-To address these limitations, we rethink the modeling paradigm from the perspective of data representation and break away from the conventional approach of implicitly discovering unknown temporal relationships.
-Unlike existing methods that rely on implicit learning, we advocate for explicitly modeling temporal relationships that have clear physical or physiological foundations.
-Based on this principle, we propose the Channel-Imposed Fusion (CIF) method, which encodes prior causal structures directly into the feature representation.
-Specifically, CIF constructs new features through a linear combination of the form , where and represent signals from two different channels, and and are predefined coefficients derived from domain knowledge.
-By embedding such structured prior knowledge into the model, CIF not only enhances the ability to capture critical pathological patterns but also significantly improves robustness and interpretability when dealing with noisy or irregular medical time-series data.
-This approach provides a more intuitive and effective framework for medical time-series analysis, contributing to the development of reliable clinical decision support systems.
-Based on the design philosophy of the CIF method, we aim for the model to maintain high efficiency while ensuring good interpretability and simplifying the structure as much as possible.
-To this end, we propose a simple yet effective Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN) architecture.
-This architecture utilizes a hierarchical causal convolution module, which explicitly encodes the coupling relationships between channels and temporal dynamics, effectively avoiding the issues of poor interpretability and structural complexity caused by the implicit modeling of temporal dependencies in Transformer-based methods.
-Moreover, it is worth noting that the CIF method is not limited to the HM-BiTCN architecture; it has high portability and can be seamlessly integrated with existing Transformer architectures to enhance the model’s adaptability and interpretability in medical time-series data.
-Our contributions are as follows:
-(1) Proposing the Channel-Imposed Fusion (CIF) method: We propose the CIF method to explicitly model the relationships between channel dimensions in medical time-series data, particularly for signals with clear physiological structures, such as EEG and ECG.
-(2) Designing the HM-BiTCN model based on CIF: We leverage the CIF method to construct the Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN), which explicitly models channel coupling and temporal dependencies.
-By adopting this TCN-based approach, HMBiTCN breaks the dominance of transformer models (SOTA) in MedTS classification and surpasses them across multiple datasets.
-(3) Providing methodological transferability: The CIF method is not only applicable to the HM-BiTCN model but can also be seamlessly integrated into existing Transformer architectures.
+By mapping sequential data into high-dimensional token embeddings, Transformers are able to implicitly model complex temporal dependencies.
+Despite their success across a wide range of time series tasks, applying Transformer architectures to MedTS classification still faces several challenges, which can be summarized as follows: (1) Misalignment between domain-specific knowledge and generic architectures.
+Mainstream time series models, such as Autoformer Wu et al.
+(2021), Crossformer Zhang & Yan (2022), and Reformer Kitaev et al.
+(2019), have demonstrated strong performance in general domains such as weather forecasting and finance.
+However, as illustrated in Fig.
+1, these approaches fail to achieve comparable effectiveness in MedTS classification tasks.
+This raises the urgent question of how to enhance the applicability of general-purpose models in medical scenarios.
+Moreover, MedTS often encode critical physiological characteristics—for example, conduction delays across ECG leads Auricchio et al.
+(2014) and rhythmic synchrony in EEG signals Palva & Palva (2014); Fries (2015)—which inherently reflect channel-level relationships.
+Unfortunately, such physiological dependencies are rarely considered in generic time series modeling frameworks.
+(2) Overemphasis on model optimization while neglecting the intrinsic low SNR of MedTS.
+Unlike general-purpose time series tasks, MedTS are characterized by pronounced low-SNR conditions Del Rio et al.
+(2011); Sraitih et al.
+(2022); Sharma (2017); Mohd Apandi et al.
+(2020); Jia et al.
+(2024), where noise and artifacts can easily overshadow critical physiological features.
+In such conditions, complex Transformer architectures do not always succeed in stably extracting effective representations, while simpler models (e.g., TCNs Bai et al.
+(2018)) may also experience more severe performance degradation.
+Indeed, recent Transformer-based methods tailored for MedTS, such as MedGNN Fan et al.
+(2025) and Medformer Wang et al.
+(2024a), primarily rely on architectural innovations, yet they fall short in fundamentally addressing the low-SNR challenge.
+This raises a key question: should breakthroughs in MedTS classification come from increasingly complex architectures, or from more principled data processing and representation strategies?
+To address the aforementioned limitations, we depart from the traditional model-centric paradigm that relies on increasingly complex architectures to capture temporal dependencies, and instead propose a data-centric approach grounded in the physiological properties of medical time series.
+Following this principle, we introduce the Channel-Imposed Fusion (CIF) method, which explicitly encodes prior causal structures into feature representations.
+Specifically, CIF constructs new features through a linear combination of signals from different channels:
+where and denote signals from two distinct channels, and and are coefficients predefined based on domain knowledge.
+When and take fixed values, they are not learned directly from patient data, but instead derived from two domain-specific prior hypotheses: (1) Physiological Coupling Hypothesis.
+For ECG signals, when two leads are highly correlated (e.g., P-wave polarity and morphology are consistent Platonov (2012)), setting achieves in-phase summation, thereby enhancing target signal components and improving the SNR.
+(2) Noise Suppression Hypothesis.
+In EEG recordings, ocular artifacts such as blinks often appear highly correlated in frontal electrodes Fp1 and Fp2 Croft & Barry (2000).
+To suppress such noise, we set , applying a differential fusion strategy to cancel common-mode interference.
+Here, the coefficients and serve as symbolic encodings of interpretable physiological principles, rather than exact data-driven estimates.
+When treated as learnable parameters, they can be fine-tuned under symbolic constraints imposed by prior knowledge (e.g., enforcing under coupling, and under noise suppression).
+This design maintains the interpretability of directional relationships (e.g., signal enhancement or cancellation) while allowing the model to adaptively adjust the magnitude of each coefficient based on the training data.
+To emphasize the importance of data-centric approaches, we deliberately designed a simple yet effective model—the Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN)—to demonstrate that excellent performance does not necessarily require model complexity.
+The combination of CIF and HM-BiTCN not only outperforms Transformer-based methods on multiple medical datasets but also achieves new state-of-the-art (SOTA) results on general time series classification benchmarks.
+More importantly, the CIF method is not limited to the HM-BiTCN architecture itself; it exhibits strong transferability and can be seamlessly integrated into existing Transformer architectures, enhancing their adaptability to MedTS data.
+Our main contributions are:
+Proposal of Channel-Imposed Fusion (CIF).
+We introduce CIF to explicitly model inter-channel relationships in medical time series, particularly suitable for signals with well-defined physiological structures such as EEG and ECG.
+Design of HM-BiTCN based on CIF.
+By integrating CIF into HM-BiTCN, our method consistently outperforms existing SOTA models across multiple publicly available medical and non-medical time series classification datasets.
+Methodological transferability.
+CIF is architecture-agnostic and can be seamlessly integrated into mainstream models such as Transformers, compensating for the limitations of traditional positional encodings in modeling channel-level correlations, and highlighting the paradigm shift from a model-centric to a data-centric perspective.
 2 Related Work
@@
 Medical time series analysis diverges fundamentally from general time series forecasting Wu et al.
-(2022); Lu et al.
+(2022a); Lu et al.
 (2024) by prioritizing pathological signature decoding over temporal extrapolation, with modalities like EEG Tang et al.
@@
 Early methods were dominated by compact CNNs such as EEGNet Lawhern et al.
-(2018), which employs depthwise separable convolutions to efficiently extract spatio–temporal features while providing preliminary interpretability via feature-map visualization.
+(2018), which employs depthwise separable convolutions to efficiently extract spatio–temporal features while providing preliminary interpretability via feature‐map visualization.
 Subsequently, temporal convolutional networks (TCNs) Bai et al.
 (2018); Lin et al.
-(2019) leveraging dilated causal convolutions achieved parallelizable computation and extended receptive fields, surpassing LSTM-based approaches Zhou et al.
-(2016); Shen and Lee (2016); Hochreiter and Schmidhuber (1997) on multiple medical signal classification benchmarks.
-Hybrid architectures such as EEG-Conformer Song et al.
-(2022) combined convolutional front-ends with Transformer self-attention to capture both local and global dependencies and enabled attention-based interpretability.
+(2019) leveraging dilated causal convolutions achieved parallelizable computation and extended receptive fields, surpassing LSTM‐based approaches Zhou et al.
+(2016); Shen & Lee (2016); Hochreiter & Schmidhuber (1997) on multiple medical signal classification benchmarks.
+Hybrid architectures such as EEG‐Conformer Song et al.
+(2022) combined convolutional front‐ends with Transformer self‐attention to capture both local and global dependencies and enabled attention‐based interpretability.
 More recently, fine-grained Transformer models such as Medformer Wang et al.
-(2024) introduced cross-channel tokenization and dual-stage self-attention, setting new SOTA accuracy on several public datasets.
+(2024a) introduced cross-channel tokenization and dual-stage self-attention, setting new SOTA accuracy on several public datasets.
 The latest MedGNN Fan et al.
-(2025) further augments attention mechanisms with multi-resolution graph learning to jointly model spatial multi-scale channel dependencies and temporal dynamics.
-Although Transformer models have significantly improved classification accuracy, their complexity, especially due to deep architectures and multi-layer attention mechanisms, results in reduced interpretability.
-In clinical applications, where model transparency and reliability are essential, it is imperative to strike a balance between high accuracy and explainability.
-Implicit modeling strategies in Transformer-based temporal methods.
-In time series analysis, Transformer-based models implicitly learn complex dependencies through diversified architectural designs: the vanilla Transformer Vaswani et al.
+(2025) further augments attention mechanisms with multi‐resolution graph learning to jointly model spatial multi‐scale channel dependencies and temporal dynamics.
+Model-centric Transformer-based time series methods.
+In time series analysis, Transformer-based models learn complex dependencies through diversified architectural designs: the vanilla Transformer Vaswani et al.
 (2017) first introduced multi-head self-attention and sinusoidal positional encoding to model temporal correlations globally; Informer Zhou et al.
@@
 (2021) proposes an Auto-Correlation mechanism that aggregates periodic subsequences to enhance the implicit capture of cyclic patterns; FEDformer Zhou et al.
-(2022) performs seasonal–trend decomposition in the frequency domain and uses compressed Fourier coefficients to enable cross-frequency attention interactions; Crossformer Zhang and Yan (2022) designs a two-stage attention mechanism across time and feature dimensions to implicitly fuse multivariate spatiotemporal couplings; iTransformer Liu et al.
+(2022) performs seasonal–trend decomposition in the frequency domain and uses compressed Fourier coefficients to enable cross-frequency attention interactions; Crossformer Zhang & Yan (2022) designs a two-stage attention mechanism across time and feature dimensions to implicitly fuse multivariate spatiotemporal couplings; iTransformer Liu et al.
 (2024b) innovatively treats time steps as channel dimensions and applies standard attention to implicitly learn nonlinear inter-variable relationships; PatchTST Nie et al.
 (2023) segments continuous time steps into patch-based tokens and uses a combination of local and global attention to capture multi-scale temporal patterns; Medformer Wang et al.
-(2024) introduces multi-granularity patch embeddings and cross-channel attention for medical signals, implicitly modeling the heterogeneous couplings of physiological metrics; and MedGNN Fan et al.
+(2024a) introduces multi-granularity patch embeddings and cross-channel attention for medical signals, implicitly modeling the heterogeneous couplings of physiological metrics; and MedGNN Fan et al.
 (2025) combines graph attention with frequency-differential networks to incorporate medical topological priors into implicit spatiotemporal dependency learning.
-Although these variants enhance time series modeling via attention-based techniques, their reliance on high-dimensional weight matrices to implicitly encode dependencies between time steps and variables hinders interpretability.
-This limitation is particularly severe in medical applications, where it undermines the trustworthiness of high-risk tasks such as disease diagnosis and risk prediction.
-3 Preliminaries and Problem Formulation
-We define the MedTS classification task as a supervised learning problem over time series data with hierarchical subject-level structure.
-The dataset is represented as , where is the -th sample with time steps and channels, is a multi-label vector over categories, and is the subject ID.
-Samples from the same subject share the same label, i.e., , where represents the set of samples from subject .
-We consider two dataset partitioning strategies: (i) Subject-Dependent Split, where the dataset is split at the sample level such that samples from the same subject may appear in both training and test sets, which can cause information leakage and yield overly optimistic performance estimates; and (ii) Subject-Independent Split, where the dataset is split at the subject level, ensuring that each subject appears only in one of the train, validation, or test sets, simulating real-world diagnostic scenarios but introducing challenges due to inter-subject variability.
-The goal is to train an encoder to obtain a latent representation , which is then passed to a classifier to predict the label .
 3 Method
 3.1 Channel-Imposed Fusion
-In this section, we analyze the data fusion process using Singular Value Decomposition (SVD) Golub and Reinsch (1971); Klema and Laub (1980).
+In this section, we analyze the data fusion process using Singular Value Decomposition (SVD) Golub & Reinsch (1971); Klema & Laub (1980); Harner (1990); Lagerlund et al.
+(1997).
 Suppose the original time-series data matrix has dimensions , where represents the temporal length and denotes the number of channels.
-We partition along the channel dimension into two submatrices: the first channels form , while the last channels form .
+As shown in Fig.2, we partition along the channel dimension into two submatrices: the first channels form , while the last channels form .
 The goal of fusion is to combine these two submatrices into a new representation:
@@
 Thus, the fusion process allows for either redundancy reduction by preserving shared patterns or diversity enhancement by integrating complementary information, depending on the correlation between the front and back segments.
+It is worth noting that no matter how the channels are selected for partitioning and fusion, this operation is meaningful.
+Each group of channels contains a portion of the time-series information, and by performing SVD decomposition and linear combination on different channel subsets, one can explore their potential similarities and complementarities, thereby achieving either redundancy reduction or diversity enhancement.
+Therefore, this partitioning strategy is generally applicable and is not limited to the first and the last channels.
 Optimizing SNR via CIF: Let two observed signals be and , where are zero-mean useful signals with variance and correlation coefficient , and are zero-mean noise components with variance and correlation coefficient .
@@
 Both modes optimize the SNR by adjusting parameters and , showcasing CIF’s ability to suppress noise and enhance the signal under varying correlations.
-Gradient Coupling Mechanism: The coefficients and are optimized through gradient descent, where their updates are coupled due to shared dependencies on and . The gradients as:
-Specifically, the gradient computations of the loss function with respect to and include information from each other’s signals, requiring their updates to simultaneously consider the influence of both forward and backward channels.
-This coupling mechanism enables the model to learn cross-channel interaction patterns, enhancing feature representation and generalization capability.
-In Appendix B, we visualize scenarios with highly and generally correlated channel information.
 3.2 HM-BiTCN Structure Design and Theoretical Analysis
-Modeling short-term and long-term dependencies in time series data is challenging.
-Traditional CNNs excel at capturing local features but have limited receptive fields, hindering long-range dependency learning.
-Transformer-based methods effectively model long-term dependencies, but their complex design lacks interpretability, which is a key issue in medical time-series classification.To address these limitations, TCNs use causal convolutions for explicit temporal modeling and dilated convolutions to expand the receptive field, overcoming the constraints of traditional CNNs.
-Building on the advantages of TCN, we propose the HM-BiTCN, which combines the benefits of dilated convolutions, bidirectional causal convolution, and residual connections.
-This approach allows for better capture of temporal dependencies while preserving causality.
-Our goal is to simplify the model structure through explicit design principles while improving its interpretability.
-4.2.1 Dilated Convolution
-Dilated convolution expands the receptive field without significantly increasing computational cost Yu and Koltun (2015).
-For a 1D input sequence , its output is defined as , where is the current time step, is the kernel size, is the dilation factor, and is the weight at the -th position in the kernel.
-Increasing effectively enlarges the receptive field, enabling the network to capture longer-term temporal dependencies.
-When stacking multiple dilated convolutional layers, the receptive field grows progressively.
-For the -th layer, the receptive field can be expressed as , where is the dilation factor of the -th layer.
-By gradually increasing , the network captures temporal dependencies across both global and local scales, offering an effective way to model long-term dependencies in time series.
-4.2.2 Bidirectional Causal Convolution Structure
-In addition to dilated convolutions, HM-BiTCN introduces a bidirectional causal convolution structure, inspired by prior bidirectional temporal modeling approaches Hanson et al.
-(2018); Hu et al.
-(2024); Yin et al.
-(2025).
-Unlike traditional TCNs that use only forward causal convolutions, our architecture applies causal convolutions in both forward and backward directions, enabling the model to capture dependencies from both past and future contexts while strictly preserving causality.
-The forward causal convolution processes the input sequence in chronological order, producing output , which depends only on current and past inputs.
-For the backward causal convolution, we first reverse the input sequence as , and then apply a causal convolution over this flipped sequence.
-This ensures that the model captures future-directed dependencies without introducing information leakage.
-The output is given by .
-These two operations are implemented using separate convolutional layers (convforward and convbackward), and their outputs are summed to form the final bidirectional result: .
-By integrating both directions under strict causality constraints, HM-BiTCN achieves superior temporal dependency modeling compared to unidirectional causal approaches.
-4.2.3 Multi-Scale Feature Learning and Residual Connections
-To further improve the model’s capacity to capture dependencies at different temporal scales, HM-BiTCN incorporates multi-scale feature learning and residual connections.
-Multi-scale Feature Learning: In HM-BiTCN, we employ a hierarchy of dilation factors that decrease layer by layer to capture temporal dependencies at multiple scales.
-Lower layers use larger dilation factors to expand the receptive field, aggregating long-range information and smoothing short-term noise in highly redundant medical time series; higher layers use smaller dilation factors to focus on local dependencies and capture fine-grained features.
-This coarse-to-fine, global-to-local design enables the network to extract broad patterns in its initial layers and refine precise details in its later layers, thereby enhancing adaptability across a wide range of time series tasks.
-Residual connections: Residual connections He et al.
-(2016) are introduced between the dilated convolutional layers to facilitate the efficient flow of information through the network.
-The residual connection is defined as , where is the convolutional output, and is the input.
-This design alleviates the vanishing gradient problem and improves the overall stability of the network during training.
+To demonstrate the value of data-centric approaches and to show that simple models can also achieve strong performance, as shown in Figure3 (a), the conventional TCN only considers unidirectional causal relationships in time series.
+In contrast, as shown in Figure3 (c), our proposed HM-BiTCN simply introduces bidirectional relationships and performs feature mixing at each layer.
+This design not only further enhances the signals with improved SNR through CIF but also captures both forward and backward causal structures, thereby making fuller use of the information contained in the data.
+See Appendix B and Appendix C for details.
+We must emphasize that such a simple modification does not constitute an innovation in itself.
+Our motivation is to demonstrate that simple models, when combined with a data-centric approach, can also surpass existing SOTA models.
 4 Experiments
-Datasets.
+Medical Time Series Datasets.
 (1) APAVA Escudero et al.
@@
 (3) ADFTD Miltiadous et al.
-(2023b, a) is an EEG dataset with a three-class label for each sample, categorizing the subject as Healthy, having Frontotemporal Dementia, or Alzheimer’s disease.
+(2023b; a) is an EEG dataset with a three-class label for each sample, categorizing the subject as Healthy, having Frontotemporal Dementia, or Alzheimer’s disease.
 (4) PTB PhysioBank (2000) is an ECG dataset where each sample is labeled with a binary indicator of Myocardial Infarction.
@@
 (2020) is an ECG dataset with a five-class label for each sample, representing various heart conditions.
-Table 4 provides information on the processed datasets.The data processing methodology is the same as that of Medformer Wang et al.
-(2024) and MedGNN Fan et al.
-(2025).
 Baselines.
 We compare with 12 state-of-the-art time series transformer methods: Autoformer Wu et al.
-(2021), Crossformer Zhang and Yan (2022), FEDformer Zhou et al.
+(2021), Crossformer Zhang & Yan (2022), FEDformer Zhou et al.
 (2022), Informer Zhou et al.
@@
 (2017), Medformer Wang et al.
-(2024), MedGNN Fan et al.
+(2024a), MedGNN Fan et al.
 (2025).
@@
 (2017).
+We consider two dataset partitioning strategies: (i) Subject-Dependent Split, where the dataset is split at the sample level such that samples from the same subject may appear in both training and test sets, which can cause information leakage and yield overly optimistic performance estimates; and (ii) Subject-Independent Split, where the dataset is split at the subject level, ensuring that each subject appears only in one of the train, validation, or test sets, simulating real-world diagnostic scenarios but introducing challenges due to inter-subject variability.
 4.1 Results of Subject-Dependent
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| ADFTD (3-Classes) Reported | Autoformer Wu et al. (2021) | 87.831.62 | 87.631.66 | 87.221.97 | 87.381.79 | 96.590.88 | 93.821.64 |
-| Crossformer Zhang and Yan (2022) | 89.351.32 | 89.001.44 | 88.791.37 | 88.881.40 | 97.520.58 | 95.451.03 | |
-| FEDformer Zhou et al. (2022) | 77.632.37 | 76.762.17 | 76.682.48 | 76.602.46 | 91.671.34 | 84.942.11 | |
-| Informer Zhou et al. (2021) | 90.930.90 | 90.740.71 | 90.501.14 | 90.600.94 | 98.190.27 | 96.510.49 | |
-| iTransformer Liu et al. (2024b) | 64.900.25 | 62.530.27 | 62.210.26 | 62.250.33 | 81.520.29 | 68.870.49 | |
-| MTST Zhang et al. (2024) | 65.080.69 | 63.850.80 | 62.710.64 | 63.030.58 | 81.360.56 | 69.340.89 | |
-| Nonformer Liu et al. (2022) | 96.120.47 | 95.940.56 | 95.990.38 | 95.960.47 | 99.590.09 | 99.080.16 | |
-| PatchTST Nie et al. (2023) | 66.260.40 | 65.080.41 | 64.970.51 | 64.950.42 | 83.070.45 | 71.700.61 | |
-| Reformer Kitaev et al. (2019) | 91.511.75 | 91.151.79 | 91.651.56 | 91.141.83 | 98.850.35 | 97.880.60 | |
-| Transformer Vaswani et al. (2017) | 97.000.43 | 96.870.53 | 96.860.36 | 96.860.44 | 99.750.04 | 99.420.07 | |
-| Medformer Wang et al. (2024) | 97.620.34 | 97.530.33 | 97.480.40 | 97.500.36 | 99.830.05 | 99.620.12 | |
-| MedGNN Fan et al. (2025) | 98.420.04 | 98.310.02 | 98.290.05 | 98.300.12 | 99.930.11 | - | |
-| Medformer + CIF | 98.870.26 | 98.770.27 | 98.860.27 | 98.810.27 | 99.960.01 | 99.920.03 | |
+| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC | ||||
+Autoformer Wu et al. (2021) | 87.831.62 | 87.631.66 | 87.221.97 | 87.381.79 | 96.590.88 | 93.821.64 | ||||
+| Crossformer Zhang & Yan (2022) | 89.351.32 | 89.001.44 | 88.791.37 | 88.881.40 | 97.520.58 | 95.451.03 | |||||
+| FEDformer Zhou et al. (2022) | 77.632.37 | 76.762.17 | 76.682.48 | 76.602.46 | 91.671.34 | 84.942.11 | |||||
+| Informer Zhou et al. (2021) | 90.930.90 | 90.740.71 | 90.501.14 | 90.600.94 | 98.190.27 | 96.510.49 | |||||
+| iTransformer Liu et al. (2024b) | 64.900.25 | 62.530.27 | 62.210.26 | 62.250.33 | 81.520.29 | 68.870.49 | |||||
+| MTST Zhang et al. (2024) | 65.080.69 | 63.850.80 | 62.710.64 | 63.030.58 | 81.360.56 | 69.340.89 | |||||
+| Nonformer Liu et al. (2022) | 96.120.47 | 95.940.56 | 95.990.38 | 95.960.47 | 99.590.09 | 99.080.16 | |||||
+| PatchTST Nie et al. (2023) | 66.260.40 | 65.080.41 | 64.970.51 | 64.950.42 | 83.070.45 | 71.700.61 | |||||
+| Reformer Kitaev et al. (2019) | 91.511.75 | 91.151.79 | 91.651.56 | 91.141.83 | 98.850.35 | 97.880.60 | |||||
+| Transformer Vaswani et al. (2017) | 97.000.43 | 96.870.53 | 96.860.36 | 96.860.44 | 99.750.04 | 99.420.07 | |||||
+| Medformer Wang et al. (2024a) | 97.620.34 | 97.530.33 | 97.480.40 | 97.500.36 | 99.830.05 | 99.620.12 | |||||
+| MedGNN Fan et al. (2025) | 98.420.04 | 98.310.02 | 98.290.05 | 98.300.12 | 99.930.11 | - | |||||
+| Medformer + CIF | 98.870.26 | 98.770.27 | 98.860.27 | 98.810.27 | 99.960.01 | 99.920.03 | |||||
 | MedGNN + CIF | 99.600.09 | 99.600.11 | 99.580.09 | 99.590.10 | 99.990.01 | 99.970.01 |
-To further validate our proposed method, we reproduced 12 baseline approaches.
-Table 1 presents a comparison of the reported experimental results for these methods, while Table 5 in Appendix E provides detailed results from our reproduction of the 12 methods.
-The experimental results demonstrate that by incorporating our proposed CIF method into the Medformer model, we not only achieve significantly better performance than the original methods, but also outperform the current state-of-the-art (SOTA) approach, MedGNN.
-These findings strongly confirm the effectiveness and superiority of our method.
+We reproduced 12 baselines.
+Table 1 lists their reported results.
+Experimental results show that integrating the CIF method into MedGNN and Medformer outperforms existing approaches, fully demonstrating its effectiveness and superiority.
 4.2 Results of Subject-Independent
-In the Subject-Independent setup, the training, validation, and test sets are split based on subjects.
-All subjects and their corresponding samples are distributed into these three sets according to a predetermined ratio or subject IDs.
-Samples from the same subject exclusively appear in one of the three sets.
-This setup simulates real-world MedTS-based disease diagnosis, aiming to train a model on subjects with known labels and then test it on unseen subjects to determine if they have a specific disease.
-All five datasets are evaluated under this setup.
-Table 2 presents the results reported by various methods in the subject-independent setting, while Table 6 shows the results of our reproduction of these methods.
-Our method achieves the highest scores in six metrics across four out of the five datasets.
-For the PTB-XL dataset, compared to the results reported by other methods, our approach achieves the highest AUROC and AUPRC score and ranks second in Accuracy.
-When compared to the results of our reproduction, our method ranks first in Accuracy, Precision, AUROC, and AUPRC, and second in Recall.
+| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC | ||||
+Autoformer Wu et al. (2021) | 68.641.82 | 68.482.10 | 68.772.27 | 68.061.94 | 75.943.61 | 74.384.05 | ||||
+| Crossformer Zhang & Yan (2022) | 73.771.95 | 79.294.36 | 68.861.70 | 68.931.85 | 72.393.33 | 72.053.65 | |||||
+| FEDformer Zhou et al. (2022) | 74.942.15 | 74.591.50 | 73.563.55 | 73.513.39 | 83.721.97 | 82.942.37 | |||||
+| Informer Zhou et al. (2021) | 73.114.40 | 75.176.06 | 69.174.56 | 69.475.06 | 70.464.91 | 70.755.27 | |||||
+| iTransformer Liu et al. (2024b) | 74.551.66 | 74.772.10 | 71.761.72 | 72.301.79 | 85.591.55 | 84.391.57 | |||||
+| MTST Zhang et al. (2024) | 71.141.59 | 79.300.97 | 65.272.28 | 64.013.16 | 68.872.34 | 71.061.60 | |||||
+| Nonformer Liu et al. (2022) | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |||||
+| PatchTST Nie et al. (2023) | 67.031.65 | 78.761.28 | 59.912.02 | 55.973.10 | 65.650.28 | 67.990.76 | |||||
+| Reformer Kitaev et al. (2019) | 78.702.00 | 82.503.95 | 75.001.61 | 75.931.82 | 73.941.40 | 76.041.14 | |||||
+| Transformer Vaswani et al. (2017) | 76.304.72 | 77.645.95 | 73.095.01 | 73.755.38 | 72.506.60 | 73.237.60 | |||||
+| Medformer Wang et al. (2024a) | 78.740.64 | 81.110.84 | 75.400.66 | 76.310.71 | 83.200.91 | 83.660.92 | |||||
+| MedGNN Fan et al. (2025) | 82.600.35 | 87.700.22 | 78.930.09 | 80.250.16 | 85.930.26 | - | |||||
+| HM-BiTCN + CIF | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | |||||
+Autoformer Wu et al. (2021) | 87.333.79 | 88.063.56 | 87.333.79 | 87.263.84 | 93.812.26 | 93.322.42 | ||||
+| Crossformer Zhang & Yan (2022) | 81.562.19 | 81.972.25 | 81.562.19 | 81.502.20 | 91.201.78 | 91.511.71 | |||||
+| FEDformer Zhou et al. (2022) | 78.131.98 | 78.521.91 | 78.131.98 | 78.042.01 | 86.561.86 | 86.481.99 | |||||
+| Informer Zhou et al. (2021) | 89.022.50 | 89.432.14 | 89.022.50 | 88.982.54 | 96.640.68 | 96.750.63 | |||||
+| iTransformer Liu et al. (2024b) | 74.671.06 | 74.711.06 | 74.671.06 | 74.651.06 | 83.371.14 | 83.731.27 | |||||
+| MTST Zhang et al. (2024) | 76.963.76 | 77.243.59 | 76.963.76 | 76.883.83 | 85.274.46 | 82.815.64 | |||||
+| Nonformer Liu et al. (2022) | 87.882.48 | 88.861.84 | 87.882.48 | 87.782.56 | 97.050.68 | 96.990.68 | |||||
+| PatchTST Nie et al. (2023) | 79.253.79 | 79.604.09 | 79.253.79 | 79.203.77 | 87.954.96 | 86.366.67 | |||||
+| Reformer Kitaev et al. (2019) | 87.922.01 | 88.641.40 | 87.922.01 | 87.852.08 | 96.300.54 | 96.400.45 | |||||
+| Transformer Vaswani et al. (2017) | 87.171.67 | 87.991.68 | 87.171.67 | 87.101.68 | 96.280.92 | 96.340.81 | |||||
+| Medformer Wang et al. (2024a) | 89.620.81 | 89.680.78 | 89.620.81 | 89.620.81 | 96.410.35 | 96.510.33 | |||||
+| MedGNN Fan et al. (2025) | 91.040.09 | 91.150.12 | 91.040.20 | 91.040.08 | 96.740.04 | - | |||||
+| HM-BiTCN + CIF | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 | |||||
+Autoformer Wu et al. (2021) | 45.251.48 | 43.671.94 | 42.962.03 | 42.591.85 | 61.021.82 | 43.102.30 | ||||
+| Crossformer Zhang & Yan (2022) | 50.452.31 | 45.571.63 | 45.881.82 | 45.501.70 | 66.452.03 | 48.332.05 | |||||
+| FEDformer Zhou et al. (2022) | 46.300.59 | 46.050.76 | 44.221.38 | 43.911.37 | 62.621.75 | 46.111.44 | |||||
+| Informer Zhou et al. (2021) | 48.451.96 | 46.541.68 | 46.061.84 | 45.741.38 | 65.871.27 | 47.601.30 | |||||
+| iTransformer Liu et al. (2024b) | 52.601.59 | 46.791.27 | 47.281.29 | 46.791.13 | 67.261.16 | 49.531.21 | |||||
+| MTST Zhang et al. (2024) | 45.602.03 | 44.701.33 | 45.051.30 | 44.311.74 | 62.500.81 | 45.160.85 | |||||
+| Nonformer Liu et al. (2022) | 49.951.05 | 47.710.97 | 47.461.50 | 46.961.35 | 66.231.37 | 47.331.78 | |||||
+| PatchTST Nie et al. (2023) | 44.370.95 | 42.401.13 | 42.061.48 | 41.971.37 | 60.081.50 | 42.491.79 | |||||
+| Reformer Kitaev et al. (2019) | 50.781.17 | 49.641.49 | 49.891.67 | 47.940.69 | 69.171.58 | 51.731.94 | |||||
+| Transformer Vaswani et al. (2017) | 50.472.14 | 49.131.83 | 48.011.53 | 48.091.59 | 67.931.59 | 48.932.02 | |||||
+| Medformer Wang et al. (2024a) | 53.271.54 | 51.021.57 | 50.711.55 | 50.651.51 | 70.931.19 | 51.211.32 | |||||
+| MedGNN Fan et al. (2025) | 56.120.11 | 55.070.09 | 55.470.34 | 55.000.24 | 74.680.33 | - | |||||
+| HM-BiTCN + CIF | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 | |||||
+Autoformer Wu et al. (2021) | 73.352.10 | 72.112.89 | 63.243.17 | 63.693.84 | 78.543.48 | 74.253.53 | ||||
+| Crossformer Zhang & Yan (2022) | 80.173.79 | 85.041.83 | 71.256.29 | 72.757.19 | 88.553.45 | 87.313.25 | |||||
+| FEDformer Zhou et al. (2022) | 76.052.54 | 77.583.61 | 66.103.55 | 67.144.37 | 85.934.31 | 82.595.42 | |||||
+| Informer Zhou et al. (2021) | 78.691.68 | 82.871.02 | 69.192.90 | 70.843.47 | 92.090.53 | 90.020.60 | |||||
+| iTransformer Liu et al. (2024b) | 83.890.71 | 88.251.18 | 76.391.01 | 79.061.06 | 91.181.16 | 90.930.98 | |||||
+| MTST Zhang et al. (2024) | 76.591.90 | 79.881.90 | 66.312.95 | 67.383.71 | 86.862.75 | 83.752.84 | |||||
+| Nonformer Liu et al. (2022) | 78.660.49 | 82.770.86 | 69.120.87 | 70.901.00 | 89.372.51 | 86.672.38 | |||||
+| PatchTST Nie et al. (2023) | 74.741.62 | 76.941.51 | 63.892.71 | 64.363.38 | 88.790.91 | 83.390.96 | |||||
+| Reformer Kitaev et al. (2019) | 77.962.13 | 81.721.61 | 68.203.35 | 69.653.88 | 91.130.74 | 88.421.30 | |||||
+| Transformer Vaswani et al. (2017) | 77.371.02 | 81.840.66 | 67.141.80 | 68.472.19 | 90.081.76 | 87.221.68 | |||||
+| Medformer Wang et al. (2024a) | 83.502.01 | 85.190.94 | 77.113.39 | 79.183.31 | 92.811.48 | 90.321.54 | |||||
+| MedGNN Fan et al. (2025) | 84.530.28 | 87.350.45 | 77.900.66 | 80.400.62 | 93.310.46 | - | |||||
+| HM-BiTCN + CIF | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 | |||||
+Autoformer Wu et al. (2021) | 61.682.72 | 51.601.64 | 49.101.52 | 48.852.27 | 82.041.44 | 51.931.71 | ||||
+| Crossformer Zhang & Yan (2022) | 73.300.14 | 65.060.35 | 61.230.33 | 62.590.14 | 90.020.06 | 67.430.22 | |||||
+| FEDformer Zhou et al. (2022) | 57.209.47 | 52.386.09 | 49.047.26 | 47.898.44 | 82.134.17 | 52.317.03 | |||||
+| Informer Zhou et al. (2021) | 71.430.32 | 62.640.60 | 59.120.47 | 60.440.43 | 88.650.09 | 64.760.17 | |||||
+| iTransformer Liu et al. (2024b) | 69.280.22 | 59.590.45 | 54.620.18 | 56.200.19 | 86.710.10 | 60.270.21 | |||||
+| MTST Zhang et al. (2024) | 72.140.27 | 63.840.72 | 60.010.81 | 61.430.38 | 88.970.33 | 65.830.51 | |||||
+| Nonformer Liu et al. (2022) | 70.560.55 | 61.570.66 | 57.750.72 | 59.100.66 | 88.320.36 | 63.400.79 | |||||
+| PatchTST Nie et al. (2023) | 73.230.25 | 65.700.64 | 60.820.76 | 62.610.34 | 89.740.19 | 67.320.22 | |||||
+| Reformer Kitaev et al. (2019) | 71.720.43 | 63.121.02 | 59.200.75 | 60.690.18 | 88.800.24 | 64.720.47 | |||||
+| Transformer Vaswani et al. (2017) | 70.590.44 | 61.570.65 | 57.620.35 | 59.050.25 | 88.210.16 | 63.360.29 | |||||
+| Medformer Wang et al. (2024a) | 72.870.23 | 64.140.42 | 60.600.46 | 62.020.37 | 89.660.13 | 66.390.22 | |||||
+| MedGNN Fan et al. (2025) | 73.870.18 | 66.260.29 | 61.130.29 | 62.540.20 | 90.210.15 | - | |||||
+| HM-BiTCN + CIF | 73.730.30 | 65.410.67 | 60.701.08 | 61.890.91 | 90.530.22 | 67.750.75 |
+Table 2 presents the results reported by various methods in the subject-independent setting.
+Our method achieves the highest average scores across six metrics on four out of the five datasets.
+On PTB-XL, our method tops AUROC and AUPRC and ranks second in Accuracy versus reported results, and ranks first in Accuracy, Precision, AUROC, AUPRC, and second in Recall versus our reproduced results.
 Additionally, it is worth noting that in the subject-independent setup, the F1 score of ADFTD is 55.42%, which is significantly lower than the 99.59% achieved in the subject-dependent setup.
 This comparison highlights the challenges of the subject-independent setup, which better simulates real-world scenarios.
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| APAVA (2-Classes) Reported | Autoformer Wu et al. (2021) | 68.641.82 | 68.482.10 | 68.772.27 | 68.061.94 | 75.943.61 | 74.384.05 |
-| Crossformer Zhang and Yan (2022) | 73.771.95 | 79.294.36 | 68.861.70 | 68.931.85 | 72.393.33 | 72.053.65 | |
-| FEDformer Zhou et al. (2022) | 74.942.15 | 74.591.50 | 73.563.55 | 73.513.39 | 83.721.97 | 82.942.37 | |
-| Informer Zhou et al. (2021) | 73.114.40 | 75.176.06 | 69.174.56 | 69.475.06 | 70.464.91 | 70.755.27 | |
-| iTransformer Liu et al. (2024b) | 74.551.66 | 74.772.10 | 71.761.72 | 72.301.79 | 85.591.55 | 84.391.57 | |
-| MTST Zhang et al. (2024) | 71.141.59 | 79.300.97 | 65.272.28 | 64.013.16 | 68.872.34 | 71.061.60 | |
-| Nonformer Liu et al. (2022) | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |
-| PatchTST Nie et al. (2023) | 67.031.65 | 78.761.28 | 59.912.02 | 55.973.10 | 65.650.28 | 67.990.76 | |
-| Reformer Kitaev et al. (2019) | 78.702.00 | 82.503.95 | 75.001.61 | 75.931.82 | 73.941.40 | 76.041.14 | |
-| Transformer Vaswani et al. (2017) | 76.304.72 | 77.645.95 | 73.095.01 | 73.755.38 | 72.506.60 | 73.237.60 | |
-| Medformer Wang et al. (2024) | 78.740.64 | 81.110.84 | 75.400.66 | 76.310.71 | 83.200.91 | 83.660.92 | |
-| MedGNN Fan et al. (2025) | 82.600.35 | 87.700.22 | 78.930.09 | 80.250.16 | 85.930.26 | - | |
-| HM-BiTCN + CIF | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | |
-| TDBrain (2-Classes) Reported | Autoformer Wu et al. (2021) | 87.333.79 | 88.063.56 | 87.333.79 | 87.263.84 | 93.812.26 | 93.322.42 |
-| Crossformer Zhang and Yan (2022) | 81.562.19 | 81.972.25 | 81.562.19 | 81.502.20 | 91.201.78 | 91.511.71 | |
-| FEDformer Zhou et al. (2022) | 78.131.98 | 78.521.91 | 78.131.98 | 78.042.01 | 86.561.86 | 86.481.99 | |
-| Informer Zhou et al. (2021) | 89.022.50 | 89.432.14 | 89.022.50 | 88.982.54 | 96.640.68 | 96.750.63 | |
-| iTransformer Liu et al. (2024b) | 74.671.06 | 74.711.06 | 74.671.06 | 74.651.06 | 83.371.14 | 83.731.27 | |
-| MTST Zhang et al. (2024) | 76.963.76 | 77.243.59 | 76.963.76 | 76.883.83 | 85.274.46 | 82.815.64 | |
-| Nonformer Liu et al. (2022) | 87.882.48 | 88.861.84 | 87.882.48 | 87.782.56 | 97.050.68 | 96.990.68 | |
-| PatchTST Nie et al. (2023) | 79.253.79 | 79.604.09 | 79.253.79 | 79.203.77 | 87.954.96 | 86.366.67 | |
-| Reformer Kitaev et al. (2019) | 87.922.01 | 88.641.40 | 87.922.01 | 87.852.08 | 96.300.54 | 96.400.45 | |
-| Transformer Vaswani et al. (2017) | 87.171.67 | 87.991.68 | 87.171.67 | 87.101.68 | 96.280.92 | 96.340.81 | |
-| Medformer Wang et al. (2024) | 89.620.81 | 89.680.78 | 89.620.81 | 89.620.81 | 96.410.35 | 96.510.33 | |
-| MedGNN Fan et al. (2025) | 91.040.09 | 91.150.12 | 91.040.20 | 91.040.08 | 96.740.04 | - | |
-| HM-BiTCN + CIF | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 | |
-| ADFTD (3-Classes) Reported | Autoformer Wu et al. (2021) | 45.251.48 | 43.671.94 | 42.962.03 | 42.591.85 | 61.021.82 | 43.102.30 |
-| Crossformer Zhang and Yan (2022) | 50.452.31 | 45.571.63 | 45.881.82 | 45.501.70 | 66.452.03 | 48.332.05 | |
-| FEDformer Zhou et al. (2022) | 46.300.59 | 46.050.76 | 44.221.38 | 43.911.37 | 62.621.75 | 46.111.44 | |
-| Informer Zhou et al. (2021) | 48.451.96 | 46.541.68 | 46.061.84 | 45.741.38 | 65.871.27 | 47.601.30 | |
-| iTransformer Liu et al. (2024b) | 52.601.59 | 46.791.27 | 47.281.29 | 46.791.13 | 67.261.16 | 49.531.21 | |
-| MTST Zhang et al. (2024) | 45.602.03 | 44.701.33 | 45.051.30 | 44.311.74 | 62.500.81 | 45.160.85 | |
-| Nonformer Liu et al. (2022) | 49.951.05 | 47.710.97 | 47.461.50 | 46.961.35 | 66.231.37 | 47.331.78 | |
-| PatchTST Nie et al. (2023) | 44.370.95 | 42.401.13 | 42.061.48 | 41.971.37 | 60.081.50 | 42.491.79 | |
-| Reformer Kitaev et al. (2019) | 50.781.17 | 49.641.49 | 49.891.67 | 47.940.69 | 69.171.58 | 51.731.94 | |
-| Transformer Vaswani et al. (2017) | 50.472.14 | 49.131.83 | 48.011.53 | 48.091.59 | 67.931.59 | 48.932.02 | |
-| Medformer Wang et al. (2024) | 53.271.54 | 51.021.57 | 50.711.55 | 50.651.51 | 70.931.19 | 51.211.32 | |
-| MedGNN Fan et al. (2025) | 56.120.11 | 55.070.09 | 55.470.34 | 55.000.24 | 74.680.33 | - | |
-| HM-BiTCN + CIF | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 | |
-| PTB (2-Classes) Reported | Autoformer Wu et al. (2021) | 73.352.10 | 72.112.89 | 63.243.17 | 63.693.84 | 78.543.48 | 74.253.53 |
-| Crossformer Zhang and Yan (2022) | 80.173.79 | 85.041.83 | 71.256.29 | 72.757.19 | 88.553.45 | 87.313.25 | |
-| FEDformer Zhou et al. (2022) | 76.052.54 | 77.583.61 | 66.103.55 | 67.144.37 | 85.934.31 | 82.595.42 | |
-| Informer Zhou et al. (2021) | 78.691.68 | 82.871.02 | 69.192.90 | 70.843.47 | 92.090.53 | 90.020.60 | |
-| iTransformer Liu et al. (2024b) | 83.890.71 | 88.251.18 | 76.391.01 | 79.061.06 | 91.181.16 | 90.930.98 | |
-| MTST Zhang et al. (2024) | 76.591.90 | 79.881.90 | 66.312.95 | 67.383.71 | 86.862.75 | 83.752.84 | |
-| Nonformer Liu et al. (2022) | 78.660.49 | 82.770.86 | 69.120.87 | 70.901.00 | 89.372.51 | 86.672.38 | |
-| PatchTST Nie et al. (2023) | 74.741.62 | 76.941.51 | 63.892.71 | 64.363.38 | 88.790.91 | 83.390.96 | |
-| Reformer Kitaev et al. (2019) | 77.962.13 | 81.721.61 | 68.203.35 | 69.653.88 | 91.130.74 | 88.421.30 | |
-| Transformer Vaswani et al. (2017) | 77.371.02 | 81.840.66 | 67.141.80 | 68.472.19 | 90.081.76 | 87.221.68 | |
-| Medformer Wang et al. (2024) | 83.502.01 | 85.190.94 | 77.113.39 | 79.183.31 | 92.811.48 | 90.321.54 | |
-| MedGNN Fan et al. (2025) | 84.530.28 | 87.350.45 | 77.900.66 | 80.400.62 | 93.310.46 | - | |
-| HM-BiTCN + CIF | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 | |
-| PTB-XL (5-Classes) Reported | Autoformer Wu et al. (2021) | 61.682.72 | 51.601.64 | 49.101.52 | 48.852.27 | 82.041.44 | 51.931.71 |
-| Crossformer Zhang and Yan (2022) | 73.300.14 | 65.060.35 | 61.230.33 | 62.590.14 | 90.020.06 | 67.430.22 | |
-| FEDformer Zhou et al. (2022) | 57.209.47 | 52.386.09 | 49.047.26 | 47.898.44 | 82.134.17 | 52.317.03 | |
-| Informer Zhou et al. (2021) | 71.430.32 | 62.640.60 | 59.120.47 | 60.440.43 | 88.650.09 | 64.760.17 | |
-| iTransformer Liu et al. (2024b) | 69.280.22 | 59.590.45 | 54.620.18 | 56.200.19 | 86.710.10 | 60.270.21 | |
-| MTST Zhang et al. (2024) | 72.140.27 | 63.840.72 | 60.010.81 | 61.430.38 | 88.970.33 | 65.830.51 | |
-| Nonformer Liu et al. (2022) | 70.560.55 | 61.570.66 | 57.750.72 | 59.100.66 | 88.320.36 | 63.400.79 | |
-| PatchTST Nie et al. (2023) | 73.230.25 | 65.700.64 | 60.820.76 | 62.610.34 | 89.740.19 | 67.320.22 | |
-| Reformer Kitaev et al. (2019) | 71.720.43 | 63.121.02 | 59.200.75 | 60.690.18 | 88.800.24 | 64.720.47 | |
-| Transformer Vaswani et al. (2017) | 70.590.44 | 61.570.65 | 57.620.35 | 59.050.25 | 88.210.16 | 63.360.29 | |
-| Medformer Wang et al. (2024) | 72.870.23 | 64.140.42 | 60.600.46 | 62.020.37 | 89.660.13 | 66.390.22 | |
-| MedGNN Fan et al. (2025) | 73.870.18 | 66.260.29 | 61.130.29 | 62.540.20 | 90.210.15 | - | |
-| HM-BiTCN + CIF | 73.730.30 | 65.410.67 | 60.701.08 | 61.890.91 | 90.530.22 | 67.750.75 |
 4.2.1 Efficiency Analysis
@@
 4.3 Ablation Study
-(1) Effectiveness of CIF: Table 3 demonstrates the excellent performance of combining HM-BiTCN with CIF, confirming the compatibility of the HM-BiTCN with CIF. See Appendix F.1 for details.
+(1) Effectiveness of CIF: Table 3 demonstrates the excellent performance of combining HM-BiTCN with CIF, confirming the compatibility of the HM-BiTCN with CIF.
+Appendix D presents ablation studies on the HM-BiTCN architecture and the performance benefits of integrating CIF into its components.
 | Datasets | APAVA | ADFTD | PTB | TDBRAIN | ||||
@@
 If transferred settings underperform, we further fine-tune them for adaptation.
-See Appendices F.2 and F.3.
 Figure 5 illustrates the outstanding performance of CIF when combined with other models.
+(3) Exploring Physiologically-Informed CIF: In Appendix E, we further conducted experiments on CIF from the perspective of emphasizing more prominent biological features.
+The results indicate that incorporating specific biological characteristics can further enhance classification performance.
+(4)Results on general time series classification tasks
+Accuracy | 75.840.52 | 71.651.27 | 74.961.25 | 66.481.66 | 74.800.96 | 64.872.78 | 76.440.64 | 76.080.81 | 76.821.32 | |||||||||||||||||||
+| F1 Score | 75.520.66 | 71.141.45 | 74.491.39 | 65.291.74 | 74.350.85 | 64.142.70 | 76.250.65 | 75.540.94 | 76.391.18 | ||||||||||||||||||||
+Accuracy | 89.741.08 | 88.442.02 | 88.861.65 | 93.080.95 | 91.441.01 | 87.781.10 | 91.650.74 | 93.720.73 | 93.780.32 | |||||||||||||||||||
+| F1 Score | 89.701.10 | 88.341.98 | 88.801.67 | 93.190.88 | 91.470.98 | 87.721.10 | 91.610.75 | 93.690.76 | 93.740.34 | ||||||||||||||||||||
+To evaluate the performance of our method on general time series, we follow the design of Medformer Wang et al.
+(2024a) and test it on two human activity recognition (HAR) datasets: FLAAP(13,123 samples, 10 classes) Kumar & Suresh (2022) and UCI-HAR(10,299 samples, 6 classes) Anguita et al.
+(2013).
+Additionally, to conduct a more comprehensive evaluation, following TimeMixer++ Wang et al.
+(2025a), we used 10 multivariate datasets from the UEA Time Series Classification Archive (2018) for the assessment of classification tasks.
+As shown in Table 4, and Fig.
+6, the combination of HM-BiTCN and CIF consistently outperforms other architectures in general time series classification, achieving a 5.5% improvement over the original TCN and surpassing current SOTA methods.
+Although CIF was originally designed for MedTS, its integration with HM-BiTCN significantly outperforms Transformer-based models in both medical and general time series classification tasks, demonstrating the effectiveness of our data-centric approach.
+The results in Tables 4, 7, 3 and Figure 5 show that CIF achieves significant improvements in MedTS classification, while the gains on non-medical data are relatively limited.
+This observation further demonstrates that a data-driven perspective is particularly effective for MedTS classification with physiological characteristics.
 5 Conclusion
-In this paper, we propose a simple yet effective temporal classification representation method called Channel-Imposed Fusion (CIF) and combine it with a straightforward TCN structure.
-This approach outperforms current SOTA methods on multiple medical time-series datasets.
-Additionally, the CIF method demonstrates strong adaptability, enabling integration with cutting-edge methods to further enhance their performance.
-Through the CIF method, we aim to show researchers in the field that, rather than solely modifying models to process high-level features, exploring the underlying core features of the data itself is also a research path worth delving into.
-Appendix A Proof of SNR Optimization via CIF.
-Consider two observed signals:
-where:
-are zero-mean useful signals with variance and correlation coefficient
-are zero-mean noise components with variance and correlation coefficient
-Signal and noise components are mutually uncorrelated
-Derivation of Output SNR
-For the linear combination :
+In this work, we propose a simple and effective method for medical time series classification, Channel-Imposed Fusion (CIF), which explicitly encodes physiological causal relationships between channels in the feature representations while enhancing the SNR of the original signals.
+Combined with the simple HM-BiTCN architecture, CIF surpasses existing SOTA methods on multiple medical datasets and performs strongly on general time series classification tasks, demonstrating that data-centric design enables simple models to outperform more complex architectures.
+More importantly, CIF exemplifies the shift from the traditional model-centric paradigm to a data-centric perspective, where structured representations grounded in physiological priors are both efficient and scalable for medical time series classification.
+CIF also exhibits strong transferability and can be seamlessly integrated into mainstream models such as Transformers, enhancing their applicability in medical scenarios.
+We hope this work encourages the community to reconsider the core of medical time series classification: should it be driven primarily by data-centric strategies or by model-centric design or both?
+Appendix
+Appendix A Explanation of SNR Optimization via CIF.
+Consider the linear combination of the two observed signals:
+where and are real coefficients. The observed signals are given by
+with zero-mean signal and noise components:
+and mutually uncorrelated signal and noise components: .
 1. Signal Power Calculation
+The power of a zero-mean random signal is given by its variance:
+This is why, for zero-mean signals, the SNR can be expressed as a ratio of variances (or mean-square values) Kay (1993).
+For the linear combination of signals:
+where .
 2. Noise Power Calculation
-3. SNR Ratio
-The output SNR becomes:
+Similarly, the noise power of the linear combination is:
 where .
-Optimization Conditions
-SNR improvement occurs when:
-Difference Mode ():
-Cooperative Mode ():
-Appendix B Visualization of the Channel-Imposed Fusion (CIF) Process.
-In Figure 5 6, we visualize scenarios with highly correlated channel information and those with generally correlated channel information.
-These visualizations intuitively demonstrate the strong representation capability of our proposed method.
-Appendix C Data Preprocessing
-| Datasets | #-Subject | #-Sample | #-Class | #-Channel | #-Timestamps | Sampling Rate | Modality | File Size |
-| APAVA | 23 | 5,967 | 2 | 16 | 256 | 256Hz | EEG | 186MB |
-| ADFTD | 88 | 69,752 | 3 | 19 | 256 | 256Hz | EEG | 2.52GB |
-| TDBrain | 72 | 6,240 | 2 | 33 | 256 | 256Hz | EEG | 571MB |
-| PTB | 198 | 64,356 | 2 | 15 | 300 | 250Hz | ECG | 2.15GB |
-| PTB-XL | 17,596 | 191,400 | 5 | 12 | 250 | 250Hz | ECG | 4.28GB |
-Appendix D Implementation Details
-We implement our method and all the baselines based on the Time-Series-Library project111https://github.com/thuml/Time-Series-Library from Tsinghua University Wu et al.
-[2023], which integrates all methods under the same framework and training techniques to ensure a relatively fair comparison.
-The 12 baseline time series transformer methods are Autoformer Wu et al.
-[2021], Crossformer Zhang and Yan [2022], FEDformer Zhou et al.
-[2022], Informer Zhou et al.
-[2021], iTransformer Liu et al.
-[2024b], MTST Zhang et al.
-[2024], Nonformer Liu et al.
-[2022], PatchTST Nie et al.
-[2023], Reformer Kitaev et al.
-[2019], vanilla Transformer Vaswani et al.
-[2017], Medformer Wang et al.
-[2024], and MedGNN Fan et al.
-[2025].
-Autoformer Autoformer Wu et al.
-[2021] employs an auto-correlation mechanism to replace self-attention for time series forecasting.
-Additionally, they use a time series decomposition block to separate the time series into trend-cyclical and seasonal components for improved learning.
-The raw source code is available at https://github.com/thuml/Autoformer.
-Crossformer Crossformer Zhang and Yan [2022] designs a single-channel patching approach for token embedding.
-They utilize two-stage self-attention to leverage both temporal features and channel correlations.
-A router mechanism is proposed to reduce time and space complexity during the cross-dimension stage.
-The raw code is available at https://github.com/Thinklab-SJTU/Crossformer.
-FEDformer FEDformer Zhou et al.
-[2022] leverages frequency domain information using the Fourier transform.
-They introduce frequency-enhanced blocks and frequency-enhanced attention, which are computed in the frequency domain.
-A novel time series decomposition method replaces the layer norm module in the transformer architecture to improve learning.
-The raw code is available at https://github.com/MAZiqing/FEDformer.
-Informer Informer Zhou et al.
-[2021] is the first paper to employ a one-forward procedure instead of an autoregressive method in time series forecasting tasks.
-They introduce ProbSparse self-attention to reduce complexity and memory usage.
-The raw code is available at https://github.com/zhouhaoyi/Informer2020.
-iTransformer iTransformer Liu et al.
-[2024b] questions the conventional approach of embedding attention tokens in time series forecasting tasks and proposes an inverted approach by embedding the whole series of channels into a token.
-They also invert the dimension of other transformer modules, such as the layer norm and feed-forward networks.
-The raw code is available at https://github.com/thuml/iTransformer.
-MTST MTST Zhang et al.
-[2024] uses the same token embedding method as Crossformer and PatchTST.
-It highlights the importance of different patching lengths in forecasting tasks and designs a method that can take different sizes of patch tokens as input simultaneously.
-The raw code is available at https://github.com/networkslab/MTST.
-Nonformer Nonformer Liu et al.
-[2022] analyzes the impact of non-stationarity in time series forecasting tasks and its significant effect on results.
-They design a de-stationary attention module and incorporate normalization and denormalization steps before and after training to alleviate the over-stationarization problem.
-The raw code is available at https://github.com/thuml/Nonstationary_Transformers.
-PatchTST PatchTST Nie et al.
-[2023] embeds a sequence of single-channel timestamps as a patch token to replace the attention token used in the vanilla transformer.
-This approach enlarges the receptive field and enhances forecasting ability.
-The raw code is available at https://github.com/yuqinie98/PatchTST.
-Reformer Reformer Kitaev et al.
-[2019] replaces dot-product attention with locality-sensitive hashing.
-They also use a reversible residual layer instead of standard residuals.
-The raw code is available at https://github.com/lucidrains/reformer-pytorch.
-Transformer Transformer Vaswani et al.
-[2017], commonly known as the vanilla transformer, is introduced in the well-known paper "Attention is All You Need." It can also be applied to time series by embedding each timestamp of all channels as an attention token.
-The PyTorch version of the code is available at https://github.com/jadore801120/attention-is-all-you-need-pytorch.
-Medformer Medformer Wang et al. [2024] uses cross-channel patch embedding to model spatiotemporal dependencies.The raw code is available at https://github.com/DL4mHealth/Medformer
-MedGNN MedGNN Fan et al.
-[2025] employs multi-resolution spatiotemporal graph learning to extract dynamic features across multiple time scales.
-The raw code is available at https://github.com/aikunyi/MedGNN.
-D.1 Evaluation metrics
+3. Output SNR
+Using the definition of SNR as the ratio of signal power to noise power:
+where .
+4. SNR Improvement Condition
+For SNR improvement relative to individual channels:
+Difference Mode (): — suppress correlated noise while possibly attenuating some correlated signal.
+Cooperative Mode (): — amplify correlated signals relative to less-correlated noise.
+A.1 Evaluation metrics
 For all methods, the optimizer used is Adam, with a learning rate of 1e-4.
@@
 It gives a single metric that reflects both precision and recall performance.
-Appendix E Reproduced Results of Existing Methods
-We employ six evaluation metrics: accuracy, precision (macro-averaged), recall (macro-averaged), F1 score (macro-averaged), AUROC (macro-averaged), and AUPRC (macro-averaged).
-The training process is conducted with five random seeds (41-45) on fixed training, validation, and test sets to compute the mean and standard deviation of the models.
-All experiments are run on an NVIDIA RTX 3090 GPU.
-E.1 Results of Subject-Dependent
-In Table 5, we present the results of various models under the Subject-Dependent Setup on the ADFTD (3-Classes) dataset.
-The results indicate that our proposed CIF method significantly improves the performance of models, particularly in accuracy, F1 score, AUROC, and AUPRC metrics.
-Specifically, the combination of MedGNN + CIF outperforms MedGNN alone across all key performance metrics, and similarly, Medformer + CIF surpasses Medformer in every critical evaluation metric.
-These results convincingly demonstrate the effectiveness of the CIF method in enhancing model performance.
-By integrating the CIF method, not only are the models’ performances significantly improved across multiple evaluation metrics, but the broad applicability and substantial benefits of the method across different model architectures are also highlighted.
-Notably, MedGNN + CIF achieves a remarkable Accuracy of 99.60%, far surpassing other models, highlighting the advantage of our approach in handling imbalanced datasets.
-Furthermore, MedGNN + CIF shows significant improvements in F1 score and precision, indicating that the CIF method effectively enhances the model’s ability to recognize both positive and negative samples.
-Thus, our experimental results validate the effectiveness of the CIF method in boosting model generalization and precision.
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| ADFTD (3-Classes) Reproduced | Autoformer Wu et al. [2021] | 87.792.87 | 87.632.95 | 86.843.10 | 87.173.05 | 96.451.15 | 93.522.05 |
-| Crossformer Zhang and Yan [2022] | 89.271.19 | 88.981.27 | 88.611.35 | 88.771.29 | 97.530.48 | 95.470.87 | |
-| FEDformer Zhou et al. [2022] | 77.552.42 | 76.712.14 | 76.462.52 | 76.442.52 | 91.661.36 | 84.992.15 | |
-| Informer Zhou et al. [2021] | 91.720.13 | 91.500.16 | 91.360.21 | 91.420.14 | 98.400.07 | 96.920.11 | |
-| iTransformer Liu et al. [2024b] | 64.600.77 | 62.220.64 | 61.950.35 | 61.870.41 | 81.460.29 | 68.810.39 | |
-| MTST Zhang et al. [2024] | 65.430.79 | 64.470.65 | 63.490.55 | 63.710.55 | 81.670.51 | 69.740.69 | |
-| Nonformer Liu et al. [2022] | 96.000.72 | 95.720.80 | 95.900.69 | 95.810.75 | 99.560.13 | 99.020.23 | |
-| PatchTST Nie et al. [2023] | 66.550.48 | 65.610.60 | 65.260.78 | 65.260.60 | 83.270.33 | 71.960.44 | |
-| Reformer Kitaev et al. [2019] | 90.742.18 | 90.872.15 | 90.612.01 | 90.442.25 | 98.690.50 | 97.570.85 | |
-| Transformer Vaswani et al. [2017] | 96.620.54 | 96.460.54 | 96.450.59 | 96.450.56 | 99.690.08 | 99.310.13 | |
-| Medformer Wang et al. [2024] | 95.761.03 | 95.541.10 | 95.621.05 | 95.581.08 | 99.530.18 | 99.090.36 | |
-| Medformer + CIF | 98.870.26 | 98.770.27 | 98.860.27 | 98.810.27 | 99.960.01 | 99.920.03 | |
-| MedGNN Fan et al. [2025] | 99.440.21 | 99.430.21 | 99.410.22 | 99.420.21 | 99.980.01 | 99.970.01 | |
-| MedGNN + CIF | 99.600.09 | 99.600.11 | 99.580.09 | 99.590.10 | 99.990.01 | 99.970.01 |
-E.2 Results of Subject-Independent
-Table 2 presents the results of various methods reported under the subject-independent setup, while Table 6 shows the results of our reproduction of these methods.
-Our approach achieves the highest scores on six metrics across four out of the five datasets.
-For the PTB-XL dataset, compared to the results reported by other methods, our method achieves the highest AUROC, with Accuracy ranking second.
-In comparison with our reproduced results, our method ranks first in Accuracy, AUROC, and AUPRC, and second in Precision and Recall.
-Additionally, it is noteworthy that under the subject-independent setup, the F1 score for ADFTD is 55.42%, significantly lower than the 99.59% F1 score achieved under the subject-dependent setup.
-This comparison highlights the challenge of the subject-independent setup, which better simulates real-world scenarios.
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| APAVA (2-Classes) Reproduced | Autoformer Wu et al. [2021] | 73.187.33 | 73.876.72 | 73.016.10 | 72.407.03 | 81.647.24 | 81.107.75 |
-| Crossformer Zhang and Yan [2022] | 72.762.04 | 79.642.45 | 67.412.62 | 66.883.61 | 71.814.06 | 71.643.74 | |
-| FEDformer Zhou et al. [2022] | 75.161.67 | 74.980.69 | 73.342.97 | 73.502.90 | 83.891.54 | 83.271.62 | |
-| Informer Zhou et al. [2021] | 72.202.78 | 73.924.80 | 68.482.51 | 68.742.70 | 70.143.43 | 70.843.80 | |
-| iTransformer Liu et al. [2024b] | 74.551.66 | 74.772.10 | 71.761.72 | 72.301.79 | 85.591.55 | 84.391.57 | |
-| MTST Zhang et al. [2024] | 69.241.24 | 75.872.80 | 63.281.81 | 61.622.75 | 66.093.27 | 68.082.93 | |
-| Nonformer Liu et al. [2022] | 71.814.20 | 71.314.40 | 70.153.38 | 70.383.74 | 71.542.73 | 72.792.50 | |
-| PatchTST Nie et al. [2023] | 68.272.11 | 78.561.88 | 61.532.60 | 58.524.07 | 64.612.18 | 67.142.06 | |
-| Reformer Kitaev et al. [2019] | 78.422.85 | 80.894.52 | 75.202.28 | 76.092.54 | 75.482.79 | 77.522.64 | |
-| Transformer Vaswani et al. [2017] | 75.534.28 | 76.905.05 | 72.144.87 | 72.645.44 | 72.306.04 | 73.047.15 | |
-| Medformer Wang et al. [2024] | 77.852.42 | 80.313.21 | 74.382.49 | 75.212.67 | 80.853.80 | 81.623.24 | |
-| MedGNN Fan et al. [2025] | 77.405.77 | 82.774.46 | 73.247.06 | 73.299.01 | 81.312.94 | 82.802.91 | |
-| HM-BiTCN + CIF | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | |
-| TDBrain (2-Classes) Reproduced | Autoformer Wu et al. [2021] | 90.383.03 | 91.162.42 | 90.383.03 | 90.313.09 | 95.832.14 | 95.432.31 |
-| Crossformer Zhang and Yan [2022] | 82.152.60 | 82.812.11 | 82.152.60 | 82.042.70 | 91.202.23 | 91.472.16 | |
-| FEDformer Zhou et al. [2022] | 77.601.23 | 78.251.52 | 77.601.23 | 77.481.19 | 86.311.23 | 86.481.36 | |
-| Informer Zhou et al. [2021] | 88.422.99 | 89.012.45 | 88.422.99 | 88.363.05 | 96.540.90 | 96.660.85 | |
-| iTransformer Liu et al. [2024b] | 74.691.02 | 74.761.04 | 74.691.02 | 74.671.02 | 83.351.24 | 83.651.41 | |
-| MTST Zhang et al. [2024] | 77.673.58 | 78.974.37 | 77.673.58 | 77.453.55 | 86.474.84 | 84.996.43 | |
-| Nonformer Liu et al. [2022] | 88.102.39 | 88.761.74 | 88.102.39 | 88.042.47 | 96.560.91 | 96.361.21 | |
-| PatchTST Nie et al. [2023] | 77.982.64 | 79.303.73 | 77.982.64 | 77.762.65 | 86.674.03 | 84.935.47 | |
-| Reformer Kitaev et al. [2019] | 88.502.30 | 89.011.80 | 88.502.30 | 88.452.35 | 96.100.63 | 96.190.55 | |
-| Transformer Vaswani et al. [2017] | 85.131.86 | 86.391.56 | 85.131.86 | 84.991.93 | 95.611.05 | 95.630.91 | |
-| Medformer Wang et al. [2024] | 88.771.24 | 88.911.11 | 88.771.24 | 88.761.25 | 96.380.34 | 96.440.30 | |
-| MedGNN Fan et al. [2025] | - | - | - | - | - | - | |
-| HM-BiTCN + CIF | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 | |
-| ADFTD (3-Classes) Reproduced | Autoformer Wu et al. [2021] | 46.902.89 | 45.592.37 | 44.912.23 | 44.342.52 | 63.492.44 | 45.632.29 |
-| Crossformer Zhang and Yan [2022] | 50.181.97 | 45.971.84 | 46.301.73 | 45.901.84 | 66.681.67 | 48.651.89 | |
-| FEDformer Zhou et al. [2022] | 45.750.78 | 45.711.29 | 44.271.28 | 43.511.00 | 62.641.64 | 45.881.35 | |
-| Informer Zhou et al. [2021] | 48.421.99 | 46.941.60 | 46.410.99 | 45.760.43 | 65.991.14 | 47.491.07 | |
-| iTransformer Liu et al. [2024b] | 52.851.36 | 46.971.05 | 47.311.03 | 46.840.78 | 67.460.96 | 49.900.89 | |
-| MTST Zhang et al. [2024] | 45.771.70 | 44.391.73 | 43.701.82 | 43.361.98 | 61.381.57 | 44.011.60 | |
-| Nonformer Liu et al. [2022] | 50.811.06 | 48.711.40 | 48.551.47 | 48.361.38 | 66.951.54 | 48.081.82 | |
-| PatchTST Nie et al. [2023] | 43.320.53 | 41.950.38 | 41.451.26 | 40.751.62 | 60.210.29 | 42.490.57 | |
-| Reformer Kitaev et al. [2019] | 51.282.60 | 49.682.75 | 49.642.02 | 48.452.06 | 69.202.53 | 51.743.24 | |
-| Transformer Vaswani et al. [2017] | 50.530.94 | 49.310.87 | 48.571.23 | 48.421.28 | 67.980.90 | 49.071.35 | |
-| Medformer Wang et al. [2024] | 53.701.18 | 51.511.32 | 50.491.48 | 50.351.53 | 70.481.17 | 50.911.13 | |
-| MedGNN Fan et al. [2025] | 50.223.21 | 48.653.72 | 47.504.57 | 47.334.40 | 67.184.39 | 48.844.11 | |
-| HM-BiTCN + CIF | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 | |
-| PTB (2-Classes) Reproduced | Autoformer Wu et al. [2021] | 71.992.74 | 69.603.85 | 61.504.23 | 61.435.07 | 74.291.89 | 70.262.00 |
-| Crossformer Zhang and Yan [2022] | 78.063.44 | 81.533.13 | 68.625.63 | 69.766.53 | 88.312.07 | 85.812.43 | |
-| FEDformer Zhou et al. [2022] | 74.542.27 | 77.994.10 | 63.143.29 | 63.284.36 | 84.634.27 | 80.915.55 | |
-| Informer Zhou et al. [2021] | 79.590.65 | 83.330.77 | 70.580.95 | 72.581.08 | 92.770.48 | 90.890.57 | |
-| iTransformer Liu et al. [2024b] | 83.431.19 | 88.061.47 | 75.641.55 | 78.291.70 | 91.381.41 | 91.081.30 | |
-| MTST Zhang et al. [2024] | 75.532.45 | 78.721.87 | 64.784.06 | 65.304.81 | 87.764.09 | 83.603.92 | |
-| Nonformer Liu et al. [2022] | 78.931.46 | 82.481.53 | 69.682.15 | 71.502.56 | 90.540.59 | 87.781.46 | |
-| PatchTST Nie et al. [2023] | 75.282.44 | 77.052.44 | 64.864.05 | 65.415.28 | 88.112.59 | 82.652.87 | |
-| Reformer Kitaev et al. [2019] | 78.111.65 | 82.700.80 | 68.172.68 | 69.683.34 | 90.771.56 | 88.141.20 | |
-| Transformer Vaswani et al. [2017] | 76.431.98 | 81.251.15 | 65.643.29 | 66.444.39 | 90.211.24 | 87.281.49 | |
-| Medformer Wang et al. [2024] | 80.990.75 | 83.010.72 | 73.351.16 | 75.471.21 | 93.101.18 | 90.691.04 | |
-| MedGNN Fan et al. [2025] | - | - | - | - | - | - | |
-| HM-BiTCN + CIF | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 | |
-| PTB-XL (5-Classes) Reproduced | Autoformer Wu et al. [2021] | 60.371.72 | 49.830.50 | 49.170.72 | 48.430.47 | 81.350.56 | 50.830.72 |
-| Crossformer Zhang and Yan [2022] | 73.150.17 | 64.620.31 | 61.280.45 | 62.490.34 | 89.940.15 | 67.140.29 | |
-| FEDformer Zhou et al. [2022] | 56.0110.27 | 51.615.58 | 48.777.64 | 47.238.76 | 82.014.20 | 52.237.14 | |
-| Informer Zhou et al. [2021] | 71.280.27 | 62.280.53 | 59.180.54 | 60.380.38 | 88.570.09 | 64.570.18 | |
-| iTransformer Liu et al. [2024b] | 69.180.33 | 59.440.48 | 54.840.25 | 56.440.23 | 86.650.17 | 60.270.44 | |
-| MTST Zhang et al. [2024] | 72.140.29 | 64.010.61 | 59.310.27 | 61.070.32 | 88.820.19 | 65.520.52 | |
-| Nonformer Liu et al. [2022] | 70.510.79 | 61.421.03 | 58.150.57 | 59.310.63 | 88.200.43 | 63.330.77 | |
-| PatchTST Nie et al. [2023] | 73.130.20 | 65.380.54 | 60.610.56 | 62.390.27 | 89.660.16 | 66.970.24 | |
-| Reformer Kitaev et al. [2019] | 71.240.46 | 61.780.79 | 59.670.99 | 60.510.65 | 88.670.21 | 64.290.29 | |
-| Transformer Vaswani et al. [2017] | 70.460.42 | 61.560.58 | 57.860.65 | 59.160.46 | 88.180.14 | 63.250.27 | |
-| Medformer Wang et al. [2024] | 72.940.15 | 64.390.32 | 59.980.49 | 61.610.33 | 89.670.10 | 66.250.25 | |
-| MedGNN Fan et al. [2025] | - | - | - | - | - | - | |
-| HM-BiTCN + CIF | 73.730.30 | 65.410.67 | 60.701.08 | 61.890.91 | 90.530.22 | 67.750.75 |
-Appendix F Ablation Study
-F.1 Module Study of CIF
-We conduct experiments by integrating our Channel-Imposed Fusion principle with other methods.
-In the CIF structure, there are four different hyperparameters.
-First, is the switch for forward and backward feature fusion.
-When , it replaces the forward features with the fused features; when , it replaces the backward features with the fused features.
-The second hyperparameter, , determines the number of selected channels.
-Additionally, there are two scaling factors associated with the channels: , the scaling factor for the first channels, and , the scaling factor for the remaining channels.
-Both and can either be learnable or fixed parameters.
-We use
-to represent learnable parameters, and
-to represent non-learnable parameters.
-The ablation study(Table 7 8 9) reveals that the performance of the CIF architecture is highly sensitive to its hyperparameters, with optimal configurations varying significantly across datasets.
-For the APAVA dataset (16 channels), the best results (85.16% accuracy, 94.06% AUROC) are achieved using backward fusion () on 9 selected channels (), where the first 7 channels apply fixed reinforcement weights (, non-learnable) and the rest use fixed suppression weights (, non-learnable).
-Conversely, the ADFTD dataset (19 channels) performs best (58.56% accuracy, 76.07% AUROC) with forward fusion () over 10 selected channels () using dynamically learned weights for negative suppression () and positive enhancement ().
-These differences, also observed in PTB (APAVA: 78–85% accuracy vs.
-ADFTD: 50–58%), emphasize the critical role of dataset-specific tuning.
-In particular, the number of selected channels () and fusion direction () are pivotal, while adaptive weighting () further improves robustness.
-Overall, the results highlight the necessity of careful hyperparameter optimization—especially in channel selection and fusion strategy—for maximizing CIF performance across diverse medical time series tasks.
-| t | n | a | b | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| - | - | - | - | 52.052.22 | 50.453.00 | 50.402.55 | 49.482.70 | 69.432.84 | 50.993.15 |
-| 1 | 10 | -0.20 |
--0.30 |
-57.330.95 | 54.061.47 | 54.441.33 | 53.731.47 | 74.821.36 | 58.241.78 |
-| 1 | 10 | -0.25 |
--0.40 |
-57.081.46 | 53.310.98 | 53.241.14 | 52.870.98 | 73.771.28 | 56.861.47 |
-| 1 | 11 | -0.20 |
--0.30 |
-56.681.30 | 53.042.00 | 53.452.03 | 52.842.30 | 73.811.55 | 57.341.46 |
-| 1 | 10 | -0.19 |
--0.27 |
-58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 |
-| 1 | 7 | 0.20 |
--0.60 |
-56.360.58 | 53.911.10 | 54.341.03 | 53.780.91 | 74.190.69 | 56.621.28 |
-| 1 | 9 | -0.20 |
--0.60 |
-55.201.50 | 53.051.14 | 53.211.06 | 52.761.02 | 73.091.13 | 55.501.71 |
-| 1 | 10 | -0.20 |
--0.20 |
-57.110.92 | 55.041.43 | 55.531.57 | 54.581.29 | 75.131.33 | 58.471.43 |
-| 1 | 11 | -0.20 |
--0.30 |
-57.831.16 | 54.321.22 | 54.511.18 | 54.011.21 | 74.741.12 | 57.961.08 |
-| 1 | 11 | -0.18 |
--0.25 |
-57.001.32 | 53.002.28 | 53.171.94 | 52.322.19 | 73.971.89 | 57.001.93 |
-| t | n | a | b | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| - | - | - | - | 81.871.87 | 86.501.24 | 73.492.90 | 75.843.20 | 94.200.29 | 93.040.45 |
-| 1 | 8 | 0.21 |
--0.50 |
-88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 |
-| 1 | 8 | 0.20 |
--0.50 |
-84.452.15 | 88.341.15 | 77.403.52 | 79.893.51 | 93.150.70 | 91.940.67 |
-| 1 | 9 | 0.22 |
--0.50 |
-85.140.22 | 87.421.08 | 79.050.52 | 81.420.35 | 93.821.19 | 92.731.27 |
-| 1 | 10 | 0.22 |
--0.50 |
-85.402.08 | 88.171.96 | 79.173.12 | 81.573.07 | 94.571.13 | 93.531.44 |
-| 1 | 11 | 0.22 |
--0.50 |
-84.470.84 | 88.020.59 | 77.501.26 | 80.101.28 | 92.770.89 | 92.020.78 |
-| 1 | 8 | 0.23 |
--0.50 |
-83.965.41 | 87.133.96 | 76.898.22 | 78.669.51 | 92.492.91 | 91.143.79 |
-| 1 | 8 | 0.25 |
--0.50 |
-85.562.09 | 88.961.77 | 79.033.03 | 81.613.14 | 93.611.11 | 93.071.03 |
-| t | n | a | b | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| - | - | - | - | 82.491.40 | 82.381.79 | 81.201.32 | 81.601.39 | 91.101.63 | 91.301.71 |
-| -1 | 5 | 1 |
-1 |
-81.481.96 | 81.992.38 | 79.522.00 | 80.212.07 | 90.931.74 | 91.171.77 |
-| -1 | 6 | 1 |
-1 |
-84.002.37 | 84.332.61 | 82.372.44 | 83.022.50 | 92.672.57 | 92.872.55 |
-| -1 | 7 | 1 |
-1 |
-81.772.31 | 82.673.24 | 79.782.04 | 80.492.26 | 89.922.60 | 90.232.53 |
-| 1 | 9 | 1 |
-1 |
-85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 |
-| -1 | 10 | 1 |
-1 |
-82.601.82 | 82.302.04 | 81.922.29 | 81.921.95 | 91.941.55 | 92.161.50 |
-| -1 | 11 | 1 |
-2 |
-78.980.97 | 78.571.09 | 77.600.93 | 77.930.97 | 87.290.93 | 87.350.96 |
-F.2 Generalizability of Hyperparameters
-We transfer the key hyperparameters obtained from HM-BiTCN—such as the fusion coefficients (, ), and segment length ()—to other baseline models.
-This experiment evaluates whether these hyperparameters retain effectiveness when applied to different architectures, indicating their robustness and broader applicability.
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| APAVA (2-Classes) Reproduced | Autoformer | 73.187.33 | 73.876.72 | 73.016.10 | 72.407.03 | 81.647.24 | 81.107.75 |
-| Autoformer + CIF | 75.962.68 | 76.333.47 | 75.111.10 | 74.971.90 | 83.421.99 | 83.152.63 | |
-| Crossformer | 72.762.04 | 79.642.45 | 67.412.62 | 66.883.61 | 71.814.06 | 71.643.74 | |
-| Crossformer + CIF | 82.322.60 | 85.351.83 | 79.213.17 | 80.293.30 | 90.391.58 | 90.021.47 | |
-| FEDformer | 75.161.67 | 74.980.69 | 73.342.97 | 73.502.90 | 83.891.54 | 83.271.62 | |
-| FEDformer + CIF | 77.202.17 | 76.972.07 | 77.022.60 | 76.552.40 | 86.701.73 | 86.531.76 | |
-| Informer | 72.202.78 | 73.924.80 | 68.482.51 | 68.742.70 | 70.143.43 | 70.843.80 | |
-| Informer + CIF | 79.782.07 | 82.293.03 | 76.552.05 | 77.532.23 | 78.561.33 | 78.581.14 | |
-| iTransformer | 74.551.66 | 74.772.10 | 71.761.72 | 72.301.79 | 85.591.55 | 84.391.57 | |
-| iTransformer + CIF | 74.950.87 | 74.400.75 | 73.791.78 | 73.811.43 | 84.300.88 | 82.490.96 | |
-| MTST | 69.241.24 | 75.872.80 | 63.281.81 | 61.622.75 | 66.093.27 | 68.082.93 | |
-| MTST + CIF | 76.202.39 | 81.461.11 | 71.673.06 | 72.063.73 | 77.653.37 | 77.982.73 | |
-| Nonformer | 71.814.20 | 71.314.40 | 70.153.38 | 70.383.74 | 71.542.73 | 72.792.50 | |
-| Nonformer + CIF | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |
-| Reformer | 78.422.85 | 80.894.52 | 75.202.28 | 76.092.54 | 75.482.79 | 77.522.64 | |
-| Reformer + CIF | 81.510.57 | 84.900.86 | 78.180.60 | 79.320.64 | 79.102.50 | 80.772.21 | |
-| Transformer | 75.534.28 | 76.905.05 | 72.144.87 | 72.645.44 | 72.306.04 | 73.047.15 | |
-| Transformer + CIF | 77.962.82 | 79.343.86 | 75.072.52 | 75.872.73 | 74.751.62 | 74.762.33 | |
-| Medformer | 77.852.42 | 80.313.21 | 74.382.49 | 75.212.67 | 80.853.80 | 81.623.24 | |
-| Medformer + CIF | 81.061.58 | 82.972.23 | 78.261.52 | 79.231.65 | 85.741.85 | 86.321.48 | |
-| HM-BiTCN | 82.491.40 | 82.381.79 | 81.201.32 | 81.601.39 | 91.101.63 | 91.301.71 | |
-| HM-BiTCN + CIF | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 |
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| TDBrain (2-Classes) Reproduced | Autoformer | 90.383.03 | 91.162.42 | 90.383.03 | 90.313.09 | 95.832.14 | 95.432.31 |
-| Autoformer + CIF | 93.422.49 | 93.712.27 | 93.422.49 | 93.402.51 | 97.461.08 | 97.211.15 | |
-| Crossformer | 82.152.60 | 82.812.11 | 82.152.60 | 82.042.70 | 91.202.23 | 91.472.16 | |
-| Crossformer + CIF | 89.401.26 | 89.831.18 | 89.401.26 | 89.371.27 | 96.760.62 | 96.800.62 | |
-| FEDformer | 77.601.23 | 78.251.52 | 77.601.23 | 77.481.19 | 86.311.23 | 86.481.36 | |
-| FEDformer + CIF | 78.871.94 | 79.192.00 | 78.881.94 | 78.821.94 | 88.121.87 | 88.391.92 | |
-| Informer | 88.422.99 | 89.012.45 | 88.422.99 | 88.363.05 | 96.540.90 | 96.660.85 | |
-| Informer + CIF | 89.381.66 | 89.711.37 | 89.381.66 | 89.351.68 | 96.920.66 | 97.030.64 | |
-| iTransformer | 74.691.02 | 74.761.04 | 74.691.02 | 74.671.02 | 83.351.24 | 83.651.41 | |
-| iTransformer + CIF | 72.791.12 | 72.891.16 | 72.791.12 | 72.761.11 | 81.081.03 | 81.310.93 | |
-| MTST | 77.673.58 | 78.974.37 | 77.673.58 | 77.453.55 | 86.474.84 | 84.996.43 | |
-| MTST + CIF | 87.211.99 | 87.302.06 | 87.211.99 | 87.201.99 | 94.141.83 | 93.242.63 | |
-| Nonformer | 88.102.39 | 88.761.74 | 88.102.39 | 88.042.47 | 96.560.91 | 96.361.21 | |
-| Nonformer + CIF | 88.002.06 | 88.911.68 | 88.002.06 | 87.922.12 | 97.101.10 | 97.171.09 | |
-| PatchTST | 77.982.64 | 79.303.73 | 77.982.64 | 77.762.65 | 86.674.03 | 84.935.47 | |
-| PatchTST + CIF | 79.580.86 | 80.220.82 | 79.580.86 | 79.470.88 | 87.201.57 | 85.811.17 | |
-| Reformer | 88.502.30 | 89.011.80 | 88.502.30 | 88.452.35 | 96.100.63 | 96.190.55 | |
-| Reformer + CIF | 89.081.19 | 89.470.70 | 89.081.19 | 89.051.23 | 96.880.35 | 97.000.35 | |
-| Transformer | 85.131.86 | 86.391.56 | 85.131.86 | 84.991.93 | 95.611.05 | 95.630.91 | |
-| Transformer + CIF | 89.961.57 | 90.531.26 | 89.961.57 | 89.921.60 | 97.730.45 | 97.770.46 | |
-| Medformer | 88.771.24 | 88.911.11 | 88.771.24 | 88.761.25 | 96.380.34 | 96.440.30 | |
-| Medformer + CIF | 90.880.87 | 90.940.81 | 90.880.87 | 90.870.87 | 97.390.30 | 97.460.28 | |
-| HM-BiTCN | 84.902.60 | 86.022.00 | 84.902.60 | 84.762.74 | 93.941.92 | 94.201.85 | |
-| HM-BiTCN + CIF | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 |
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| ADFTD (3-Classes) Reproduced | Autoformer | 46.902.89 | 45.592.37 | 44.912.23 | 44.342.52 | 63.492.44 | 45.632.29 |
-| Autoformer + CIF | 45.922.78 | 44.812.44 | 44.662.25 | 44.262.38 | 62.692.21 | 44.932.68 | |
-| Crossformer | 50.181.97 | 45.971.84 | 46.301.73 | 45.901.84 | 66.681.67 | 48.651.89 | |
-| Crossformer + CIF | 54.581.22 | 47.850.67 | 48.960.89 | 48.220.72 | 69.101.05 | 52.230.98 | |
-| FEDformer | 45.750.78 | 45.711.29 | 44.271.28 | 43.511.00 | 62.641.64 | 45.881.35 | |
-| FEDformer + CIF | 48.631.99 | 46.971.58 | 46.871.63 | 46.691.54 | 65.562.14 | 48.302.30 | |
-| Informer | 48.421.99 | 46.941.60 | 46.410.99 | 45.760.43 | 65.991.14 | 47.491.07 | |
-| Informer + CIF | 50.120.66 | 47.230.47 | 46.770.39 | 46.620.38 | 65.130.52 | 46.840.54 | |
-| iTransformer | 52.851.36 | 46.971.05 | 47.311.03 | 46.840.78 | 67.460.96 | 49.900.89 | |
-| iTransformer + CIF | 50.760.50 | 47.110.67 | 47.290.59 | 47.100.66 | 67.000.55 | 49.600.69 | |
-| MTST | 45.771.70 | 44.391.73 | 43.701.82 | 43.361.98 | 61.381.57 | 44.011.60 | |
-| MTST + CIF | 46.360.93 | 45.040.65 | 45.470.88 | 44.870.81 | 63.571.24 | 47.351.36 | |
-| Nonformer | 50.811.06 | 48.711.40 | 48.551.47 | 48.361.38 | 66.951.54 | 48.081.82 | |
-| Nonformer + CIF | 51.812.11 | 49.930.81 | 49.660.36 | 49.100.57 | 68.590.60 | 50.440.94 | |
-| Reformer | 51.282.60 | 49.682.75 | 49.642.02 | 48.452.06 | 69.202.53 | 51.743.24 | |
-| Reformer + CIF | 52.500.91 | 52.102.91 | 49.012.74 | 47.534.23 | 68.251.79 | 50.442.59 | |
-| Transformer | 50.530.94 | 49.310.87 | 48.571.23 | 48.421.28 | 67.980.90 | 49.071.35 | |
-| Transformer + CIF | 52.842.50 | 51.272.33 | 51.532.37 | 51.102.21 | 70.252.42 | 52.352.94 | |
-| Medformer | 53.701.18 | 51.511.32 | 50.491.48 | 50.351.53 | 70.481.17 | 50.911.13 | |
-| Medformer + CIF | 55.880.82 | 51.911.90 | 50.801.63 | 50.291.92 | 70.451.33 | 53.731.70 | |
-| HM-BiTCN | 52.052.22 | 50.453.00 | 50.402.55 | 49.482.70 | 69.432.84 | 50.993.15 | |
-| HM-BiTCN + CIF | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 |
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| PTB (2-Classes) | Autoformer | 71.992.74 | 69.603.85 | 61.504.23 | 61.435.07 | 74.291.89 | 70.262.00 |
-| Autoformer + CIF | 77.710.63 | 77.151.29 | 70.321.62 | 71.771.60 | 81.204.15 | 78.134.80 | |
-| Crossformer | 78.063.44 | 81.533.13 | 68.625.63 | 69.766.53 | 88.312.07 | 85.812.43 | |
-| Crossformer + CIF | 84.554.80 | 85.963.63 | 78.647.49 | 80.257.89 | 92.172.99 | 91.163.70 | |
-| FEDformer | 74.542.27 | 77.994.10 | 63.143.29 | 63.284.36 | 84.634.27 | 80.915.55 | |
-| FEDformer + CIF | 77.931.86 | 80.653.50 | 68.592.20 | 70.232.63 | 86.003.51 | 83.424.76 | |
-| Informer | 79.590.65 | 83.330.77 | 70.580.95 | 72.581.08 | 92.770.48 | 90.890.57 | |
-| Informer + CIF | 83.632.32 | 85.311.80 | 77.283.67 | 79.373.45 | 93.870.48 | 92.290.63 | |
-| iTransformer | 83.431.19 | 88.061.47 | 75.641.55 | 78.291.70 | 91.381.41 | 91.081.30 | |
-| iTransformer + CIF | 81.622.96 | 87.482.03 | 72.744.41 | 74.994.91 | 90.253.29 | 89.803.28 | |
-| MTST | 75.532.45 | 78.721.87 | 64.784.06 | 65.304.81 | 87.764.09 | 83.603.92 | |
-| MTST + CIF | 81.801.60 | 84.192.05 | 74.282.05 | 76.522.25 | 91.302.48 | 88.413.30 | |
-| Nonformer | 78.931.46 | 82.481.53 | 69.682.15 | 71.502.56 | 90.540.59 | 87.781.46 | |
-| Nonformer + CIF | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |
-| PatchTST | 75.282.44 | 77.052.44 | 64.864.05 | 65.415.28 | 88.112.59 | 82.652.87 | |
-| PatchTST + CIF | 80.801.90 | 82.461.75 | 73.232.81 | 75.242.95 | 92.450.91 | 88.231.61 | |
-| Reformer | 78.111.65 | 82.700.80 | 68.172.68 | 69.683.34 | 90.771.56 | 88.141.20 | |
-| Reformer + CIF | 81.952.09 | 84.631.07 | 74.443.53 | 76.553.64 | 93.360.87 | 91.531.05 | |
-| Transformer | 76.431.98 | 81.251.15 | 65.643.29 | 66.444.39 | 90.211.24 | 87.281.49 | |
-| Transformer + CIF | 78.802.44 | 82.452.36 | 69.423.62 | 71.104.33 | 92.501.17 | 89.071.39 | |
-| Medformer | 80.990.75 | 83.010.72 | 73.351.16 | 75.471.21 | 93.101.18 | 90.691.04 | |
-| Medformer + CIF | 78.740.64 | 81.110.84 | 75.400.66 | 76.310.71 | 83.200.91 | 83.660.92 | |
-| HM-BiTCN | 81.871.87 | 86.501.24 | 73.492.90 | 75.843.20 | 94.200.29 | 93.040.45 | |
-| HM-BiTCN + CIF | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 |
-During the process of transferring the optimized CIF hyperparameter configurations from the HM-BiTCN architecture to other model structures, we observed significant improvements in performance across most models and evaluation metrics.
-This result strongly demonstrates that the CIF module not only exhibits high generalizability but also possesses well-transferable hyperparameters that can be effectively adapted to various architectures.
-These hyperparameters enhance feature extraction and representation capabilities across different model designs.
-This finding further highlights the potential of the CIF architecture in diverse tasks and provides valuable insights for future model design and parameter sharing.
-F.3 Fine-tuning Transferred Parameters
-For models where the transferred hyperparameters from HM-BiTCN do not yield optimal results, we perform additional fine-tuning.
-This step investigates the adaptability of CIF-related parameters and explores how they can be optimized for other model structures.
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| TDBrain (2-Classes) Reproduced | iTransformer | 74.691.02 | 74.761.04 | 74.691.02 | 74.671.02 | 83.351.24 | 83.651.41 |
-| iTransformer + CIF(HM-BiTCN) | 72.791.12 | 72.891.16 | 72.791.12 | 72.761.11 | 81.081.03 | 81.310.93 | |
-| iTransformer + CIF(New) | 76.100.76 | 76.160.75 | 76.100.76 | 76.090.76 | 84.791.15 | 85.131.11 | |
-| Nonformer | 88.102.39 | 88.761.74 | 88.102.39 | 88.042.47 | 96.560.91 | 96.361.21 | |
-| Nonformer + CIF(HM-BiTCN) | 88.002.06 | 88.911.68 | 88.002.06 | 87.922.12 | 97.101.10 | 97.171.09 | |
-| Nonformer + CIF (New) | 88.671.30 | 89.340.93 | 88.671.30 | 88.611.33 | 96.890.31 | 96.910.37 |
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| ADFTD (3-Classes) Reproduced | Autoformer | 46.902.89 | 45.592.37 | 44.912.23 | 44.342.52 | 63.492.44 | 45.632.29 |
-| Autoformer + CIF((HM-BiTCN)) | 45.922.78 | 44.812.44 | 44.662.25 | 44.262.38 | 62.692.21 | 44.932.68 | |
-| Autoformer + CIF(New) | 48.491.31 | 46.421.07 | 45.391.80 | 44.941.97 | 63.751.46 | 46.101.84 | |
-| iTransformer | 52.851.36 | 46.971.05 | 47.311.03 | 46.840.78 | 67.460.96 | 49.900.89 | |
-| iTransformer + CIF(HM-BiTCN) | 50.760.50 | 47.110.67 | 47.290.59 | 47.100.66 | 67.000.55 | 49.600.69 | |
-| iTransformer + CIF(New) | 53.931.52 | 48.440.89 | 49.131.17 | 48.560.84 | 68.700.94 | 50.881.20 | |
-| Reformer | 51.282.60 | 49.682.75 | 49.642.02 | 48.452.06 | 69.202.53 | 51.743.24 | |
-| Reformer + CIF(HM-BiTCN) | 52.500.91 | 52.102.91 | 49.012.74 | 47.534.23 | 68.251.79 | 50.442.59 | |
-| Reformer + CIF(New) | 53.111.71 | 52.141.42 | 52.421.90 | 51.732.01 | 70.471.64 | 51.881.92 | |
-| Medformer | 53.701.18 | 51.511.32 | 50.491.48 | 50.351.53 | 70.481.17 | 50.911.13 | |
-| Medformer + CIF(HM-BiTCN) | 55.880.82 | 51.911.90 | 50.801.63 | 50.291.92 | 70.451.33 | 53.731.70 | |
-| Medformer + CIF(New) | 55.211.42 | 52.842.36 | 52.602.26 | 52.432.65 | 72.112.42 | 54.103.09 |
-| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
-| PTB (2-Classes) Reproduced | iTransformer | 83.431.19 | 88.061.47 | 75.641.55 | 78.291.70 | 91.381.41 | 91.081.30 |
-| iTransformer + CIF(HM-BiTCN) | 81.622.96 | 87.482.03 | 72.744.41 | 74.994.91 | 90.253.29 | 89.803.28 | |
-| iTransformer + CIF(New) | 83.561.15 | 86.001.03 | 76.781.94 | 79.101.87 | 90.400.75 | 89.800.93 | |
-| Nonformer | 78.931.46 | 82.481.53 | 69.682.15 | 71.502.56 | 90.540.59 | 87.781.46 | |
-| Nonformer + CIF(HM-BiTCN) | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |
-| Nonformer + CIF(New) | 80.462.20 | 84.222.44 | 71.793.10 | 73.893.50 | 92.740.97 | 90.301.26 | |
-| Medformer | 80.990.75 | 83.010.72 | 73.351.16 | 75.471.21 | 93.101.18 | 90.691.04 | |
-| Medformer + CIF(HM-BiTCN) | 78.740.64 | 81.110.84 | 75.400.66 | 76.310.71 | 83.200.91 | 83.660.92 | |
-| Medformer + CIF(New) | 83.980.81 | 85.210.62 | 78.001.30 | 80.111.22 | 93.320.97 | 90.971.69 |
-For models that did not initially benefit from the direct transfer of CIF hyperparameters optimized for HM-BiTCN, we conducted further fine-tuning of the key CIF parameters.
-The results demonstrate that, after targeted adjustment, these models also achieved notable performance improvements.
-This process not only reinforces the adaptability of the CIF architecture across diverse models but also highlights the high tunability and flexibility of its hyperparameters.
-By appropriately modifying the number of selected channels, fusion direction, and scaling factors, CIF can be effectively tailored to different network structures, thereby maximizing its strengths in feature modeling and discriminative capability.
+The Area Under the Receiver Operating Characteristic Curve (AUROC) measures the ability of a model to distinguish between classes, defined as
+where
+The Area Under the Precision–Recall Curve (AUPRC) summarizes the trade-off between precision and recall across different thresholds, defined as
+where
+Appendix B HM-BiTCN Structure Design and Theoretical Analysis
+Modeling short-term and long-term dependencies in time series data is challenging.
+Traditional CNNs excel at capturing local features but have limited receptive fields, hindering long-range dependency learning.
+Transformer-based methods effectively model long-term dependencies, but their complex design lacks interpretability, which is a key issue in medical time-series classification.To address these limitations, TCNs use causal convolutions for explicit temporal modeling and dilated convolutions to expand the receptive field, overcoming the constraints of traditional CNNs.
+Building on the advantages of TCN, we propose the HM-BiTCN, which combines the benefits of dilated convolutions, bidirectional causal convolution, and residual connections.
+This approach allows for better capture of temporal dependencies while preserving causality.
+B.1 Dilated Convolution
+Dilated convolution expands the receptive field without significantly increasing computational cost Yu & Koltun (2015).
+For a 1D input sequence , its output is defined as , where is the current time step, is the kernel size, is the dilation factor, and is the weight at the -th position in the kernel.
+Increasing effectively enlarges the receptive field, enabling the network to capture longer-term temporal dependencies.
+When stacking multiple dilated convolutional layers, the receptive field grows progressively.
+For the -th layer, the receptive field can be expressed as , where is the dilation factor of the -th layer.
+By gradually increasing , the network captures temporal dependencies across both global and local scales, offering an effective way to model long-term dependencies in time series.
+B.2 Bidirectional Causal Convolution Structure
+In addition to dilated convolutions, HM-BiTCN introduces a bidirectional causal convolution structure, inspired by prior bidirectional temporal modeling approaches Hanson et al.
+(2018); Hu et al.
+(2024); Yin et al.
+(2025).
+Unlike traditional TCNs that use only forward causal convolutions, our architecture applies causal convolutions in both forward and backward directions, enabling the model to capture dependencies from both past and future contexts while strictly preserving causality.
+The forward causal convolution processes the input sequence in chronological order, producing output , which depends only on current and past inputs.
+For the backward causal convolution, we first reverse the input sequence as , and then apply a causal convolution over this flipped sequence.
+This ensures that the model captures future-directed dependencies without introducing information leakage.
+The output is given by .
+These two operations are implemented using separate convolutional layers (convforward and convbackward), and their outputs are summed to form the final bidirectional result: .
+By integrating both directions under strict causality constraints, HM-BiTCN achieves superior temporal dependency modeling compared to unidirectional causal approaches.
+B.3 Multi-Scale Feature Learning and Residual Connections
+To further improve the model’s capacity to capture dependencies at different temporal scales, HM-BiTCN incorporates multi-scale feature learning and residual connections.
+Multi-scale Feature Learning: In HM-BiTCN, we employ a hierarchy of dilation factors that decrease layer by layer to capture temporal dependencies at multiple scales.
+Lower layers use larger dilation factors to expand the receptive field, aggregating long‐range information and smoothing short‐term noise in highly redundant medical time series; higher layers use smaller dilation factors to focus on local dependencies and capture fine‐grained features.
+This coarse‐to‐fine, global‐to‐local design enables the network to extract broad patterns in its initial layers and refine precise details in its later layers, thereby enhancing adaptability across a wide range of time series tasks.
+Residual connections: Residual connections He et al.
+(2016) are introduced between the dilated convolutional layers to facilitate the efficient flow of information through the network.
+The residual connection is defined as , where is the convolutional output, and is the input.
+This design alleviates the vanishing gradient problem and improves the overall stability of the network during training.
+Appendix C Pseudocode of CIF Method and Key Components of HM-BiTCN
+Appendix D Ablation Experiments of the HM-BiTCN Structure
+| Datasets | Models | CIF | Forward | Backward | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC | |||
+HM-BiTCN | ✓ | 82.312.34 | 83.292.50 | 80.392.65 | 81.022.63 | 91.501.80 | 91.661.82 | |||||
+| HM-BiTCN | ✓ | 79.453.51 | 80.693.04 | 77.144.47 | 77.584.75 | 87.953.82 | 88.413.74 | ||||||
+| HM-BiTCN | ✓ | ✓ | 82.491.40 | 82.381.79 | 81.201.32 | 81.601.39 | 91.101.63 | 91.301.71 | |||||
+HM-BiTCN | ✓ | ✓ | 80.435.60 | 80.465.23 | 79.565.98 | 79.505.97 | 89.234.44 | 89.624.25 | ||||
+| HM-BiTCN | ✓ | ✓ | 79.393.44 | 79.493.76 | 78.092.94 | 78.353.33 | 87.623.09 | 88.112.91 | |||||
+| HM-BiTCN | ✓ | ✓ | ✓ | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | ||||
+HM-BiTCN | ✓ | 53.321.35 | 52.011.54 | 51.462.19 | 51.211.99 | 70.781.78 | 53.162.20 | |||||
+| HM-BiTCN | ✓ | 52.801.18 | 50.160.77 | 49.231.22 | 49.241.02 | 68.650.71 | 49.950.97 | ||||||
+| HM-BiTCN | ✓ | ✓ | 52.052.22 | 50.453.00 | 50.402.55 | 49.482.70 | 69.432.84 | 50.993.15 | |||||
+HM-BiTCN | ✓ | ✓ | 56.060.47 | 53.211.03 | 53.541.36 | 52.821.33 | 72.930.88 | 55.711.03 | ||||
+| HM-BiTCN | ✓ | ✓ | 56.541.33 | 54.280.96 | 54.631.06 | 53.911.11 | 73.461.17 | 56.121.61 | |||||
+| HM-BiTCN | ✓ | ✓ | ✓ | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 | ||||
+HM-BiTCN | ✓ | 87.232.87 | 87.752.48 | 87.232.87 | 87.172.93 | 95.551.69 | 95.731.60 | |||||
+| HM-BiTCN | ✓ | 86.923.46 | 87.413.17 | 86.923.46 | 86.863.51 | 95.281.78 | 95.421.70 | ||||||
+| HM-BiTCN | ✓ | ✓ | 84.902.60 | 86.022.00 | 84.902.60 | 84.762.74 | 93.941.92 | 94.201.85 | |||||
+HM-BiTCN | ✓ | ✓ | 93.291.73 | 93.341.73 | 93.291.73 | 93.291.73 | 98.500.63 | 98.560.60 | ||||
+| HM-BiTCN | ✓ | ✓ | 93.691.52 | 93.831.42 | 93.691.52 | 93.681.53 | 98.560.67 | 98.590.64 | |||||
+| HM-BiTCN | ✓ | ✓ | ✓ | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 | ||||
+HM-BiTCN | ✓ | 82.561.74 | 86.161.51 | 74.912.88 | 77.242.92 | 95.690.64 | 94.560.76 | |||||
+| HM-BiTCN | ✓ | 81.074.24 | 85.362.71 | 72.506.59 | 74.336.71 | 92.832.38 | 91.282.79 | ||||||
+| HM-BiTCN | ✓ | ✓ | 81.871.87 | 86.501.24 | 73.492.90 | 75.843.20 | 94.200.29 | 93.040.45 | |||||
+HM-BiTCN | ✓ | ✓ | 87.331.41 | 90.261.24 | 81.642.04 | 84.191.97 | 96.211.30 | 95.671.52 | ||||
+| HM-BiTCN | ✓ | ✓ | 84.352.28 | 87.422.07 | 77.543.30 | 79.983.41 | 91.251.92 | 90.422.28 | |||||
+| HM-BiTCN | ✓ | ✓ | ✓ | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 | ||||
+HM-BiTCN | ✓ | 70.812.31 | 72.581.33 | 69.812.79 | 70.072.24 | 95.890.28 | 76.901.21 | |||||
+| HM-BiTCN | ✓ | 70.292.04 | 72.772.09 | 68.862.39 | 69.561.98 | 95.610.26 | 76.561.56 | ||||||
+| HM-BiTCN | ✓ | ✓ | 76.080.81 | 76.050.83 | 75.950.84 | 75.540.94 | 96.490.10 | 81.190.65 | |||||
+HM-BiTCN | ✓ | ✓ | 72.301.50 | 72.981.61 | 71.651.35 | 71.541.50 | 95.920.55 | 77.872.27 | ||||
+| HM-BiTCN | ✓ | ✓ | 72.811.04 | 74.050.80 | 71.861.25 | 72.121.17 | 96.200.21 | 79.221.25 | |||||
+| HM-BiTCN | ✓ | ✓ | ✓ | 76.821.32 | 77.380.85 | 76.521.24 | 76.391.18 | 96.480.06 | 81.770.81 | ||||
+HM-BiTCN | ✓ | 91.940.98 | 92.360.90 | 92.020.96 | 91.980.93 | 99.300.08 | 97.310.47 | |||||
+| HM-BiTCN | ✓ | 93.030.62 | 93.280.63 | 93.120.60 | 93.050.62 | 99.360.19 | 97.720.46 | ||||||
+| HM-BiTCN | ✓ | ✓ | 93.720.73 | 94.020.72 | 93.750.70 | 93.690.76 | 99.600.09 | 98.310.40 | |||||
+HM-BiTCN | ✓ | ✓ | 92.180.45 | 92.420.47 | 92.210.44 | 92.140.44 | 99.170.11 | 97.040.15 | ||||
+| HM-BiTCN | ✓ | ✓ | 92.620.90 | 92.880.86 | 92.680.88 | 92.630.89 | 99.250.18 | 97.150.57 | |||||
+| HM-BiTCN | ✓ | ✓ | ✓ | 93.780.32 | 94.080.26 | 93.790.32 | 93.740.34 | 99.340.19 | 97.600.46 |
+From the table 5, it can be observed that when both the Forward and Backward parts of the HM-BiTCN structure are used simultaneously, the performance drops significantly compared to using only one of them individually.
+We speculate that this is mainly due to the presence of substantial noise within medical time-series data.
+When both parts of the structure are applied at the same time, it is akin to capturing noise from two different directions simultaneously.
+Instead of enhancing the representation, this leads to noise accumulation, which ultimately results in degraded performance.
+However, after processing the data with CIF to improve the signal-to-noise ratio, the combination of the Forward and Backward parts of the HM-BiTCN structure eventually outperforms the use of either part alone.
+This result strongly demonstrates the feature-capturing capability of the HM-BiTCN when both directions are utilized together.
+It indicates that once noise interference is effectively reduced, the bidirectional structure of HM-BiTCN can better leverage its strengths, thereby improving overall performance.
+Further observations show that using only the Forward part of HM-BiTCN outperforms the Backward part.
+This is closely related to the inherent unidirectionality of medical time-series signals such as EEG and ECG, where information typically propagates forward in time (e.g., neural signal transmission in EEG or atrial-to-ventricular activation in ECG).
+Such characteristics enable the Forward structure to capture key features and temporal evolution more effectively, yielding better performance.
+This finding not only deepens the understanding of medical signal processing but also provides insights for optimizing HM-BiTCN in related applications.
+To evaluate the performance of our method on general time series, we follow the design of Medformer Wang et al.
+(2024a) and test it on two human activity recognition (HAR) datasets: FLAAP(13,123 samples, 10 classes) Kumar & Suresh (2022) and UCI-HAR(10,299 samples, 6 classes) Anguita et al.
+(2013).
+Additionally, on the non-medical datasets FLAAP and UCI-HAR, we observed that integrating the bidirectional structure significantly improves performance.
+This indicates that in high-SNR scenarios, bidirectional modeling can more effectively capture both forward and backward feature information, enhancing overall model performance.
+In contrast, CIF provides relatively limited gains on these high-SNR datasets.
+This observation further highlights the design advantage of CIF: it is specifically tailored for low-SNR medical time series, explicitly fusing inter-channel physiological information to enhance signal quality and discriminative power, while its marginal benefit is smaller for low-noise non-medical data.
+Overall, these findings not only reveal the differential adaptability of model architectures under varying data characteristics but also underscore the unique value of CIF in complex medical scenarios.
+Appendix E Further exploration of physiological structures
+The parameters in CIF are explicit hyperparameters that can be directly set and adjusted based on experience.
+For example, Figure.
+7 shows the corresponding locations of EEG channels on the human brain, we have adjusted the AFAVA dataset, which comprises 16 channels: C3, C4, F3, F4, F7, F8, Fp1, Fp2, O1, O2, P3, P4, T3, T4, T5, and T6.
+For the first six channels, we performed pairwise fusion as follows:
+Here, the channels C3, F3, and F7 exhibit physiological symmetry with C4, F4, and F8, respectively.
+We denote this type of fusion as Physiological Symmetry Fusion (PSF).
+In contrast, the channel fusion previously described in our paper, which was adjusted via the hyperparameter , lacked physiological symmetry and is referred to as Random Fusion (RF).
+| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC | |||
+HM-BiTCN + CIF(RF) | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | |||
+| HM-BiTCN + CIF(PSF) | 86.232.09 | 85.822.14 | 86.042.06 | 85.832.12 | 94.591.08 | 94.641.08 |
+The results in the table 6 reveal that explicit fusion leveraging the prior knowledge of channels can more effectively integrate channel features, thereby yielding more accurate classification outcomes.
+Many previous methods, especially various general time series models, are unable to incorporate such medical prior knowledge in a "controllable" manner.
+Appendix F Results on General Time Series
+Datasets / Models |
+RNN |
+TCN |
+Transformers |
+MLP |
+CNN |
+LSTM |
+LSTNet |
+LSSL |
+TCN |
+Trans. |
+Re. |
+In. |
+Pyra. |
+Auto. |
+Station. |
+FED. |
+ETS. |
+Flow. |
+iTrans. |
+DLinear |
+LightTS. |
+TiDE |
+TimesNet |
+Time Mixer++ |
+HM-BiTCN |
+HM-BiTCN
+EthanolConcentration |
+32.3 |
+39.9 |
+31.1 |
+28.9 |
+32.7 |
+31.9 |
+31.6 |
+30.8 |
+31.6 |
+32.7 |
+28.1 |
+31.2 |
+33.8 |
+28.1 |
+32.6 |
+29.7 |
+27.1 |
+35.7 |
+39.9 |
+31.9 |
+32.3 |
+FaceDetection |
+57.7 |
+65.7 |
+66.7 |
+52.8 |
+67.3 |
+68.6 |
+67.0 |
+65.7 |
+68.4 |
+68.0 |
+66.0 |
+66.3 |
+67.6 |
+66.3 |
+68.0 |
+67.5 |
+65.3 |
+68.6 |
+71.8 |
+66.8 |
+67.2 |
+Handwriting |
+15.2 |
+25.8 |
+24.6 |
+53.3 |
+32.0 |
+27.4 |
+32.8 |
+29.4 |
+36.7 |
+31.6 |
+28.0 |
+32.5 |
+33.8 |
+24.2 |
+27.0 |
+26.1 |
+23.2 |
+32.1 |
+26.5 |
+49.5 |
+51.2 |
+Heartbeat |
+72.2 |
+77.1 |
+72.7 |
+75.6 |
+76.1 |
+77.1 |
+80.5 |
+75.6 |
+74.6 |
+73.7 |
+73.7 |
+71.2 |
+77.6 |
+75.6 |
+75.1 |
+75.1 |
+74.6 |
+78.0 |
+79.1 |
+74.6 |
+77.5 |
+JapaneseVowels |
+79.7 |
+98.1 |
+98.4 |
+98.9 |
+98.7 |
+97.8 |
+98.9 |
+98.4 |
+96.2 |
+99.2 |
+98.4 |
+95.9 |
+98.9 |
+96.6 |
+96.2 |
+96.2 |
+95.6 |
+98.4 |
+97.9 |
+97.8 |
+98.3 |
+PEMS-SF |
+39.9 |
+86.7 |
+86.1 |
+68.8 |
+82.1 |
+82.7 |
+81.5 |
+83.2 |
+82.7 |
+87.3 |
+80.9 |
+86.0 |
+83.8 |
+87.9 |
+75.1 |
+88.4 |
+86.9 |
+89.6 |
+91.0 |
+82.6 |
+86.1 |
+SelfRegulationSCP1 |
+68.9 |
+84.0 |
+90.8 |
+84.6 |
+92.2 |
+90.4 |
+90.1 |
+88.1 |
+84.0 |
+89.4 |
+88.7 |
+89.6 |
+92.5 |
+90.2 |
+87.3 |
+89.8 |
+89.2 |
+91.8 |
+93.1 |
+89.7 |
+91.1 |
+SelfRegulationSCP2 |
+46.6 |
+52.8 |
+52.2 |
+55.6 |
+53.9 |
+56.7 |
+53.3 |
+53.3 |
+50.6 |
+57.2 |
+54.4 |
+55.0 |
+56.1 |
+54.4 |
+50.5 |
+51.1 |
+53.4 |
+57.2 |
+65.6 |
+61.6 |
+62.2 |
+SpokenArabicDigits |
+31.9 |
+100.0 |
+100.0 |
+95.6 |
+98.4 |
+97.0 |
+100.0 |
+99.6 |
+100.0 |
+100.0 |
+100.0 |
+100.0 |
+98.8 |
+96.0 |
+81.4 |
+100.0 |
+95.0 |
+99.0 |
+99.8 |
+99.5 |
+99.6 |
+UWaveGestureLibrary |
+41.2 |
+87.8 |
+85.9 |
+88.4 |
+85.6 |
+85.6 |
+85.6 |
+83.4 |
+85.9 |
+87.5 |
+85.3 |
+85.0 |
+86.6 |
+85.9 |
+82.1 |
+80.3 |
+84.9 |
+85.3 |
+88.2 |
+92.1 |
+92.8 |
+Average Accuracy |
+48.6 |
+71.8 |
+70.9 |
+70.3 |
+71.9 |
+71.5 |
+72.1 |
+70.8 |
+71.1 |
+72.7 |
+70.7 |
+71.0 |
+73.0 |
+70.5 |
+67.5 |
+70.4 |
+69.5 |
+73.6 |
+75.3 |
+74.6 |
+75.8 |
+We compared our method with various approaches on the general time series UEA dataset.
+The results of these methods were provided by TimeMixer++ Wang et al.
+(2025a).
+Experimental results show that CIF can enhance the performance of HM-BiTCN on general time series classification tasks.
+Moreover, the combination of HM-BiTCN with CIF achieves SOTA performance.
 Appendix G Limitations and Future Work
@@
 This integration should foster both deeper theoretical insights and stronger practical applicability, providing robust solutions for real-world medical problems.
-Furthermore, integrating domain-specific prior knowledge into medical time-series analysis can reveal and model relationships between channels more precisely.
-By combining medical expertise, clinical experience, and existing pathological data, we can enhance model interpretability and predictive performance, supporting clinical decision-making and interventions.
-Future research could focus on incorporating such prior knowledge to advance intelligent medical analytics.
+Furthermore, incorporating domain-specific prior knowledge into medical time-series analysis can more precisely reveal and model relationships between channels.
+By integrating medical expertise, clinical experience, and existing pathological data, the interpretability and predictive performance of models can be enhanced, thereby supporting clinical decision-making and interventions.
+On this basis, frequency-domain analysis Hu et al.
+(2025); Nason & Sachs (1999); Yi et al.
+(2025) offers an additional perspective: by applying Fourier transform or wavelet decomposition to the signals, physiological features at different frequency components can be identified, revealing patterns that are difficult to capture in the time domain.
+This is particularly valuable for noise reduction, extraction of periodic signals, and detection of pathological events, and can also provide richer feature representations for model inputs.
+Future research could further explore how to combine time-domain and frequency-domain information, integrating domain priors to improve the accuracy and robustness of intelligent medical analytics.
 Finally, we must acknowledge that the development trends in the field of artificial intelligence highlight the importance of architectural innovation.
@@
 For example, the local feature extraction capabilities of CNNs LeCun et al.
-[1989], the temporal stability of TCNs Bai et al.
-[2018] for long sequences, the long-term dependency modeling of RNNs Rumelhart et al.
-[1986] and LSTMs Hochreiter and Schmidhuber [1997], the global modeling efficiency of Transformers Vaswani et al.
-[2017], the resource-efficient computation of Mamba Gu and Dao [2023], and the hybrid recurrence-attention structure of RWKV Peng et al.
-[2023].
+(1989), the temporal stability of TCNs Bai et al.
+(2018) for long sequences, the long-term dependency modeling of RNNs Rumelhart et al.
+(1986) and LSTMs Hochreiter & Schmidhuber (1997), the global modeling efficiency of Transformers Vaswani et al.
+(2017), the resource-efficient computation of Mamba Gu & Dao (2023), and the hybrid recurrence-attention structure of RWKV Peng et al.
+(2023).
 By adapting and integrating these methods, we aim to build a powerful model that is not only deeply compatible with the CIF framework but also capable of efficiently handling complex medical time-series data.
-Appendix H Broader Impacts
-In this study, we take the first step toward a paradigm shift in medical time-series analysis, transitioning from a focus on model architecture improvements to emphasizing the intrinsic structure of the data.
-Our experimental results, which outperform several state-of-the-art methods across multiple datasets, validate the effectiveness of this approach.
-We emphasize that uncovering the essential characteristics in medical time-series data is a valuable research direction, particularly given the critical need for model interpretability in medical applications.
-However, practitioners should be cognizant of the method’s limitations.